A Complete Bibliography of Publications by, and about, Leopold Infeld

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

23 January 2025 Version 2.26

Title word cross-reference

$2.00 [Ano50a]. $2.50 [Kae38]. $25 [Goo98]. $3 [Mal41]. $3.50 [Nag48]. $7.50 [Ber61]. $8 [Ber62]. ^{e} [Bie73b]. D [Wol93]. δ [IP56d, IP57]. k [Inf61c]. U(1) [Wol91].

-approximation [Inf61c]. -functions [IP56d, IP57].

życiu [Inf50h, Inf51c, Inf98].

150e [Tit82]. 150th [Tit82]. 1907 [PT01]. 1933 [Szc33]. 1959 [Rox07]. 1986 [HS89]. 1995 [Flo08]. 1998 [Baż98].

21st [Edw12]. 21st-century [Edw12]. 2nd [Fra81].

9-10 [Inf74b].

Ē [EI78a].

A-secrets [Sch50b]. åsikter [EI38d]. Abelian [BdRS01]. Abenteuer [EI38f, EI49c, N.38]. Abhandlungen [Bor63]. Abolition [Ano55, BEI^{+}55b, RE63, Whi55]. abstract [Inf38b, Inf47f, IS49a, IS50]. Académie [Tit82]. Academy [Ber61, Inf56f, Tit82]. accelerated [Inf63e]. According [Inf53d, SW50, Baż56, STW53]. Achievements [Bel37]. Action [Inf57e, HI37, Inf36b, Inf37b, Inf57d]. addresses [DFKI41]. AdS [Li16]. adventure [EI38f, EI39c, EI49c, EI58, EI65a]. Aether [Inf52, Dir52]. Affairs [GR63]. Again [Ano50g]. Against [Foc53a]. Agar [Des46]. Age [GR63, Des46]. Agency [Ols61a]. ai [EI65b, EI13]. aideshi [Inf69a]. aimé [Inf57f]. Ajnstajn [Inf79a]. Al [Inf47a]. Albert [Ano39, Arm38, Bro39, Inf50b, Kae38, Len39, N.38, Qui40, Ano50a, Bič79, Inf50a, Inf50b, Inf53a, Inf53g, Inf54c, Inf55a, Inf55b, Inf56a, Inf56b, Inf57b, Inf57a, Inf58a, Inf60a, Inf61b, Inf61a, Inf63b, Inf68a, Inf74a, Inf79a, Inf79b, Inf84, Inf92, Nag50, Sch49a, Sch49b, Sch50a, C.51]. äldre [EI38d]. Alfred [Ber79a]. Algebraic [Arc11, Gar21, Gal30a, Ver34, Gal29]. algébrique [Gal29, Gal30a]. algébriques [Ver34]. alla [EI65b, EI13]. allgemeinen [Foc53b, IvdW33]. am [KIR62]. Among [Whi55]. Analyse [Gal29, Gal30a]. Analyses [Coh80]. Analysis [GK78, Gal29, Gal31a, Gal90, Gal30a]. Andover [HS89]. Angered [Ano50c]. Angle [Ano50b]. Angular [Coi56]. Anisfield [Ano42]. anniversaire [Bie73b, Tit82]. Anniversary [Tit82]. Antenna [Inf49b, IS56]. Antennas [Inf47c]. anti [Ano66]. anti-communist [Ano66]. antykomuniscie [Ano66]. Appeal [Ano55, BEI^{+}55a, BEI^{+}55b, RE63]. application [BR04, Inf50d]. Approach [Inf45, IS45a, IS46]. Appropriate [CC67]. approximation [Inf53f, Inf61c]. Arbeit [Inf28a, Inf30a]. Articles [GR63]. Artin [Ehr11c]. Ask [Whi55]. Associate [Ano50g]. Associated [IP56a]. asua [Inf57a, Inf60a, Inf61a]. ate [EI39b]. Atom [MH12]. Atomic [Ano50b, Des46, GR63, Inf46a, Sch54, Des46, Ols61a]. Atoms [Sal19]. attitude [KI60, KIR62]. attitude/ [KI60]. Auguste [Gal32b]. Ausgewählte [Bor63]. Authority [Ber68]. Authors [Ano42]. autobiography [Inf42b, Inf80, Inf06, Fra81, Roe81]. aux [EI38e, EI48a, EI63, EI78b, EI89b, EI93, EI15]. aventura [EI65a]. Awards [Ano42]. axisymmetric [Bre89].

B [Nag50, Rob38, IS46]. Back [Goo98]. Background [Gar21]. Ban [Whi55]. barricades [Tot93]. barricate [Tot93]. Based [HS89]. Basque [EI00]. Bazański [McC63]. Be [Liv05, Ano68a, Gel68, SB52]. be-Frihata [Gel68]. Beautiful [Goo98]. Bedeutung [IIK^{+}54]. Begriff [Inf57c]. Behind [Kae38]. Bemerkung [Inf30a]. Bemerkungen [Inf55d, Inf28a]. Benefits [Sch54]. Bessel [ISC47]. between [Bou11, Inf99]. Beyond [Inf48b]. Bhautiki [Inf65b]. bibliography [Inf70]. birth [Bie73a]. bis [EI61b]. black [HF11, Hen08, Li16]. black-body [HF11]. Bloc [Sch54]. Boden [Inf31b]. bodies [Inf54e, IP62a]. Body [IMT69b, HF11, Rob38]. bogov [Inf81]. bogów [Inf50h, Inf51c, Inf98]. Bomb [Des46]. Bondi [Rox07]. Book [Ano39, Ano50a, Arm38, Bad79, Bel48, Ber61, Ber62, Ber79a, Bor48, Bro39, C.51, Cai42a, Dav82, Des46, Fra81, Goo98, Inf39, Inf40a, Inf43a, Inf47b, Inf50c, Kac79, Kae38, Len39, Mal41, McC63, Mon41, N.38, Nag48, Pag65, Qui40, Roe81, Sch50a, Ste82, Tay79, Tro79, Cur12, Inf74b, Sil10, Coh80]. Books [Ano42, Gal31b]. Born [Bre89, Hen08, Kie11, Lom05, Spe12, ANN^{+}89, BdRS01, BB99, BS99, BR04, BI34e, BY05, CH03, CP13a, CP13b, CY14, CL07, Dir60, FOP02, HF11, Li16, MH12, MP14, PR13, Wol91, Wol93]. branes [Hen08]. Brans [Hen08]. Briefly [CCK50]. Broglie [Cas75, IS31, SI31, SI32]. Broglie’a [SI32]. Brooklyn [DFKI41]. Brownian [Inf40b]. Building [Sch54]. Bulletin [GR63]. Butsurigaku [EI40a].

C [Inf39]. calculation [Inf32b, Inf32a]. calculus [IP55b]. Calls [Ols61a]. Camille [Bre11]. can [SB52]. Canada [Bad79, Dav82, Kac79, dBR78, Inf78e, Ber79a]. Canadian [Ano50b, Ano50c]. candidat [Ehr08]. candidate [Ehr08]. Card [Ehr21]. Carnegie [Des46]. Cartesian [Inf61i]. Cartesian-like [Inf61i]. Cassius [Cai42a, Cai42b]. Catalan [Gal84]. causality [Inf30c]. Causation [Inf29, Inf30b]. centenario [Bie73a]. centenary [Bie73a]. Centennial [Baż98]. centuries [EI38a]. century [Edw12, SZI53]. certain [IP55a]. chaînon [Bou11]. champs [Inf32c]. Chaos [IR89]. characteristic [BR04]. Charge [IS31, Wol91, Wol93]. Charles [Ano50a, Gol11, Nag50]. Chevalier [Gal32b]. China [Sch54]. chmury [SI32]. choice [HI37]. Chosen [Inf81, Inf78a]. Christian [KIR62]. chrześcijaninem [KIR62]. CIP [Ano66]. city [Ric97]. Clarity [Kae38]. class [IP55a]. classes [Inf74b]. classical [Inf27, Inf28c, Inf31b, Inf53h, IP56f, Baż56]. classique [Inf27, Inf28c]. Clocks [Inf43b]. Cloud [Inf53d, SI31, SI32]. Co [Mal41, Inf61i, Des46]. Co-Operate [Des46]. co-ordinate [Inf61i]. cœur [Fer17]. Cold [Sal19]. Collaborator [Kae38]. collection [Ber62]. Comments [Inf28a]. Committee [Des46]. common [Inf28b]. commune [Inf28b]. Communist [Ols61a, Ano66, Ano66]. Community [Gel68]. concave [CP13a]. conceitos [EI39b]. Concept [Foc55]. concepte [EI57]. conceptos [EI65a]. Concepts [Bro39, EI38b, EI38c, EI42, EI66b, Kae38, EI38e, EI39b, EI47, EI48a, EI57, EI59, EI61d, EI62a, EI63, EI65a, EI65b, EI66a, EI71, EI78b, EI89b, EI93, EI94, EI98b, EI08b, EI11, EI13, EI15]. concetti [EI65b, EI13]. conditions [Gal0 , Inf53b, Inf54a]. Conducting [ISS49]. Conference [HS89, Inf64a, Pag65, Inf64a, Inf54d, Con15]. connection [Inf30c]. Constant [Inf36a, IS45b, Inf61h, Inf61j, Inf63d, Inf61g, Inf63a]. constitution [Ehr11c]. constraints [PR13]. contemporâneo [Inf57a, Inf60a, Inf61a]. contemporaneo [Inf63b, Inf68a, Inf74a, Inf92]. contemporary [Inf57a, Inf60a, Inf61a, Inf63b, Inf68a, Inf74a, Inf92]. Continued [Tan07, Gal29]. continues [Gal29]. Contours [Inf69b]. Contributions [IPS^{+}49]. Control [Des46]. Convergent [Kie11]. Cooperation [Sal19]. coordinate [Inf53b, Inf54a]. coordinates [Inf57g]. Copernic [Bie73b]. Copernican [Foc53b]. Copernico [Bie73a, Bie73a]. Copernicus [BBI^{+}65, Inf55h, Inf73]. Correspondence [Sta99]. Cosmology [Inf45, IS45a, IS46]. Couldn’t [Liv05]. Coulomb [Inf47e]. Covariance [Foc55]. covariant [IP56b]. CPES [Con15]. creation [Arc11]. Criticism [Foc53a]. Critics [Ano50b]. cuantelor [EI57]. cuantos [EI65a]. Cultural [Nag50]. Curbs [Ano64]. Curl [CP13b]. Curvature [IS45b]. Czech [EI58, Inf58a, Inf59b].

dai [EI65b, EI13]. d’analyse [Gal31a]. Danish [EI39a, EI61a]. dans [Inf27, Inf28b, Inf28c]. Darwin [Nag50]. data [CH03]. Dawn [Ale10, Sil10]. Dead [Ano68b]. death [Tit82, Inf53c, Inf55b]. Debate [Ols61b]. dedicated [Ber62]. Deduction [BI34a, BI34a]. defense [Inf54c]. del [EI65a, Inf57l]. dela [Inf79a]. delivered [DFKI41]. della [EI48b, EI65b, EI13, Inf63b, Inf68a, Inf74a, Inf92]. Demand [Ano64]. Demonstration [Gal29, Gal29]. Denies [Ano50b]. Denounce [Ano64]. Derivation [BY05, IP56f]. derivatives [Inf57h]. desarrollo [EI65a]. desde [EI39b, EI65a]. desenvolvimento [EI39b]. Destruction [Gel68, Inf54b]. desvoltarea [EI57]. d’Evariste [Con05, Coh80, Ano17, Ber02, BA62, BA76, Che32b, Dup96a, Dup96b, Dup03, Dup92, Gal84, Inf78b, Pic97, Pic51, Tit82, PT01]. Development [EI39a, EI62c, Inf55g, EI38d, EI39b, EI57, EI59, EI61a, EI62a, EI65a, EI65b, EI66a, EI98b, EI13, Inf53h, Lie95]. Developments [Ano62, Ber62, vK63, McC63]. développement [Lie95]. Dicke [Hen08]. dieux [Inf57f]. Differential [Arc11, MP14, Her81, Inf50d, PR13, SB52]. dílo [Inf58a]. dimension [Inf38c]. dimensions [Wol93]. dioses [Inf78a]. dipole [IP56e]. Dirac [BI34a, Inf32b, Inf32a, Inf34a, Inf53d, Inf53f, IP56d, IP57]. Diracschen [Inf32a]. Directed [Ano68b]. Disciplines [Inf61f]. Discourses [Ano42]. Discovered [Liv05]. Discovery [Des46]. Discussion [GK78, Gal90]. Dlaczego [KIR62]. dlja [Inf74b]. do [EI59, EI62a, EI62c, EI66a, EI98b, Inf55h]. dobrodruzstvi [EI58]. doc [Szc33]. d’ondes [BI34a]. Doran [Mal41]. Doubleday [Mal41]. Doubt [Sch50b]. Douglas [Cai42a, Cai42b]. d’ouvrages [Coh80]. dr [Szc33]. Dr. [Ano42, Ano50f]. Drie [EI38a]. Drogi [Szc33, Inf33, Inf57i]. Dublin [Inf49c]. Duel [Ale10, Liz16, Sil10]. duelo [Liz16]. d’un [Gal29, Gal30a]. d’une [Ehr11b]. Dupuy [Ber02]. Dutch [EI38a]. Dvizhenie [IP62a]. dvizhenija [Inf64c]. Dynamics [Inf58b]. Dzieciece [Nat32]. dzieło [Inf56a].

E451 [BTC^{+}41]. E456 [IA41a, BTC^{+}41, IA41b]. earliest [EI59, EI62a, EI98b]. Early [Bro39, EI38b, EI38c, EI42, EI66b, Kae38, Pie98, EI47, EI61d, EI71, EI94, EI08b, EI11, Sal19]. eboluzioa [EI00]. École [Ehr08]. Écrits [BA62, BA76, Gal97]. ed [Fra81]. Edited [Ber79a, Dav82, Pag65]. eeuwen [EI38a]. effect [SI32]. effects [BS99]. efforts [Arc11]. ego [Inf64c]. eigenvalue [Inf41a, Inf42a]. EIH [Inf61c].

Einfluß [Inf53a, Inf56b, Inf57b]. einheitlichen [Inf28d]. Einige [Inf55d]. Einshteine [Inf56c]. Einstein [Ano39, Ano50a, Ano06, Arm38, Bro39, C.51, Inf53a, Inf56b, Inf57b, Inf69b, Inf99, Kae38, Len39, N.38, Nag50, Qui40, Rob38, Sch50a, Ano50g, Ano68b, Ban51, Bič79, BI67, BI69, Bre89, Cas75, CC67, GMW80, HS89, Inf31b, Inf47d, Inf50a, Inf50b, Inf50g, Inf50e, Inf50f, Inf51b, Inf51a, Inf53a, Inf53g, Inf54c, Inf55a, Inf55b, Inf55h, Inf56a, Inf56b, Inf56c, Inf56d, Inf57b, Inf57a, Inf58a, Inf59b, Inf60a, Inf61b, Inf61a, Inf61d, Inf63b, Inf68a, Inf69b, Inf73, Inf74a, Inf79a, Inf79b, Inf83, Inf84, Inf92, Inf99, Inf15, Kae38, Lom05, Mal79, Sch49a, Sch49b, Sta99, Whi55, Inf50b]. Einsteina [Inf55h]. Einsteinie [Inf56d]. Einsteinschen [Inf31b]. Elected [Rox07]. Electric [Inf53d]. électricité [Inf28b]. Electricity [Inf28d, Inf28b, Inf28e]. Electrodynamics [BB99, Dir60, Inf36a, IW40, IP53, IP54, BI34a, BI34d, Inf37a, Inf53h, Inf53f, IP56f, STW53, SW50]. electrodynamique [BI34a, BI34d]. Electromagnetic [BI33a, Inf38a, Inf38b, Vaz77]. electromagnetism [BY05]. Electron [BI34e, IvdW33, BS99, Inf31a, Inf64c, SI32]. Electronic [Inf53d]. electrons [SI31]. electrostatic [FOP02]. elegido [Inf78a]. Elektrizität [Inf28e, Inf28d]. Elektrodynamik [Inf53h]. Elektrodynamika [SW50, STW53]. Elektronenwelle [Inf31a]. elektronowej [SI32]. Elektrons [IvdW33]. Elementary [BTC^{+}41, IA41b]. elements [SW79]. Eltern [Ric97]. Embattled [Sal19]. Endowment [Des46]. Energy [BR04, Des46, Inf46a, Inf67, Ols61a, Des46]. Engineering [SBS^{+}61]. England [Inf49d]. entre [Bou11]. Entwicklung [Inf53h, IIK^{+}54]. Equation [IvdW33, Inf34a, Liv05, ANN^{+}89, BI34a, Inf32b, IP56c, Inf60c, SB52, BI34a]. Equations [Arc11, BI34c, BI63b, IW40, IS49a, Inf50e, Inf50f, IS51, Inf54a, Inf55e, Inf56e, IP56a, Inf57e, Inf59a, MP14, SI49, BR04, Bre89, CH03, CL07, EIH38a, EIH38b, EI40b, FOP02, Gal30a, Gal0 , Gal30b, Gal32a, Inf28b, Inf32a, Inf50d, Inf51a, Inf53b, IP56b, IP56f, Inf57c, Inf57j, Inf57d, Inf64c, Kie11, Pie98, Rob38, Ver34, EI49a, Gal30a, Gal30b, Gal32a, Inf28b, Ver34]. Erinnerung [Inf69b]. Erinnerungen [BI67, BI69]. Erkenntnis [EI38f, EI49c, N.38]. Erratum [Kos95a]. espace [Inf27, Inf28c]. Euler [BR04]. Evarist [Inf81]. E‘variste [Tit82, Bel48, Bor48, Cur12, Inf53i, Inf54f, Inf55i, Inf57m, Nag48, PT01, SC06, Tay79, Tro79, Ano17, Ast94, Ber02, BA62, BA76, Bre11, Che32b, Con05, Dal56, Dup96a, Dup96b, Dup03, Dup92, Ehr08, Ehr11b, Ehr11a, Fer17, Gal84, Inf48c, Inf53i, Inf54f, Inf55i, Inf57f, Inf57m, Inf61e, Inf74b, Inf78a, Inf78b, Inf78d, Inf81, Liz16, Neu11, Pic97, Pic51, PT01, Sar37, Tan06a, Tan06b, Tan07, Tan08, Tit82, Tot93, Tot96, Ver34, WA11, Gal11a, PT01]. eventyr [EI39c]. Everist [Inf60b, Inf74b]. Evoliutsiia [EI66a]. evolucão [EI39b, EI08a]. Evolutia [EI57]. Evolution [Ano39, Arm38, Bro39, EI38b, EI38c, EI42, EI50b, EI50a, EI56, EI59, EI61b, EI62a, EI66a, EI66b, EI68, EI78a, EI84, EI87, EI88, EI89a, EI92, EI95, EI98b, EI98a, EI04, EI08a, Inf65b, Kae38, Len39, Mal41, Mon41, Qui40, Ste82, EI38e, EI39b, EI47, EI48b, EI48a, EI61c, EI61d, EI63, EI65b, EI71, EI78b, EI89b, EI93, EI94, EI08b, EI11, EI13, EI15, Inf41b, Inf41c, Inf89, EI57, EI00]. evrimi [EI94, EI11]. Ewarysta [Inf50h, Inf51c, Inf98]. Ewolucja [EI59, EI62a, EI98b]. Ex [Ano50g]. Ex-Associate [Ano50g]. example [KI60, KIR62]. exelixē [EI78a]. existence [CH03]. Expansion [IP56a]. Explains [Kae38]. exposé [Tit82]. Extended [CP13b, CP13a, Pen99].

fabrication [Ehr11b]. Factorization [Coi56, HI48, IH51, Her81, Inf42a, Inf50d]. factorized [SB52]. fait [Tit82]. fali [SI32]. Fano [Ehr21]. Feature [Ban51]. Feldtheorie [Inf28e, Inf28d]. Ferment [Ols61b]. Festschrift [Inf58b]. Field [Ban51, Ber68, BI34c, BI63a, BI63b, Inf28d, IS49a, Inf53d, SI49, Tem01, Baż56, BI33b, BI34b, BI35, CL07, HI37, Inf28e, Inf32c, Inf36b, Inf37b, Inf55e, Inf61g, Kie11, Vaz77]. field-theories [Inf55e]. Fields [Gol11, Ols61b, FOP02]. fifth [Bie73a]. fig [Szc33]. Figure [Sch50b]. filosofia [Cas75]. final [Liz16]. Fine [Inf36a]. Finnish [EI62b]. first [Bon13, EI38e, EI48a, EI57, EI63, EI65a, EI78b, EI89b, EI93, EI15, GMW80]. Fisica [Cas75, EI39b, EI48b, EI65a, EI65b, EI13, EI68, EI84, EI08a]. Fisikaren [EI00]. fizicii [EI57]. Fizigin [EI94, EI11]. fizike [EI62c]. fiziki [EI66a]. fizyka [Inf68b]. fizyki [EI59, EI62a, EI98b]. fizyków [Inf54d]. Flowering [Gel68]. fluid [Bre89]. Fock [Inf53g]. Force [Nag50]. Foreword [Ber79a]. form [Inf60c]. formalism [Con05]. formalisme [Con05]. formation [KI60, KIR62]. forms [SB52]. Formula [Inf67]. formulas [Inf47e]. formulation [IP56b]. Fought [Ano42]. Foundations [BI33b, BI34b, BI63a]. fourth [Inf38c]. fra [EI39a, EI61a]. från [EI38d]. fractions [Gal29]. Franck [WHS68]. francuzskom [Inf74b]. Franklin [Cai42a, Cai42b]. free [KI60, KIR62]. free-thinking [KI60, KIR62]. Freedom [Ano64, Inf61f, Inf62, Ols61a, Ols61b]. French [Inf74b, Ano17, Arc11, Ber02, BI34a, BI34d, Bou11, BA62, BA76, Che32b, Coh80, Con05, Dal56, Dup96a, Dup96b, Dup03, Dup92, Ehr08, Ehr11c, Ehr11b, EI38e, EI48a, EI63, EI78b, EI89b, EI93, EI15, Fer17, Gal29, Gal30a, Gal0 , Gal31a, Gal30b, Gal30c, Gal31b, Gal32a, Gal46, Inf27, Inf28b, Inf28c, Inf32c, Inf57f, Inf78b, Lie95, Pic97, Pic51, PT01, Tan06a, Tan06b, Tan07, Tan08, Tit82, Ver34, WA11]. Frenkel [BI34e]. Frieden [Inf54b]. Frihata [Gel68]. Front [Kae38]. FRS [Rox07]. fulgurante [Fer17]. function [HI37, Inf36b, Inf47e, Inf57g]. Functionals [CP13b]. Functions [ISC47, IP56a, Inf37b, IP56d, IP57]. Fundamental [Inf53e]. further [IP57]. Fysiikan [EI62b]. Fysikens [EI38d]. Fysikken [EI39c]. Fysikkens [EI39a, EI61a]. Fyzika [EI58].

G [Ano42, Inf30a]. G.R [IS49a]. G.R.T. [Inf63c]. Galilean [CC67]. Galilei [EI38a, EI39a, EI61a]. Galileo [EI38a, EI39a, EI61a]. ’Galois [Ehr11c, Cai42b, PT01, Ano17, Arc11, Ast94, Ber02, Bir37, Bon13, Bou11, BA62, BA76, Bre11, Che32b, Con05, Dal56, DFKI41, Dup96a, Dup96b, Dup03, Dup92, Edw12, Ehr21, Ehr08, Ehr11c, Ehr11b, Ehr11a, Fer17, Gal32b, Gal84, Gal11a, Gar86, Gar21, Gol11, Her81, Inf48c, Inf50h, Inf51c, Inf53i, Inf54f, Inf55i, Inf57f, Inf57l, Inf57m, Inf61e, Inf74b, Inf78a, Inf78b, Inf78d, Inf81, Inf98, Kis07, Kos95b, Lie95, Lie32, Liz16, Nag48, Neu11, Pic97, Pic51, Pie98, Sar37, Tan06a, Tan06b, Tan07, Tan08, Tit82, Tot93, Tot96, Ver34, WA11, Bel48, Bor48, Coh80, PT01, Tay79, Tro79, Cai42a, Cur12, Kos95a, SC06]. Galua [Inf60b, Inf74b, Inf81]. Gamow [Inf48b]. Gap [Inf47c, IS56]. Garoa [Inf69a]. gauge [Wol91]. Gauss [Bou11]. General [Ano62, Ber62, EI49a, EI49b, HS89, IvdW33, Inf34a, Inf46b, IS49b, Inf56e, vK63, McC63, Rob38, SI49, Foc53b, Inf54e, IP56e, Inf57d, IP62a, Inf63c, Pen99, Hir74, Inf57e]. generalised [Inf32b]. generalization [Inf42a]. Generalized [Ban51, Inf65c, Inf32a]. Genius [Bel48, Liv05, Nag48]. Geocentric [Anoxxa]. Geodesic [IS49a, SI49, Inf60c]. geometer [Dal56]. géomètre [Dal56]. Geometrie [Inf31c, Inf57c]. geometry [Inf31b, Inf31c]. George [Inf48b]. German [Bor63, BI67, BI69, EI38f, EI49c, EI50a, EI56, EI61b, EI87, EI88, EI89a, EI92, EI95, EI98a, EI04, Foc53b, Foc54, Inf28a, Inf28e, Inf28d, Inf29, Inf30a, Inf30b, Inf31b, Inf31c, Inf31a, Inf32a, IvdW33, Inf53a, Inf53h, Inf53i, Inf54b, Inf54f, Inf55d, Inf55c, Inf55i, Inf56b, Inf57b, Inf57c, Inf57m, Inf69b, Inf99, Ric97]. Geschichte [Inf53i, Inf54f, Inf55c, Inf55i, Inf57m]. Gleichungen [Inf32a]. Global [CH03]. Goddesses [Inf50h, Inf51c, Inf98]. Gods [Bel48, Bor48, Inf48c, Inf78d, Inf81, Nag48, Tay79, Tro79, Inf53i, Inf54f, Inf55i, Inf57f, Inf57m, Inf78a]. Gordon [IP56a, IP56c]. Götter [Inf53i, Inf54f, Inf55i, Inf57m]. Government [Inf46a, Des46]. Grandfather [Inf15]. graph [Kos95a, Kos95b]. Gravitation [Inf28d, Inf64a, Inf28b, Inf28e, Inf28e, Inf28d, Inf64a, Pag65]. Gravitational [Inf38a, Inf38b, IS50, IS51, Inf59a, IMT69b, SI50, CY14, EIH38a, EIH38b, EI40b, Inf61g, Rob38]. gravity [Inf28b]. Great [Bel37]. Greek [EI78a]. Ground [MH12, Inf31b]. Group [Bir37, Inf37b, SW79, Wol91]. groups [Lie32]. Growth [Bro39, EI38b, EI38c, EI42, Kae38, EI61d, EI71]. Guide [Inf49b, ISS49]. Guides [IPS^{+}49].

H [Des46, MH12]. H-Atom [MH12]. Haitian [Ano42]. hajimerareta [EI40a]. Halecki [Ano66]. Haleckim [Ano66]. Hamiltonian [ANN^{+}89]. Hamming [Ehr21]. Hard [Sch50b]. harmonic [Inf54a]. hasta [EI65a]. Hattori [Inf28a]. Hawking [Li16]. Heads [Ols61a]. heart [Fer17]. Hebrew [Gel68]. Heisenberg [Cas75]. held [Baż98]. Helen [Bad79, Ber79a, Dav82]. Hermann [Rox07]. Hermite [Gol11]. Heroes [Ale10, Sil10]. Herrn [Inf28a, Inf30a]. Hertha [WHS68]. hidden [CL07]. Higgs [Mal79]. higher [Inf57h]. Hill [HS89]. him [Gaj99]. Himself [Kae38]. Hindi [Inf65b]. histoire [Fer17]. historia [Inf78a]. History [HS89, Inf55c, Inf55f, Pie98, Inf55g]. hito [Inf69a]. hits [Ano50b]. Hoffman [Mal79]. Hoffmann [Rob38, CC67]. hole [Li16]. Homogeneous [Inf53d]. Honored [Ano42]. honour [Pen99]. hope [Ric99]. hours [Inf57l]. House [Nag48]. Hull [Her81]. hundred [GMW80]. Hurbana [Gel68]. Hydrogen [HI48]. Hyman [Inf40a].

icône [Ehr11b]. icon [Ehr11b]. icosahedron [Kos95a, Kos95b]. ideōn [EI78a]. idea [EI38d]. Ideas [Bro39, EI38b, EI38c, EI42, Kae38, EI38e, EI39b, EI48a, EI57, EI59, EI61d, EI62a, EI63, EI65a, EI65b, EI66a, EI71, EI78b, EI89b, EI93, EI98b, EI13, EI15]. idee [EI65b, EI13]. idées [EI38e, EI48a, EI63, EI78b, EI89b, EI93, EI15]. idei [EI66a]. ideias [EI39b]. ideilor [EI57]. idéutveckling [EI38d]. II [BI35, EI40b, Inf28c, Inf30b, Inf37b]. ikani [EI40a]. iliskinlige [EI94]. ilk [EI94, EI11]. Illus [Ber61]. im [Foc53b]. immortalità [Inf57l]. immortality [Inf57l]. Impact [Des46, Inf79a]. Impedance [Inf49b]. Inclusion [MP14]. Indispensable [Ols61a]. individual [KI60, KIR62]. indywidualny [KI60, KIR62]. inédits [PT01, Tan06a, Tan06b, Tan07]. Infeld [Bad79, Baż98, Ber61, Ber79a, Cas75, Coh80, Dav82, Des46, Kae38, Mal41, Mal79, Nag48, Rob38, Sch50a, WHS68, A.68, Ano39, Ano50a, Ano50e, Ano50d, Ano50f, Ano66, Ano68b, Ano08, Anoxxa, Anoxxb, ANN^{+}89, Arm38, Baż56, Bel48, Ber68, Ber79b, BdRS01, BB99, Bič79, BS99, BR04, Bor48, BY05, Bre89, Bro39, C.51, Cai42a, Cai42b, CH03, CC67, CP13a, CP13b, CY14, CL07, Coi56, Dir52, Dir60, FOP02, Fra81, Gaj99, HF11, Hen08, Her81, Inf70, IBBT78, Kac79, Kie11, Len39, Li16, Lom05, MH12, Mon41, MP14, N.38, Pag65, PR13, Pen99, Qui40, Ric99, dBR71, Sch50b, Spe12, Sta99, SB52, Ste82, SW50, STW53, Szc33, Tay79, Tem01, Tro79, Wol91, Wol93, Ber62, Nag50]. Infelda [Baż56, SW50, STW53]. Infield [Ano42]. Infinity [Inf48b, Inf48a]. influência [Inf57a, Inf60a, Inf61a]. Influence [C.51, IS31, Inf47c, Lie95, Nag50, Sch50a, Ano50a, Inf50a, Inf53a, Inf56a, Inf56b, Inf57b, Inf57a, Inf58a, Inf60a, Inf61b, Inf61a, Inf61d, Inf63b, Inf68a, Inf74a, Inf83, Inf92, SI31, Lie95]. influencia [Inf61d, Inf83]. influenza [Inf63b, Inf68a, Inf74a, Inf92]. initial [CH03, EI65b, EI13]. iniziali [EI65b, EI13]. Injustice [Ano68a]. Institute [DFKI41, Ano68b]. Intellectual [Ols61b]. Intensities [HI48]. International [Des46, Pag65]. internationale [Inf64a]. Interpretation [Inf31b]. Introduction [Dav82, Ber79a, Inf43a]. inverse [Mal79]. Island [DFKI41]. Istoriya [Inf55g]. Italian [Cas75, EI48b, EI65b, EI13, Inf57l, Inf63b, Inf68a, Inf74a, Inf92, Tot93]. IV [Inf49b]. Izbrannik [Inf81]. izluchajushhego [Inf64c].

J [Ano42, Coh80, Szc33]. ja [Inf68b, KI60, KIR62, KIR81]. Jabłonna [Inf64a]. Jackson [Cai42a, Cai42b]. Jahre [WHS68]. jako [EI58]. Japanese [EI40a, Inf69a]. jazyke [Inf74b]. jego [Inf56a]. jeho [Inf58a]. jelektrona [Inf64c]. Jerzy [Ber61]. Jesse [Cai42a, Cai42b]. jestem [KIR62]. Jewish [Ano42]. Jews [AJP86]. John [Ano42, Flo08]. Jordan [Des46, Bre11]. Josef [WHS68]. Joseph [Inf53c, SC06]. juin [Tit82]. July [Inf64a]. June [Baż98, Tit82]. jüngste [Inf53h]. jusqu’à [Ehr11c].

kappa [BdRS01]. kappa-symmetry [BdRS01]. Kausalgesetz [Inf29, Inf30b]. kavramlardan [EI94, EI11]. KCB [Rox07]. Kehilat [Gel68]. kehitys [EI62b]. Kepler [IS45b]. Keyser [Cai42a, Cai42b]. Kinematic [IS45a, IS46]. kinlige [EI11]. kkamigami [Inf69a]. klassach [Inf74b]. klassischen [Inf31b, Inf53h]. Klasyczna [Baż56]. Klein [IP56a, IP56c]. Kniga [Inf74b]. knowledge [EI38f, EI49c, EI58]. Komunizujacy [Ano66]. konferencji [Inf54d]. Konin [Gel68, Ric95, Ric96, Ric97]. Kontur [Inf69b]. Kopernik [BBI^{+}65]. Kopernika [Bie73b, Inf55h]. Kopernikanische [Foc53b]. Kordian [Inf68b]. kształtowania [KI60, KIR62]. kuantumlara [EI94, EI11]. kvantemekanik [EI61a, EI39a]. kvantne [EI62c]. kvantov [EI66a]. kvantteorierna [EI38d]. kvanttiteoriaan [EI62b]. kwantów [EI59, EI62a, EI98b]. Kwanty [Szc33, Inf33].

L. [Baż56]. Ladder [SW79]. Lagrangian [Inf57g, Inf57h, Inf57k, Inf63c, Inf64c]. lagranzhian [Inf64c]. Language [Liv05, Inf74b]. last [Kos95a, Kos95b]. lata [Nat32]. Later [Inf50b]. Laws [Inf53e]. Leave [Ano50d, Ano50g]. Leben [Inf99, Inf69b]. Lecture [Pen99]. Lectures [Ano50b, DFKI41, SW50, STW53, Cai42a, Cai42b]. Left [Bad79, Dav82, Kac79, dBR78, Ber79a, Inf78e]. Legal [Des46]. l’Enseignement [Gal31b]. Leonardo [Inf53e]. Leopold [A.68, Ano39, Ano42, Ano50a, Arm38, Bad79, Bel48, Ber61, Ber62, Ber79a, Bor48, Bro39, C.51, Cai42a, Cai42b, Coh80, Des46, Fra81, Kac79, Kae38, Len39, Mal41, Mon41, N.38, Nag48, Nag50, Pag65, Qui40, Rob38, Ste82, Szc33, Tay79, Tro79, Ano66, Ano68b, Ano08, Anoxxa, Anoxxb, Ber68, Bič79, Dav82, Dir52, Gaj99, Inf70, IBBT78, Pen99, Ric99, dBR71, Sch50a, SW50, STW53, Tem01, WHS68]. Leopolda [SW50, STW53]. Lere [Inf47a]. Letter [BI33a, BI33b, Inf45, Inf48a, Inf50e, Inf52, IS56, Inf61h, Kos95a, Kos95b]. Lettre [Gal32b]. L’evolució [EI84]. L’evolución [EI68]. L’Évolution [EI38e, EI48a, EI63, EI78b, EI89b, EI93, EI15]. L’evoluzione [EI65b, EI13, EI48b]. Levy [Inf40a]. Lewis [Bad79, Ber79a, Dav82]. Leyburn [Ano42]. liang [Inf36c]. Liberty [Ols61a]. library [SZI53]. Lichte [Foc53b]. lieben [Inf53i, Inf54f, Inf55i, Inf57m]. Life [Inf50h, Inf51c, Inf98, Nag48, Ano17, Ber02, Dup96a, Dup96b, Dup03, Dup92, Fer17, Inf57l, IBBT78, Inf99, Nat32, Tit82, Tot93]. Light [Inf21b, Foc53b, Inf21a]. Lighton [Flo08]. like [Inf61i]. line [Inf60c]. linear [Inf55e]. link [Bou11]. Liouville [SC06]. Lives [Bel37]. Living [Inf69b]. Localized [Vaz77]. Long [DFKI41]. looking [Ric97]. Lorentz [Inf37a]. Loria [Inf58c]. Loss [Ano50e]. lost [Ric99]. Louis [Bou11]. Love [Bel48, Bor48, Inf48c, Inf78d, Nag48, Tay79, Tro79, Inf53i, Inf54f, Inf55i, Inf57m]. loved [Inf57f]. lower [CP13a].

M. [Gal32b]. Macmillan [Ber62]. made [Tit82]. Madness [Goo98]. Magnetic [Inf53d, Wol91]. Magnetostatic [CP13b, CP13a]. Makers [SZI53]. man [Inf63b, Inf68a, Inf74a, Inf92, Inf99]. Mandarin [Inf36c]. Manifesto [Ano06, Des46]. Mann [Inf99]. manquant [Bou11].

Manuscripts [Gal11b, PT01, Tan06a, Tan06b, Tan07, Tan08]. Manuscrits [PT01, Tan06a, Tan06b, Tan07, Tan08]. March [Flo08]. Martyrs [Ale10, Sil10]. mass [BI33a]. Massachusetts [HS89]. Matematica [Tot93]. matematico [Inf57l]. materia [Inf57i]. Material [Inf36c]. Materials [Inf54d]. Materiały [Inf54d]. materja [Inf33, Szc33]. Mathematical [BA62, BA76, Che32b, Cur12, Liv05, Nag48, Neu11, CL07, Ehr11b, Gal46, Pic97, Pic51, PT01, SC06]. Mathematician [Sch50b, Fer17, Inf57l, WA11]. Mathematicians [Bel37]. mathématicien [Fer17, WA11]. Mathematics [Ale10, DFKI41, Ehr11a, Inf48b, Bel37, Lie95, Goo98, Tot93, Sil10]. Mathematik [IIK^{+}54]. mathématique [Ehr11b]. Mathématiques [BA62, BA76, PT01, Che32b, Con15, Gal46, Gal89, Gal97, Lie95, Pic97, Pic51]. Mathesis [Szc33]. matrix [SW79]. Matter [Tem01, Inf33, Inf34b, Inf57i]. Maxwell [Inf28b, Kie11, Nat32, Spe12]. Maxwella [Nat32]. May [HS89, Sch54]. McGill [Dav82]. McGill-Queen [Dav82]. Mé [Inf59b]. measurement [Inf27, Inf28c]. measurements [Inf30c]. Mechanical [EI64, Inf57g, Inf57h]. Mechanics [EI39a, Inf29, IMT66, EI61a, Inf30a, Inf30b, Inf30c, Inf31b, Inf39]. Mechanik [Inf31b]. mechanism [Mal79]. Meeting [Baż98, Tit82]. meiner [Ric97]. Meixner [WHS68]. membrane [BY05]. Memoir [Kac79, Bon13, Gal30a, Gal0 ]. Mémoire [Gal0 , Gal30a]. Mémoires [BA62, BA76, Gal97]. Memoirs [BA62, BA76]. Memories [Inf56c, Inf56d, Inf59b, Inf64b, Inf66, Inf68b, BI67, BI69]. memory [Inf69b]. Men [Bel37, Inf47a]. mesure [Inf27, Inf28c]. meteoric [Fer17]. Method [HI48, IH51, Inf42a, Inf50d, Inf53f, IP56c, SB52]. methods [Inf61c]. Michalska [Ano16]. Michalska-Trautman [Ano16]. middle [Inf74b]. middle-school [Inf74b]. Mikołaj [BBI^{+}65]. Mikołaja [Bie73b]. Mind [Inf50b, Goo98]. missing [Bou11]. Model [CP13b, Anoxxa, CP13a]. Modern [Ale10, EI39a, Foc53a, Con05, EI61a, Foc54, Inf34b, SZI53, Sil10, Inf40a]. moderne [Con05, EI39a, EI61a, Foc54]. modernen [IIK^{+}54]. modification [IP57]. modified [IP56d]. Moi [Inf56c]. Moje [Inf56d]. Momentum [Coi56, Inf47f, BR04]. mondo [Inf63b, Inf68a, Inf74a, Inf92]. Monographs [Ber61]. monopole [Lom05]. Monopoly [Sch54]. Montreal [Dav82]. mort [Tit82]. Most [Ban51, Sch54, Inf61i]. Motion [EI49a, EI49b, IW40, Inf40b, IS49b, IS50, Inf50e, Inf50f, IS51, Inf56e, Inf57e, Inf59a, IP60, IP62b, Inf65c, IMT69a, SI50, EIH38a, EIH38b, EI40b, Inf51a, Inf53b, Inf54a, Inf54e, Inf55e, IP56b, IP56f, Inf57c, Inf57j, Inf57d, IP62a, Inf63e, Inf64c, Rob38, IP62a, Ber61]. motions [BY05]. mouvement [Con15]. Mr [Inf28a, Inf30a]. MR1503389 [Rob38]. MR2654369 [Sil10]. MR2882171 [Cur12]. Multi [ANN^{+}89]. Multi-Hamiltonian [ANN^{+}89]. mundo [Inf57a, Inf60a, Inf61a, Inf61d, Inf83]. My [Inf15, Inf50b, Ric97, Inf56c, Inf56d, Inf59b].

nach [Ric97]. nacimiento [Bie73a]. naissance [Bie73b]. najdawniejszych [EI59, EI62a, EI98b]. Nakładem [Szc33]. nas [Inf79a]. Nasar [Goo98]. Natanson [Inf78f]. National [GR63]. Naturalist [Nag50]. Nature [Inf61h, Inf61j]. nauce [Inf56a]. nauki [Inf33, Inf57i, Szc33]. Naukowe [Ber61, Ber62]. Nearly [IMT69a]. neben [Inf99]. Nécrologie [Che32a]. Negative [IS45b]. neuen [Inf31b]. Newe [Inf47a]. Newton [EI39a, EI61b, EI39a, EI61b, EI61a, EI62b, EI62c]. Newtona [EI62c]. Newtonian [CC67]. Newtonista [EI62b]. next [Inf99]. nichtholonomen [Inf31c]. Nicolas [Bie73b, Bie73a, BBI^{+}65, Bie73a]. nie [KIR62]. Nine [Whi55]. njegova [Inf79a]. njihov [Inf79a]. no [Inf57a, Inf60a, Inf61a, Inf69a]. nombres [Gal30c]. Non [BdRS01, Ols61b, Inf54a]. Non-Abelian [BdRS01]. non-harmonic [Inf54a]. Non-Political [Ols61b]. nonholonomic [Inf31c]. Nonlinear [BB99, IR89, Spe12, Vaz77]. North [HS89]. Norwegian [EI39c]. Note [Gal30b, IS45b, Rob38, Inf30a, Gal30b]. Noted [Ano55, Ols61a, Whi55]. Notes [Dav82, Gal31a, Ber79a, Gal31a, Inf32c]. nouvelle [BI34d]. novel [WA11]. November [Rox07]. Nowe [Inf33, Inf57i, Szc33]. Nuclear [BEI^{+}55a, Whi55]. nuestro [Inf61d, Inf83]. number [Gal30c]. numeric [Gal30b]. numériques [Gal30b]. NY [DFKI41].

O [SI32, Ano66, EI39b, Inf50h, Inf51c, Inf56d, Inf98]. ob [Inf56c]. Obra [Gal84, Inf57a, Inf60a, Inf61a, Inf61d, Inf83]. obshhej [IP62a]. occasion [Tit82]. Od [Inf55h, EI59, EI62a, EI62c, EI98b]. œuvre [Tit82]. Œuvres [Gal46, Gal89, Pic97, Pic51, PT01]. older [EI38d]. One [Ano42, Inf48a, Inf48b]. only [Inf38b, Inf47f, IS49a, IS50, Inf57g, Pen99]. Operate [Des46]. operational [IP56c]. Operator [CP13b]. operators [Her81, SW79]. opinions [EI38d]. Oppenheimer [Inf78c]. Oppression [Ano42]. oprac [SW50, STW53]. optics [SW50]. optyka [SW50]. opus [Gal11a]. order [Ber62, Bou11, CP13a, Inf57h]. ordinate [Inf61i]. ordre [Bou11]. ore [Inf57l]. original [EI66a]. Osgood [HS89]. Oskarze [Ano66]. Ossetia [Ano66]. ot [EI66a]. Other [Ano42]. Otnositelnosti [Inf55g, EI66a, IP62a]. Our [Sch50a, Ano50a, C.51, Inf46b, Inf49a, Inf50a, Inf53a, Inf56a, Inf56b, Inf57b, Inf58a, Inf61b, Inf61d, Inf79a, Inf83, Nag50]. outline [Pen99]. Ouvrages [Gal31b]. Own [Sch54].

p [Ano50a]. P. [Ber02]. Pacifist [Sal19]. pages [Goo98]. Państwowe [Ber61, Ber62]. Paper [BI34e, Rob38]. Papers [Tan06a, Tan06b, Tan07, Ber62, PT01]. Papiers [PT01, Tan06a, Tan06b, Tan07]. parents [Ric97]. Part [Inf27, Inf28c, Inf49b, ISS49]. particle [BY05, Inf57g, Inf57h]. Particles [EI49b, IS49b]. partie [Inf27, Inf28c]. partir [BI34a]. Partisans [Sal19]. past [Inf64b, Inf66, Ano66]. paths [Inf57i]. pathways [Inf33]. Paul [Nag50]. Peace [Des46, Sal19, Inf54b]. Peaceful [Sal19]. Pebański [Ber61]. pensamiento [EI65a]. pensée [Con05, Con15]. People [Ano42]. Perfectly [ISS49]. Pergamon [Ber61, Ber62]. Peril [Whi55]. Periodic [IMT69a, Gal29]. périodiques [Gal29]. perturbative [Kie11]. pervonachalnykh [EI66a]. Philip [Cai42a, Cai42b]. Philosopher [Sch49a, Sch49b]. Philosopher-Scientist [Sch49a, Sch49b]. philosophy [Cas75]. physica [EI38a]. Physical [Ber61, Foc53a, SBS^{+}61]. Physicist [Ano42, Ano68b, Kac79, Ols61a]. Physicists [Sal19]. Physics [Bič79, Bro39, EI38b, EI38c, EI39c, EI39a, EI42, EI50b, EI50a, EI56, EI57, EI61b, EI62b, EI65a, EI66b, EI68, EI78a, EI84, EI87, EI88, EI89a, EI92, EI95, EI98a, EI04, EI08a, Inf65b, Kae38, RW64, EI38a, EI38d, EI38e, EI39b, EI47, EI48b, EI48a, EI59, EI61c, EI61d, EI61a, EI62a, EI62c, EI63, EI65b, EI66a, EI71, EI78b, EI89b, EI93, EI94, EI98b, EI00, EI08b, EI11, EI13, EI15, Inf27, Inf28c, Inf50d, Inf54d, Inf68b, Cas75, EI38f, EI49c, EI58, Ano39, Arm38, Len39, Qui40]. Physik [EI61b, IIK^{+}54, EI38f, EI49c, EI50a, EI56, EI87, EI88, EI89a, EI92, EI95, EI98a, EI04, N.38]. physikē [EI78a]. physique [EI38e, EI48a, EI63, EI78b, EI89b, EI93, EI15, Inf27, Inf28c]. Picard [Her81]. Piecsetlecie [Bie73b]. PIW [Ano66]. Planck [Cas75, Inf58b, Inf61g, Inf61j, Inf63a, Inf63d]. Plancks [Inf61h]. Plant [Sch54]. Plea [Whi55]. pogladów [EI59, EI62a, EI98b]. Poincaré [Bel37]. Poinsot [Bou11]. point [Wol93]. points [Gal31a]. pojeć [EI59, EI62a, EI98b]. pola [Baż56]. Poland [Ber61, Ber62, AJP86, Ano50b, Ano50g, Bič79, Inf49e, Sch50b]. Pole [Ano68b]. Polish [Ano66, Ber61, Nat32, Ano64, Baż56, BBI^{+}65, EI59, EI62a, EI98b, Inf21a, Inf33, Inf50h, Inf51c, Inf54d, Inf55h, Inf56a, Inf56d, Inf56f, Inf57i, Inf64b, Inf66, Inf68b, Inf98, KI60, KIR62, KIR81, Ols61a, Ols61b, SW50, STW53, SI32]. Political [Ols61b]. politics [Inf78e, Dav82, Bad79, Kac79, dBR78]. Polskiej [Szc33]. poniatii [EI66a]. Pool [Sch54]. Portrait [Inf47d]. portraits [Bre11]. Portuguese [EI39b, EI08a, Inf57a, Inf60a, Inf61a]. possible [SB52]. Post [CC67]. Post-Galilean [CC67]. Post-Newtonian [CC67]. postawy [KI60, KIR62]. Postulate [IS49a, SI49]. Potentials [IP54, IP53]. Power [Des46, Kie11]. powieść [Inf50h, Inf51c, Inf98]. poznani [EI58]. pp [Ber61, Ber62, Kae38, Mal41, Nag48]. preceding [Rob38]. premiers [EI38e, EI48a, EI63, EI78b, EI89b, EI93, EI15]. préparatoire [Ehr08]. Preparatory [Ehr08]. presentation [Tit82]. Presentations [vK63]. Press [Ano50c, Dav82]. primele [EI57]. primeros [EI65a]. primitive [EI39b, Gal32a]. primitives [Gal32a]. primitivos [EI39b]. Principes [BI34d]. Principle [Inf57e, Inf30a, Inf57d]. Principles [BI34d, Inf39, Inf50c, Inf58b]. Problem [Ano42, Inf28d, IS45b, IS56, Inf65c, IMT69b, Tem01, EIH38a, EI40b, Inf30c, Inf32c, Mal79, PR13, Rob38]. problème [Inf32c]. Problems [BTC^{+}41, HI48, IA41b, EIH38b, Inf41a, Inf42a]. procedure [IP56e]. Proceedings [HS89, Baż98, Inf64a]. prof [Baż56]. Prof. [Ano42, Inf58c, Sch50b]. Professeurs [Gal31b]. Professor [Ols61a, A.68]. Professors [Gal31b]. Progress [GK78, Mal41, Gal90]. Projected [Ano50b]. promieniowanie [Inf57i]. przeszłości [Inf64b, Inf66, Ano66]. Przykład [KI60, KIR62]. Ptolemaic [Foc53b]. Ptolemäische [Foc53b]. Public [BEI^{+}55a]. Publications [Cai42b, vK63]. publiés [PT01]. published [PT01]. Pure [GK78, Gal90]. Pyenson [Ber79a, Dav82, Bad79].

Quanta [EI38b, EI38c, EI42, EI66b, EI38e, EI39b, EI47, EI48a, EI61d, EI63, EI65a, EI65b, EI66a, EI71, EI78b, EI89b, EI93, EI94, EI08b, EI11, EI13, EI15, Inf33, Inf34b, Kae38, Bro39]. Quantenmechanik [Inf30b, Inf29]. Quantentheorie [EI61b, Inf30a]. quanti [EI65b, EI13]. quantique [BI34a, BI34d]. Quantisation [BI63b]. Quantization [BI34c, BI35]. Quantum [EI39a, EI62b, Inf29, Inf30b, BI34a, BI34d, EI38a, EI38d, EI57, EI59, EI61b, EI61a, EI62a, EI62c, EI98b, Inf30a, Inf30c, Inf36c, Inf37a, Inf47b]. quantumtheorie [EI38a]. Queen [Dav82]. quelques [Gal31a]. Quest [Inf41b, Inf41c, Inf42b, Inf80, Inf89, Inf06, Ric95, Ric96, Fra81, Roe81]. Quest. [Mal41, Mon41, Ste82]. qui [Gal32a]. quinto [Bie73a].

radiating [Inf64c]. Radiation [Inf38a, Inf38b, ISS49, IS50, IS51, Inf59a, IMT66, Inf67, IMT69a, IMT69b, SI50, HF11, Inf57i, Inf63c, Li16]. radicals [Gal0 , Gal32a]. radicaux [Gal0 , Gal32a]. Radioactivity [Inf47f]. rapports [Tit82]. Raum [Foc54, Inf99]. Raumes [Inf57c]. razvitie [EI66a]. Razvitiya [Inf55g]. Razvoj [EI62c]. Reactor [Sch54]. readers [Edw12]. reading [Inf74b]. reality [EI39c]. Rebuttal [Inf53g]. Recenzja [Szc33]. Rectangular [Inf49b, ISS49]. Recurrence [Inf47e]. Red [Inf61h, Inf61j, Inf63d, Sch54, Inf63a]. Reflections [Bad79, Dav82, Kac79, dBR78, Inf78e]. Reformulation [Dir60]. regained [Ric99]. regular [Kie11, Lom05]. Related [HI48, MP14]. relations [Inf30c, Tit82, Des46]. relative [EI38d]. relatividad [EI65a]. relatividade [EI39b]. relativistes [Inf64a]. Relativistic [Inf45, Inf58b]. relativita [EI65b, EI13, Inf63b, Inf68a, Inf74a, Inf92]. relativitatii [EI57]. Relativitätstheorie [Foc53b, IvdW33, Inf55d, Inf55c]. relativité [EI38e, EI48a, EI63, EI78b, EI89b, EI93, EI15, Inf27, Inf28c]. relativiteitstheorie [EI38a]. relativitets [EI38d]. relativitets- [EI38d]. Relativitetsteori [EI39a, EI61a]. Relativity [Ano62, Ber62, Ber68, Bro39, EI38b, EI38c, EI39a, EI42, EI49a, EI49b, EI66b, Foc55, HS89, Inf21b, Inf21a, IvdW33, Inf34a, IS49b, Inf55d, Inf55g, Inf56e, IP60, Kae38, Nag50, SI49, EI38a, EI38e, EI39b, EI48a, EI57, EI59, EI61a, EI62a, EI63, EI65a, EI65b, EI66a, EI71, EI78b, EI89b, EI93, EI94, EI98b, EI08b, EI11, EI13, EI15, Foc53b, Inf27, Inf28c, Inf38c, Inf43a, Inf43b, Inf46b, Inf54e, Inf55f, IP56e, Inf57d, Inf57k, IP62a, IP62b, Inf63b, Inf63c, Inf63e, Inf68a, Inf74a, Inf92, Pen99, EI47, EI61d, Hir74, Inf55c, Inf57e, Rob38, Ber61, McC63, vK63]. Religia [KI60, KIR81, KIR62]. Religion [KI60, KIR62, KIR81]. reljativizm [IP62a]. Remarkable [Kae38]. Remarks [BI34e, Inf55d, Inf65c, Inf56f]. Remarques [Inf32c]. remember [Gaj99]. Repaired [Ano68a]. Reply [Dir52]. representation [Kis07]. Reprisals [Ano64]. Research [Inf50c]. résolubilité [Gal0 ]. resolution [Gal30a, Gal30b, Gal30a, Gal30b].

Resources [Des46]. Review [Ano39, Ano50a, Arm38, Bad79, Bel48, Ber61, Ber62, Bor48, Bro39, C.51, Cai42a, Dav82, Fra81, Goo98, Inf39, Inf40a, Inf47b, Inf48b, Inf50c, Kac79, Kae38, Len39, Mal41, McC63, Mon41, N.38, Nag48, Pag65, Qui40, dBR78, Roe81, Sch50a, Ste82, Tay79, Tro79, Cur12, Sil10, Inf43a]. Reviewed [CCK50]. Reviews [Ber79a, Des46, Coh80]. revolutionary [Dal56]. révolutionnaire [Dal56]. Richard [Inf39]. Riemann [PR13]. rigid [Inf43b]. Rise [Ale10, EI64, Sil10]. Rising [Ols61b]. Róża [Ano16]. rods [Inf43b]. rola [Inf56a]. Roman [Inf78b, Coh80, WA11]. Romanian [EI57]. Rotating [Hen08]. rozwój [EI59, EI62a, EI98b]. Russell [Ano06]. Russian [EI66a, Inf55g, Inf56c, IP62a, Inf64c, Inf74b, Inf81].

S [McC63, Sal19]. sa [Tit82]. Schild. [Ber79a]. school [Inf74b, Ehr08]. Schrödinger [SB52]. Schuster [Goo98, Kae38]. Science [Bad79, Dav82, Gal31b, Inf47a, Inf53e, Inf62, Kac79, dBR78, Inf33, Inf34b, Inf78e, SZI53, Inf40a]. Sciences [Ber61, Gal31b, SBS^{+}61, Inf56f]. Scientific [Inf50c, Inf61f, Inf57i, IBBT78]. Scientist [Ano50c, Ano50g, Sch49a, Sch49b, Inf41b, Inf41c, Inf63b, Inf68a, Inf74a, Inf89, Inf92, Mal41, Mon41, Ste82]. Scientists [Ano55, BEI^{+}55a, BEI^{+}55b, GR63, Ols61b, Whi55, GR63]. scienziato [Inf63b, Inf68a, Inf74a, Inf92]. screening [Wol91]. Scribner [Ano50a]. séance [Tit82]. Sears [Nag50]. secrets [Sch50b]. Section [ISS49]. See [Inf65a]. Seeks [Ano50g]. Seen [Sta99]. Sein [Inf56b, Inf53a, Inf57b]. Selected [Bor63]. Self [Bre11]. Self-portraits [Bre11]. September [Rox07]. Serbian [Inf79a]. Series [ISC47, Kie11]. ses [Tit82]. shōgai [Inf69a]. shadow [Bre11]. Shift [Inf61h, Inf61j, Inf63d, Inf63a]. shocks [BR04]. si [EI57]. Signed [Whi55]. Signers [Whi55]. Simon [Goo98, Kae38]. Simple [Inf67, IP56f]. Simplicity [Kae38]. Singular [IP56a]. Sir [Rox07]. Sketches [Ano66, Inf64b, Inf66]. skoly [Inf74b]. Slovenian [EI62c]. small [CH03]. Social [Ehr11a]. Sociologist [Ano42]. Solitons [IR89]. soluble [Gal32a]. solubles [Gal32a]. Solution [BTC^{+}41, Kie11, Spe12, Wol93]. Solutions [BTC^{+}41, IA41b, Bre89, CY14, Lom05, Vaz77]. solvability [Gal0 ]. Solved [Liv05]. solving [Inf42a, IP56c]. som [EI39c]. Some [ISC47, Gal31a, Inf41a, Inf55d]. son [Tit82]. Sons [Ano50a]. sont [Gal32a]. Sought [Ano50d]. Source [ISS49]. sources [Kie11]. Soviet [Sch54]. Space [Foc55, IS31, IS45b, Inf47f, Foc54, Inf27, Inf28c, Inf99]. Spałe [Inf54d]. Spanish [Bie73a, EI65a, EI68, EI84, Inf61d, Inf78a, Inf83, Liz16]. Speaks [Inf50b]. special [Inf57k, Hir74]. Species [Nag50]. spherically [Lom05]. spin [Inf57g, Inf57h, BS99]. spinor [Inf32a, IP55b]. Spinorenrechnung [Inf32a]. spinores [Inf32b]. Spinors [Pen99]. Sponer [WHS68]. Sponer-Franck [WHS68]. srednej [Inf74b]. stē [EI78a]. stability [Spe12]. Stadt [Ric97]. Stalin [Inf53c]. Stanisław [Inf58c]. started [EI40a]. State [MH12]. Stationary [Bre89, Kie11]. Statistical [Inf39]. Story [Bel48, Bor48, Inf47b, Inf48c, Inf50h, Inf51c, Inf78d, Inf98, Nag48, Tay79, Tro79, Fer17, Inf53i, Inf54f, Inf55i, Inf57m, Inf78a, Inf78b]. Str [Szc33].

Strange [Ban51, Inf47b]. string [BY05]. Stroll [Gol11]. Structure [IS31, Inf36a, Inf49a, ANN^{+}89, CL07, Inf31a, Inf46b, SI31, SI32]. Struktur [Inf31a]. strukture [SI32]. Studied [Ano50g]. Study [Des46]. sua [Inf57a, Inf60a, Inf61a, Inf63b, Inf68a, Inf74a, Inf92]. Subcommittee [Des46]. Suche [Ric97]. Suite [Tan07]. sul [Inf63b, Inf68a, Inf74a, Inf92]. sulle [Tot93]. Sully [Coh80]. survey [CY14]. sv [Inf58a]. svet [Inf79a]. sviluppo [EI65b, EI13]. Swedish [EI38d]. Sylvia [Goo98]. symmetric [Lom05]. symmetries [CL07]. Symmetry [Liv05, BdRS01]. Synge [Flo08]. System [Foc53b, Inf61i, PR13, Spe12]. Systems [IMT69a]. Szkicach [Ano66]. Szkice [Inf64b, Inf66].

t [Inf58a]. tōn [EI78a]. tabl [Szc33]. Teach [Ano50g]. Teaching [Gal31b]. tekście [Szc33]. tel [IP62a]. Tells [Ano42]. temps [Inf27, Inf28c]. Ten [Nat32]. tenija [Inf74b]. teoretyczna [STW53]. teoria [Baż56, EI57, Inf63b, Inf68a, Inf74a, Inf92]. Teorii [Inf55g, EI59, EI62a, EI66a, EI98b, IP62a]. teorije [EI62c]. term [CP13a]. Test [IS49b]. Texts [Ano55]. Their [Sta99, Inf30c]. theorem [Gal29]. théorème [Gal29]. theoretical [Inf50d, STW53]. Theorie [Foc54, Bou11, Ehr11c, Gal30c, Inf27, Inf28b, Inf28c, Inf32c, Ver34]. Theories [Foc53a, EI38d, EI38e, EI48a, EI63, EI78b, EI89b, EI93, EI15, Inf55e, Lom05, EI38e, EI48a, EI63, EI78b, EI89b, EI93, EI15, Inf64a]. Theory [Arc11, Ban51, Ber68, Bir37, BI34e, BI63a, CC67, EI39a, EI49b, EI62b, Foc55, Gar21, Inf21b, Inf21a, Inf28d, IvdW33, Inf34a, Inf40b, Inf47c, Inf47f, IPS^{+}49, Inf50g, Inf50e, Inf50f, Inf51b, Inf53d, Inf55d, Inf55g, Inf56e, IS56, Inf57e, Inf61h, Inf61j, Inf63d, Inf64a, MH12, Pag65, Baż56, BR04, BI33b, BI34b, BI35, Bou11, CY14, Ehr11c, EI38a, EI57, EI59, EI61b, EI61a, EI62a, EI62c, EI66a, EI98b, Foc53b, Foc54, Gal30c, Gar86, Hen08, Her81, HI37, Inf27, Inf28b, Inf28e, Inf28c, Inf32c, Inf36b, Inf37b, Inf43a, Inf43b, Inf51a, Inf54e, Inf55f, IP56e, Inf57d, Inf57k, IP62a, Inf63a, Inf63b, Inf63c, Inf68a, Inf74a, Inf92, Lie32, Pie98, Ver34, Wol91, Inf55c]. there [Dir52, Inf52]. Thin [Inf49b]. thinking [KI60, KIR62]. thought [Con05, EI65a]. Three [Inf48a, Inf48b, EI38a]. til [EI39a, EI61a]. till [EI38d]. Time [Ehr11a, Foc55, Foc54, Inf27, Inf28c, Inf99]. Tolman [Inf39]. Took [Sch50b]. tot [EI38a]. total [Inf54b]. totale [Inf54b]. Trace [Kae38]. Transcendents [Arc11]. Transformation [CC67]. transformations [Inf37a, IP55a, IP55b]. Translated [Ber79a, Dav82]. Trautman [Ano16, Bič79]. Travaux [Che32b]. treatment [Inf41a]. Tredici [Inf57l]. Tribute [Ber79b]. Trick [Ehr21]. trilogy [SZI53]. trivial [Spe12]. truncated [Kos95a, Kos95b]. tunneling [Li16]. Turkish [EI94, EI11]. twentieth [SZI53]. Two [Ano42, Inf48a, Inf48b, IMT69b, Rob38]. Two-Body [IMT69b]. type [FOP02].

u [Gel68]. u-ve-Hurbana [Gel68]. Udvikling [EI39a, EI61a]. último [Liz16]. Unbestimmtheitsrelationen [Inf30a]. uncertainty [Inf30a, Inf30c].

Unhappy [Nag48]. Unified [Inf28d]. Uniformity [Foc55]. Uniformly [Inf63e]. unitaire [Inf32c]. Unitary [IP55b, Inf32c, Inf36b, Inf37b, IP55a]. Universe [Inf49a, Inf46b]. University [Dav82, DFKI41]. Uniw [Szc33]. unknown [Ano50h, Ano50i, Foc53c]. Unpublished [PT01, Tan06a, Tan06b, Tan07]. Unscientific [Foc53a]. unsere [Inf53a, Inf56b, Inf57b]. uomo [Inf63b, Inf68a, Inf74a, Inf92]. upon [Inf47c]. Uravnenie [Inf64c]. Urge [Whi55]. Urges [Ols61a]. urodzin [Bie73b]. use [Inf53f]. uticaj [Inf79a]. Utilization [Des46]. utveckling [EI38d].

v [IP62a, Inf74b]. V. [Inf53g]. Variational [Inf58b, BR04]. Varsity [Ano50d]. ve [EI94, EI11, Gel68]. velocities [BR04]. verallgemeinerte [Inf32a]. Verdensbillede [EI39a, EI61a]. Vernichtung [Inf54b]. Vessiot [Her81]. via [Li16]. Victory [Ols61a]. vie [Ano17, Ber02, Dup96a, Dup96b, Dup03, Dup92, Fer17, Tit82]. View [EI39a, EI64, Inf53g]. vikas [Inf65b]. Vinci [Inf53e]. Virgil [Des46]. virkelighetens [EI39c]. Virtual [Sch54]. Visit [Inf49c, Inf49d, Inf49e]. vita [Inf57l, Tot93]. vliv [Inf58a]. Vospominaniya [Inf56c]. vzpomínky [Inf59b].

w [Ano66, Inf54d, Inf56a, Szc33]. wa [EI40a]. War [Ano55, BEI^{+}55a, BEI^{+}55b, RE63, Whi55, Sal19]. Warning [Whi55]. Warsaw [Ber61, Ber62, Ano68b, Ols61a]. Warszawa [Inf64a, Szc33]. Was [Whi55, EI40a]. Wataghin [Inf30a, Inf30a]. Wave [IvdW33, IPS^{+}49, Inf49b, ISS49, BR04, BI34a, Inf31a, Inf47e, SI32]. Waves [IS31, IR89, Inf21b, Inf21a, SI31]. według [Baż56, STW53]. Wellengleichung [IvdW33]. Welt [Inf53a, Inf56b, Inf57b]. Weltgeometrie [Inf31b]. Wen [Inf53i, Inf54f, Inf55i, Inf57m]. Werk [Inf53a, Inf56b, Inf57b]. wg [SW50]. which [SB52]. Whittlesey [Nag48]. Whom [Inf48c, Inf78d, Inf53i, Inf54f, Inf55i, Inf57m, Nag48, Bel48, Bor48, Tay79, Tro79]. Width [Inf47c]. Wimperis [Des46]. Win [Ano42]. without [IP53, IP54, PR13]. Władysław [Inf78f]. Wm [Des46]. wolnomyślicielskiej [KI60, KIR62]. Work [C.51, Nag50, Sch50a, Ano50a, Inf28a, Inf30a, Inf50a, Inf53a, Inf56a, Inf56b, Inf57b, Inf57a, Inf58a, Inf60a, Inf61b, Inf61a, Inf61d, IBBT78, Inf79a, Inf83, Tit82]. Worked [Ano68b]. Works [PT01, SC06, Bor63, Che32b, Gal46, Pic97, Pic51, Gal84]. World [Des46, EI39a, GR63, Inf46a, Inf31b, Inf34b, Inf50a, Inf53a, Inf56a, Inf56b, Inf57b, Inf57a, Inf58a, Inf60a, Inf61b, Inf61a, Inf61d, Inf63b, Inf68a, Inf74a, Inf79a, Inf83, Inf92, Nag50, Ano50a, C.51, Sch50a, Des46]. worldview [EI61a]. wpływie [SI32]. Writers [Ano64]. Writings [Cur12, Neu11, BA62, BA76]. wspomnienia [Inf56d, Inf64b, Inf66, Inf68b]. Wu [Inf36c]. Wybrańcy [Inf50h, Inf51c, Inf98]. Wydawnictwo [Ber61, Ber62]. wykładów [SW50, STW53]. wykładu [Baż56]. wzgledności [EI59, EI62a, EI98b].

x [Kae38, Szc33]. years [GMW80, Nat32, Inf50b]. York [Ano50a, Ber61, Ber62, Kae38, Mal41, Nag48]. yu [Inf36c].

z [Ano66, Inf54d, Inf64b, Inf66, Szc33]. Zeit [Foc54, Inf99]. Zeno [Bel37]. zhi [Inf36c]. zi [Inf36c]. Zum [Inf28d]. zur [EI61b, Inf28e, Inf31c]. zwischen [Inf99].

References

A:1968:PLI

[A.68] M. A. Professor Leopold Infeld. Bulletin on General Relativity and Gravitation, 16(1):N1, 1968.

Abramsky:1986:JP

[AJP86] Chimen Abramsky, Maciej Jachimczyk, and Antony Polonsky, editors. The Jews in Poland. B. Blackwell, Oxford, UK, 1986. ISBN 0-631-14857-4. LCCN DS135.P6 J46 1986.

Alexander:2010:DDH

[Ale10] Amir Alexander. Duel at Dawn: Heroes, Martyrs, and the Rise of Modern Mathematics. Harvard University Press, Cambridge, MA, USA, 2010. ISBN 0-674-04661-7. 307 pp. LCCN QA10.7 .A44 2010.

Arik:1989:MHS

[ANN^{+}89] Metin Arik, Fahrünisa Neyzi, Yavuz Nutku, Peter J. Olver, and John M. Verosky. Multi-Hamiltonian structure of the Born–Infeld equation. Journal of Mathematical Physics, 30(6):1338–1344, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i6/p1338_s1.

Anonymous:1939:BRE

[Ano39] Anonymous. Book review: The Evolution of Physics by Albert Einstein and Leopold Infeld. Revue de métaphysique et de morale, 46 (1):173, 1939. ISSN 0035-1571 (print), 2102-5177 (electronic). URL

http://www.jstor.org/stable/40898826.

Anonymous:1942:TAW

[Ano42] Anonymous. Two authors win [John] Anisfield Awards: Dr. Leopold Infield, physicist, and Prof. J. G. Leyburn, sociologist, honored. books fought oppression: One discourses on the Jewish problem and the other tells of the Haitian people. New York Times, ??

(??):19, February 13, 1942. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/106193980.

Anonymous:1950:BRI

[Ano50a] Anonymous. Book review: Infeld, Leopold. Albert Einstein: His work and its influence on our world. New York: Charles Scribner’s Sons, 1950. 132 p. $2.00. Science Education, 34(5):333, 1950. CO-DEN SEDUAV. ISSN 1098-237X.

Anonymous:1950:CHC

[Ano50b] Anonymous. Canadian hits at critics: Denies projected lectures in Poland have atomic angle. New York Times, ??(??):20, March 19, 1950. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/111479921/.

Anonymous:1950:CSA

[Ano50c] Anonymous. Canadian scientist angered by press. The British Columbian, ??(??):??, March 15, 1950. Reprinted in [Inf78e, page 46].

Anonymous:1950:LVS

[Ano50d] Anonymous. Leave from varsity sought by Infeld. Toronto Telegram, ??(??):??, March 18, 1950. Reprinted in [Inf78e, pages 46–47].

Anonymous:1950:LI

[Ano50e] Anonymous. The loss of Infeld. Varsity, ??(??):??, ???? ??, 1950. Reprinted in [Inf78e, pages 59–60].

Anonymous:1950:MDI

[Ano50f] Anonymous. More on Dr. Infeld. The Ensign, ??(??):??, April 1, 1950. Reprinted in [Inf78e, pages 50–51].

Anonymous:1950:SLS

[Ano50g] Anonymous. Scientist’s leave studied: Einstein ex-associate seeks to teach again in Poland. New York Times, ??(??):10, March 17, 1950. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

URL http://search.proquest.com/hnpnewyorktimes/docview/
111497409.

Anonymous:1950:Ua

[Ano50h] Anonymous. [unknown]. Globe and Mail, ??(??):??, March 16, 1950. Reprinted in [Inf78e, page 48].

Anonymous:1950:Ub

[Ano50i] Anonymous. [unknown]. The Ottawa [Evening] Citizen, ??(??):??, March 18, 1950. Reprinted in [Inf78e, pages 51–52].

Anonymous:1955:TAN

[Ano55] Anonymous. Texts of appeal by noted scientists for abolition of war. New York Times, ??(??):25, July 10, 1955. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http:/

/search.proquest.com/hnpnewyorktimes/docview/113171452/.

Anonymous:1962:RDG

[Ano62] Anonymous, editor. Recent Developments in General Relativity. Pergamon Press, London, UK, 1962. 472 pp. LCCN QC 6 .R295. This book is dedicated to Leopold Infeld in connection with his 60th birthday.

Anonymous:1964:PWD

[Ano64] Anonymous. Polish writers denounce curbs: Demand for more freedom is met with reprisals. New York Times, ??(??):2, April 18, 1964. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

URL http://search.proquest.com/hnpnewyorktimes/docview/
115564167.

Anonymous:1966:KLI

[Ano66] Anonymous. Komunizujacy Leopold Infeld o antykomuniscie Oskarze Haleckim w Szkicach z przeszłości. PIW, 1966. (polish) [Communist Leopold Infeld on anti-communist Ossetia Halecki in Sketches of the past. cip, 1966], 1966. URL https://lustronauki.

wordpress.com/2009/06/13/oskar-halecki/.

Anonymous:1968:IR

[Ano68a] Anonymous. An injustice to be repaired. Globe and Mail, ??(??):??, January 18, 1968. See [Inf78e, page 193].

Anonymous:1968:LIP

[Ano68b] Anonymous. Leopold Infeld, physicist, dead: Pole worked with Einstein – directed Warsaw Institute. New York Times, ??(??): 51, January 17, 1968. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/118493262/.

Anonymous:2006:REM

[Ano06] Anonymous. The Russell–Einstein Manifesto. The SPS Observer, 23

(2):??, Summer 2006. URL http://www.spsobserver.org/2006/

observer_ethics.pdf. The Manifesto is signed by Max Born, Percy W. Bridgman, Albert Einstein, Leopold Infeld, Frederic Joliot-Curie, Herman J. Muller, Linus Pauling, Cecil F. Powell, Joseph Rotblat, Bertrand Russell, and Hideki Yukawa.

Anonymous:2008:LI

[Ano08] Anonymous. Leopold Infeld. In Marquis Who Was Who in America 1607–1984, page ???? Marquis Who’s Who LLC, New Providence, NJ, USA, 2008. ISBN 1-84972-344-3, 0-8379-0269-X, 1-84972-397-4 (e-book). LCCN E176 .W64 2008.

Anonymous:2016:RMT

[Ano16] Anonymous. Róża michalska-trautman (1932–2015). Acta Physica Polonica A, 130(3):665, 2016. CODEN ATPLB6. ISSN 0587-4246 (print), 1898-794X (electronic).

Anonymous:2017:VVG

[Ano17] Anonymous. La vie d’Évariste Galois. (French) [The life of Évariste Galois]. 8-minute lecture video., October 4, 2017. URL https://

www.youtube.com/watch?v=yh3aa5IS_ME.

Anonymous:20xx:GML

[Anoxxa] Anonymous. Geocentric model and Leopold Infeld. Online en-

cyclopedia article, 20xx. URL http://en.unionpedia.org/c/
Geocentric_model/vs/Leopold_Infeld.

Anonymous:20xx:LI

[Anoxxb] Anonymous. Leopold Infeld. Online encyclopedia article, 20xx. URL

http://en.unionpedia.org/i/Leopold_Infeld.

Archibald:2011:DEA

[Arc11] Tom Archibald. Differential equations and algebraic transcendents: French efforts at the creation of a Galois theory of differential equations 1880–1910. Revue d’Histoire des Math’ematiques = Journal for the History of Mathematics, 17(2):373–401, ???? 2011. ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/Publications/RevueHistoireMath/17/pdf/smf_rhm_17_373-
401.pdf.

Armstrong:1938:BRE

[Arm38] E. F. Armstrong. Book review: The Evolution of Physics by Albert Einstein, Leopold Infeld. Journal of the Royal Society of Arts, 86(4477):1045–1046, September 9, 1938. CODEN JRSAA4. ISSN

0035-9114. URL http://www.jstor.org/stable/41361408.

Astruc:1994:EG

[Ast94] Alexandre Astruc. Évariste Galois. Flammarion, Paris, France, 1994. ISBN 2-08-066675-4. 223 pp. LCCN QA29.G25 A88 1994.

Bourgne:1962:EMM

[BA62] Robert Bourgne and Jean-Pierre Azra, editors. Écrits et Mémoires Mathématiques d’Évariste Galois. (French) [Writings and Mathematical Memoirs of Évariste Galois]. Gauthier-Villars et cie, Paris, France, 1962. ???? pp.

Bourgne:1976:EMM

[BA76] Robert Bourgne and Jean-Pierre Azra, editors. Écrits et Mémoires Mathématiques d’Évariste Galois. (French) [Writings and Mathematical Memoirs of Évariste Galois]. Gauthier-Villars et cie, Paris, France, second edition, 1976. ???? pp. Reprinted in 1997 by Éditions Jacques Gabay, Paris.

Badash:1979:BRW

[Bad79] Lawrence Badash. Book review: Why I Left Canada: Reflections on Science and Politics by Leopold Infeld; Helen Infeld; Lewis Pyenson. Isis, 70(4):638, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230782.

Bandyopadhyay:1951:SFE

[Ban51] Gaganbihari Bandyopadhyay. A strange feature of Einstein’s most recent generalized field theory. Nature, 167(4251):648–649, April 21, 1951. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL http://www.nature.com/nature/journal/
v167/n4251/pdf/167648a0.pdf.

Bazanski:1956:KTP

[Baż56] Stanisław Leon Bażański. Klasyczna teoria pola : według wykładu prof. L. Infelda. (Polish) [Classical field theory: according to prof. L. Infeld]. Pa stwowe Wydawnictwo Naukowe, Warszawa, Poland and Łódź, Poland, 1956. 242 pp. LCCN ????

Bazanski:1998:PIC

[Baż98] Stanisław L. Bażański, editor. Proceedings of the Infeld (1898–1968) Centennial Meeting held on June 22–23, 1998, volume 30(10) of Acta Physica Polonica B. University of Warsaw. Institute of Theoretical Physics, Polish Academy of Sciences. Center for Theoretical Physics, Warszawa, Poland, 1998. ISSN 0587-4254 (print), 1509-5770 (electronic). LCCN ???? URL http://www.actaphys.uj.

edu.pl/fulltext?series=Reg&vol=30&page=2853.

Bialynicki-Birula:1999:BIN

[BB99] Iwo Białynicki-Birula. Born–Infeld nonlinear electrodynamics. Acta Physica Polonica B, 30(10):2875–2878, October 1999. CO-DEN APOBBB. ISSN 0587-4254 (print), 1509-5770 (electronic).

URL http://www.actaphys.uj.edu.pl/fulltext?series=Reg&
vol=30&page=2875.

Billig:1965:MKP

[BBI^{+}65] Wilhelm Billig, Aleksander Birkenmajer, Leopold Infeld, Stanisław Konopka, Bogusław Leśnodorski, Edward Lipiński, Michał Łomonosow, Ludwik H. Morstin, and Włodzimierz Zonn. Mikołaj Kopernik. (Polish) [Nicolas Copernicus]. Państwowe wydawnictwo naukowe, Warszawa, Poland, 1965. 304 pp. LCCN ???? Edited by Józef Hurwic.

Bergshoeff:2001:NAB

[BdRS01] E. A. Bergshoeff, M. de Roo, and A. Sevrin. Non-Abelian Born– Infeld and kappa-symmetry. Journal of Mathematical Physics, 42(7): 2872–2888, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bridgman:1955:NWA

[BEI^{+}55a] P. W. Bridgman, Albert Einstein, Leopold Infeld, H. J. Muller, C. F. Powell, Joseph Rotblat, Bertrand Russell, and Hideki Yukawa. Nuclear war — an appeal to scientists and the public. The Lancet (London, England), 2(266)(6881):131–132, July 16, 1955. CODEN LANCAO. ISSN 0140-6736 (print), 1474-547x (electronic). This is the famous Einstein–Russell Manifesto on nuclear disarmament.

Bridgman:1955:SAA

[BEI^{+}55b] Percy W. Bridgman, Albert Einstein, Leopold Infeld, Hermann Joseph Muller, Cecil F. Powell, Bertrand Russell, and Hideki Yukawa. Scientists appeal for abolition of war. Bulletin of the Atomic Scientists, 11(7):236–237, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprint in [RE63].

Bell:1937:MML

[Bel37] Eric Temple Bell. Men of mathematics: The Lives and Achievements of the Great Mathematicians from Zeno to Poincaré. Simon and Schuster, New York, NY, USA, 1937. ISBN 0-671-62818-6. xxi + 592 pp. LCCN QA28 .B4. US$17.00.

Bell:1948:BRG

[Bel48] E. T. Bell. Book review: Genius: Whom the Gods Love. The Story of Évariste Galois by Leopold Infeld. The Scientific Monthly, 66(5): 431–432, 1948. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-

7513 (electronic). URL http://www.jstor.org/stable/19499.

Bertrand:1902:VEG

[Ber02] Joseph Bertrand. La vie d’Évariste Galois par P. Dupuy. (French) [The life of Évariste Galois by P. Dupuy], chapter 15, pages 329– 345. Hachette et cie, Paris, France, 1902. LCCN Q141 B4 1902. New series. With a historical elegy of Joseph Bertrand by Gaston Darboux.

Bergmann:1961:BRM

[Ber61] Peter G. Bergmann. Book review: Motion and Relativity (Polish Academy of Sciences, Physical Monographs). Leopold Infeld and Jerzy Pebański. Pergamon, New York; Państwowe Wydawnictwo Naukowe, Warsaw, Poland, 1961. 229 pp. Illus. $7.50. Science, 133 (3460):1250, April 21, 1961. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Bergmann:1962:BRR

[Ber62] P. Bergmann. Book review: Recent Developments in General Relativity. A collection of papers dedicated to Leopold Infeld. Państwowe Wydawnictwo Naukowe, Warsaw, Poland; Pergamon, New York, 1962 (order from Macmillan, New York). 472 pp. $8. Science, 138 (3537):134, October 12, 1962. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Bergmann:1968:LIA

[Ber68] Peter G. Bergmann. Leopold Infeld, authority on field theory and relativity. Physics Today, 21(3):113, March 1968. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Bergmann:1979:BRW

[Ber79a] Peter G. Bergmann. Book reviews: Why I left Canada, by Leopold Infeld. Translated by Helen Infeld. Edited with introduction and notes by Lewis Pyenson. Foreword by Alfred Schild. Bulletin of the Atomic Scientists, 35(3):60–61, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bergmann:1979:TI

[Ber79b] Peter G. Bergmann. Tribute to Infeld. Physics Today, 32(1):95, January 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Born:1933:LEM

[BI33a] Max Born and Leopold Infeld. Letter: Electromagnetic mass. Nature, 132(3347):970, December 23, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1933:LFN

[BI33b] Max Born and Leopold Infeld. Letter: Foundations of the new field theory. Nature, 132(3348):1004, December 30, 1933. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1934:DEO

[BI34a] Max Born and Leopold Infeld. Deduction de l’équation d’ondes de Dirac à partir de l’electrodynamique quantique. (French) [Deduction of the Dirac wave equation from quantum electrodynamics]. Comptes Rendus Hebdomadaires des Séances de l’Academie des Sciences, 199:1596–1598, November–December 1934. ISSN 0001-4036.

Born:1934:FNF

[BI34b] Max Born and Leopold Infeld. Foundations of the new field theory. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 144(A852):425– 451, March 29, 1934. ISSN 0950-1207 (print), 2053-9150 (electronic).

Born:1934:QNF

[BI34c] Max Born and Leopold Infeld. On the quantization of the new field equations. I. Proceedings of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 147(862):522–546, December 1, 1934. CODEN PRLAAZ. ISSN 1364-5021 (print), 1471-2946 (electronic).

Born:1934:PNE

[BI34d] Max Born and Leopold Infeld. Principes de la nouvelle electrodynamique quantique. (French) [principles of the new quantum electrodynamics]. Comptes rendus de l’Académie des sciences, Paris, 199:1297–1299, November–December 1934. ISSN 0001-4036.

Born:1934:RPF

[BI34e] Max Born and Leopold Infeld. Remarks on the paper by Frenkel on Born’s theory of the electron. Proceedings of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 146(859):935, October 15, 1934. CODEN PRLAAZ. ISSN 1364-5021 (print), 1471-2946 (electronic).

Born:1935:QNF

[BI35] Max Born and Leopold Infeld. On the quantization of the new field theory — II. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 150(A869):141–166, May 1, 1935. CODEN PRLAAZ. ISSN 0080-4630.

Born:1963:FNF

[BI63a] Max Born and Leopold Infeld. Foundations of the new field theory. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63], chapter 74, pages 514–540. LCCN ???? Reprinted from Proc. Roy. Soc. A 144, 425–451 (1934).

Born:1963:QNF

[BI63b] Max Born and Leopold Infeld. On the quantisation of the new field equations I. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63], chapter 75, pages 541–559. LCCN ???? Reprinted from Proc. Roy. Soc. A 147, 522–546 (1934).

Born:1967:EEG

[BI67] Max Born and Leopold Infeld. Erinnerungen an Einstein. (German) [Memories of Einstein]. Union-Verlag, Berlin, West Germany, 1967. ?? pp. LCCN ????

Born:1969:EEG

[BI69] Max Born and Leopold Infeld. Erinnerungen an Einstein. (German) [Memories of Einstein]. Union-Verlag, Berlin, West Germany, 1969. 74 pp. LCCN ????

Bicak:1979:TAE

[Bič79] Jiří Bičák. With Trautman on Albert Einstein, Leopold Infeld and on physics in Poland. Ceskoslovensky Casopis pro Fyziku Sekce A, 29(3):255–259, ???? 1979. CODEN CKCFAH. ISSN 0009-0700 (print), 1804-8536 (electronic).

Bienkowska:1973:NCQ

[Bie73a] Barbara Bieńkowska. Nicolás Copérnico en el quinto centenario de su nacimiento, 1473–1973. (Spanish) [Nicolas Copernico on the fifth centenary of his birth, 1473–1973]. Siglo Veintiuno Editores, Buenos Aires, Argentina, 1973. 189 + 3 pp. LCCN ???? Edited by Barbara Bieńkowska, with works by Stanisław Herbst and others. Translation from Polish by Estanislao J. Zembrzuski of Mikołaj Kopernik w piecsetlecie urodzin, 1473–1973.

Bienkowska:1973:PUM

[Bie73b] Barbara Bieńkowska. Piecsetlecie urodzin Mikołaja Kopernika / 500^{e} anniversaire de la naissance de Nicolas Copernic [1473–1973]. Towarzystwo Naukowe w Toruniu / Société Scientifique de Toruń, Toruń, Poland, 1973. ???? pp. LCCN ???? URL http://kpbc.

umk.pl/dlibra/docmetadata?id=44423&from=pubindex&dirids=
103&lp=2.

Birkhoff:1937:GGT

[Bir37] Garrett Birkhoff. Galois and group theory. Osiris, 3(??):260–268, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301590.

Bondoni:2013:GFM

[Bon13] Davide Bondoni. Galois’s first memoir. Notices of the American Mathematical Society, 60(5):598–599, May 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Borges:1948:BRW

[Bor48] Frederic W. Borges. Book review: Whom the Gods Love: The Story of Évariste Galois by Leopold Infeld. The Mathematics Teacher, 41 (5):243, 1948. ISSN 0025-5769 (print), 2330-0582 (electronic). URL

http://www.jstor.org/stable/27953329.

Born:1963:AAGb

[Bor63] Max Born. Ausgewählte Abhandlungen. (German) [Selected works], volume 2. Vandenhoeck & Ruprecht, Göttingen, West Germany, 1963. viii + 706 pp. LCCN ???? Mit einem Verzeichnis der wissenschaftlichen Schriften (German) [With a table of contents of scientific writings].

Boucard:2011:LPT

[Bou11] Jenny Boucard. Louis Poinsot et la théorie de l’ordre: un chaînon manquant entre Gauss et Galois?. (French) [Louis Poinsot and the theory of order: a missing link between Gauss and Galois?]. Revue d’Histoire des Math’ematiques = Journal for the History of Mathematics, 17(1):41–138, 2011. ISSN 1262-022X (print), 1777-

568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/17/pdf/smf_rhm_17_41-138.pdf.

Boillat:2004:EMW

[BR04] Guy Boillat and Tommaso Ruggeri. Energy momentum, wave velocities and characteristic shocks in Euler’s variational equations with application to the Born–Infeld theory. Journal of Mathematical Physics, 45(9):3468–3478, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i9/p3468_s1.

BretonB:1989:SAF

[Bre89] Nora Bretón B. Stationary axisymmetric fluid solutions of the Einstein–Born–Infeld equations. Journal of Mathematical Physics, 30(11):2607–2610, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i11/p2607_s1.

Brechenmacher:2011:SPE

[Bre11] Frédéric Brechenmacher. Self-portraits with Évariste Galois (and the shadow of Camille Jordan). Revue d’Histoire des Math’ematiques = Journal for the History of Mathematics, 17(2):273–371, ???? 2011. ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/17/pdf/smf_
rhm_17_273-371.pdf.

Brown:1939:BRE

[Bro39] G. Burniston Brown. Book review: The Evolution of Physics: The Growth of Ideas from the Early Concepts to Relativity and Quanta, by Albert Einstein and Leopold Infeld. Philosophy, 14(54):242, April 1939. ISSN 0031-8191 (print), 1469-817X (electronic). URL http:/

/www.jstor.org/stable/3746034.

Boillat:1999:BIE

[BS99] Guy Boillat and Alberto Strumia. On the Born–Infeld electron: Spin effects. Journal of Mathematical Physics, 40(1):1–11, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buker:1941:PSEa

[BTC^{+}41] W. E. Buker, V. Thébault, N. A. Court, C. A. Richmond, V. W. Graham, and Leopold Infeld. Problems and solutions: Elementary problems: Problems for solution: E451–E456. American Mathematical Monthly, 48(1):65, January 1941. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.

org/stable/2303183.

Brenier:2005:DPS

[BY05] Yann Brenier and Wen-An Yong. Derivation of particle, string, and membrane motions from the Born–Infeld electromagnetism. Journal of Mathematical Physics, 46(6):062305, June 2005. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/
i6/p062305_s1.

C:1951:BRA

[C.51] C. W. C. Book review: Albert Einstein. His Work and its Influence on our World by Leopold Infeld. The Journal of Philosophy, 48(7): 222, 1951. ISSN 0022-362X (print), 1939-8549 (electronic). URL

http://www.jstor.org/stable/2020458.

Cairns:1942:BRG

[Cai42a] S. S. Cairns. Book review: Galois Lectures, by Jesse Douglas, Philip Franklin, Cassius Jackson Keyser, Leopold Infeld. American Mathematical Monthly, 49(1):52, 1942. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2303783.

Cairns:1942:RPG

[Cai42b] S. S. Cairns. Recent publications: Galois Lectures, by Jesse Douglas, Philip Franklin, Cassius Jackson Keyser, and Leopold Infeld. American Mathematical Monthly, 49(1):52, January 1942. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Casertano:1975:FFE

[Cas75] Giovanni Casertano. Fisica e filosofia: Einstein, Infeld, de Broglie, Heisenberg, Planck. (Italian) [Physics and philosophy: Einstein, Infeld, de Broglie, Heisenberg, Planck]. Edizione Il Tripode, Napoli, Italia, 1975. 151 pp. LCCN QC6 .F55.

Chandrasekhar:1967:PGT

[CC67] S. Chandrasekhar and G. Contopoulos. On a post-Galilean transformation appropriate to the post-Newtonian theory of Einstein, Infeld and Hoffmann. Proceedings of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 298(1453):123–141, April 4, 1967. CODEN PRLAAZ. ISSN 1364-5021 (print), 1471-2946 (elec-

tronic). URL http://www.jstor.org/stable/2415974.

Carlson:1950:BR

[CCK50] A. J. Carlson, Edward A. Chapin, and Emanuel Krimsky. Briefly reviewed. The Scientific Monthly, 71(2):135–136, 1950. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL

http://www.jstor.org/stable/20140.

Chae:2003:GES

[CH03] Dongho Chae and Hyungjin Huh. Global existence for small initial data in the Born–Infeld equations. Journal of Mathematical Physics, 44(12):6132–6139, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chevalier:1832:N

[Che32a] Auguste Chevalier. Nécrologie. In Revue encyclopédique, ou, Analyse raisonnée des productions les plus remarquables dans la politique, les sciences, l’industrie et les beaux-arts, pages 744–752. Bureau de la Revue Encyclopédique, Paris, France, 1832. First published essay on the life of Évariste Galois, including letters that Galois wrote the evening before his fatal duel.

Chevalier:1832:TME

[Che32b] Auguste Chevalier. Travaux mathématiques d’Évariste Galois. (French) [mathematical works of Évariste Galois]. Revue encyclopédique, 55(??):566–576, September 1832.

Cirilo-Lombardo:2007:MSH

[CL07] Diego Julio Cirilo-Lombardo. On the mathematical structure and hidden symmetries of the Born–Infeld field equations. Journal of Mathematical Physics, 48(3):032301, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i3/
p032301_s1.

Cohen:1980:AOR

[Coh80] E. Cohen. Analyses d’ouvrages: Le roman d’Évariste Galois par Leopold Infeld; J. Sully. (French) [Book reviews: Le roman d’Évariste Galois par Leopold Infeld and J. Sully]. Revue d’Histoire des Sciences, 33(3):276–277, juillet 1980. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.

org/stable/23632253.

Coish:1956:IFA

[Coi56] H. R. Coish. Infeld factorization and angular momentum. Canadian Journal of Physics = Journal canadien de physique, 34(4):343–349, 1956. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (electronic).

Connes:2005:PEG

[Con05] Alain Connes. La pensée d’Evariste Galois et le formalisme moderne. (French) [The thought of Évariste Galois and modern formalism]. 61minute lecture video., June 8, 2005. URL https://www.youtube.

com/watch?v=kaG3yhrh9Pw.

Connes:2015:MPM

[Con15] Alain Connes. Les mathématiques et la pensée en mouvement: conférence CPES. Web video (1h55m), November 12, 2015. URL

https://www.youtube.com/watch?v=qrpp1Mh8EDo.

Chen:2013:EMB

[CP13a] Jun Chen and Xing-Bin Pan. An extended magnetostatic Born– Infeld model with a concave lower order term. Journal of Mathematical Physics, 54(11):111501, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2013:FOC

[CP13b] Jun Chen and Xing-Bin Pan. Functionals with operator curl in an extended magnetostatic Born–Infeld model. SIAM Journal on Mathematical Analysis, 45(4):2253–2284, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Curtis:2012:MWE

[Cur12] Charles W. Curtis. The Mathematical Writings of Évariste Galois [book review MR2882171]. Notices of the American Mathematical Society, 59(11):1565–1568, December 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). See [Neu11].

Chern:2014:SSG

[CY14] Jann-Long Chern and Sze-Guang Yang. A survey of solutions in a gravitational Born–Infeld theory. Journal of Mathematical Physics, 55(3):031501, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dalmas:1956:EGR

[Dal56] André Dalmas. Évariste Galois, révolutionnaire et géomètre. (French) [Évariste Galois, revolutionary and geometer]. Fasquelle, Paris, France, 1956. 175 pp. LCCN QA29.G25.

Davis:1981:BRL

[Dav82] Chandler Davis. Book review: Leopold Infeld, Why I Left Canada: Reflections on Science and Politics, Translated by Helen Infeld, Edited with Introduction and Notes by Lewis Pyenson (Montreal: McGill-Queen’s University Press 1978). Labour / Le Travail, 8/9: 374–376, Autumn/Spring 1981–1982. ISSN 0700-3862 (print), 1911-

4842 (electronic). URL http://www.jstor.org/stable/25140104.

Robinson:1971:LI

[dBR71] G. de B. Robinson. Leopold Infeld. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 14(??):301–304, ???? 1971. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Robinson:1978:RWL

[dBR78] G. de B. Robinson. Review of Why I Left Canada: Reflections on Science and Politics. Canadian Journal of Higher Education = La Revue canadienne d’enseignement supérieur, 8(3):94–96, 1978. ISSN 0316-1218 (print), 2293-6602 (electronic). URL http://journals.

sfu.ca/cjhe/index.php/cjhe/article/view/182786/182773.

Desch:1946:BRW

[Des46] C. H. Desch. Book reviews: World Power and Atomic Energy: The Impact on International Relations, by H. E. Wimperis. Utilization and Control of Atomic Energy by The Legal Subcommittee of the Committee on Atomic Energy of the Carnegie Endowment for International Peace. Manifesto for the Atomic Age, by Virgil Jordan. Atomic Energy and World Government by Leopold Infeld. Co-Operate or Die. A Study on World Resources and the Discovery of the Atomic Bomb. by Wm. Agar. International Affairs (Royal Institute of International Affairs 1944–), 22(4):547–548, 1946. ISSN 0020-5850 (print), 1468-2346 (electronic). URL http://www.jstor.

org/stable/3018211.

Douglas:1941:GLA

[DFKI41] Jesse Douglas, Philip Franklin, Cassius Jackson Keyser, and Leopold Infeld. Galois lectures: addresses delivered at the Galois Institute of Mathematics, Long Island University, Brooklyn, NY, volume 5 of The Scripta Mathematica library. Scripta Mathematica, Yeshiva College, New York, NY, USA, 1941. 3–124 pp.

Dirac:1952:TARb

[Dir52] P. A. M. Dirac. Is there an æther?: Reply to Leopold Infeld. Nature, 169(4304):702, April 26, 1952. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v169/n4304/abs/169702b0.html.

Dirac:1960:RBI

[Dir60] P. A. M. Dirac. A reformulation of the Born–Infeld electrodynamics. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 257(1288):32–43, August 23, 1960. CODEN

PRLAAZ. ISSN 0080-4630. URL http://www.jstor.org/stable/
2413797.

Dupuy:1896:VEGa

[Dup96a] Paul Dupuy. La vie d’Évariste Galois. (French) [The life of Évariste Galois]. Annales de l’École Normale, 13:197–266, 1896.

Dupuy:1896:VEGb

[Dup96b] Paul Dupuy. La vie d’Évariste Galois. (French) [The life of Évariste Galois]. Gauthier-Villars et cie, Paris, France, 1896. 70 pp. LCCN ????

Dupuy:1903:VDG

[Dup03] Paul Dupuy. La vie d’Évariste Galois. (French) [The life of Évariste Galois]. Cahiers de la quinzaine, Paris, France, 1903. 100 pp.

LCCN QA29.G25. URL http://catalogue.bnf.fr/ark:/12148/
cb354317921.

Dupuy:1992:VEGe

[Dup92] Paul Dupuy. La vie d’Évariste Galois. (French) [The life of Évariste Galois]. Éditions Jacques Gabay, Paris, France, 1992. ISBN 2-87647-143-4. 100 pp. LCCN ????

Edwards:2012:GCR

[Edw12] Harold M. Edwards. Galois for 21st-century readers. Notices of the American Mathematical Society, 59(7):912–923, August 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ehrhardt:2008:EGC

[Ehr08] Caroline Ehrhardt. Évariste Galois, un candidat à l’École préparatoire en 1829. (French) [Évariste Galois, a candidate for the Preparatory School in 1829]. Revue d’Histoire des Math’ematiques = Journal for the History of Mathematics, 14(2):289–328 (2009), 2008. ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/14/pdf/smf_
rhm_14_289-328.pdf.

Ehrhardt:2011:EGS

[Ehr11a] Caroline Ehrhardt. Évariste Galois and the social time of mathematics. Revue d’Histoire des Math’ematiques = Journal for the History of Mathematics, 17(2):175–210, ???? 2011. ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/Publications/RevueHistoireMath/17/pdf/smf_rhm_17_175-
210.pdf.

Ehrhardt:2011:EGF

[Ehr11b] Caroline Ehrhardt. Évariste Galois: la fabrication d’une icône mathématique. (French) [Évariste Galois: the fabrication of a mathematical icon]. Éditions de l’École des hautes études en sciences sociales, Paris, France, 2011. ISBN 2-7132-2317-2. 300 pp. LCCN QA29.G25 E37 2011.

Ehrhardt:2011:CTG

[Ehr11c] Caroline Ehrhardt. La constitution de la “théorie de Galois” jusqu’à Artin. (French) [the constitution of the ’Galois’ theory up to Artin]. 57-minute lecture video., October 25, 2011. URL https://www.

youtube.com/watch?v=xbcWOqTGego.

Ehrenborg:2021:FGH

[Ehr21] Richard Ehrenborg. Fano, Galois, Hamming and a card trick. College Mathematics Journal, 52(4):274–280, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/07468342.2021.1942678.

Einstein:1938:DEP

[EI38a] Albert Einstein and Leopold Infeld. Drie eeuwen physica: van Galilei tot relativiteitstheorie en quantumtheorie. (Dutch) [Three centuries of physics: from Galileo to relativity theory and quantum theory]. D. B. Centen, Amsterdam, The Netherlands, 1938. 319 pp. LCCN ???? Translation to Dutch of [EI38b].

Einstein:1938:EPGa

[EI38b] Albert Einstein and Leopold Infeld. The Evolution of Physics: the Growth of Ideas from Early Concepts to Relativity and Quanta. Simon and Schuster, New York, NY, USA, 1938. x + 319 + 1 pp. LCCN QC7 .E5.

Einstein:1938:EPGb

[EI38c] Albert Einstein and Leopold Infeld. The Evolution of Physics. The Growth of Ideas from the Early Concepts to Relativity and Quanta. Cambridge library of modern science. Cambridge University Press, Cambridge, UK, 1938. 319 pp. LCCN QC 7 EIN.

Einstein:1938:FUI

[EI38d] Albert Einstein and Leopold Infeld. Fysikens utveckling : idéutveckling från äldre åsikter till relativitets- och kvantteorierna. (Swedish) [The development of physics: idea development from older opinions to the relative and quantum theories]. Natur och Kultur, Stockholm, Sweden, 1938. 253 pp. LCCN ????

Einstein:1938:LIP

[EI38e] Albert Einstein and Leopold Infeld. L’Évolution des idées en physique des premiers concepts aux théories de la relativité et des quanta. (French) [The evolution of ideas in physics from the first concepts to the theories of relativity and quanta]. Flammarion, Paris, France, 1938. vi + 298 pp. LCCN ???? Translation to French by Maurice Solovine of [EI38b].

Einstein:1938:PAE

[EI38f] Albert Einstein and Leopold Infeld. Physik als Abenteuer der Erkenntnis. (German) [Physics as an adventure of knowledge]. A. W. Sijthoff’s Uitgeversmaatschappij N.V., Leiden, The Netherlands, 1938. viii + 222 pp. LCCN ???? Translation to German of [EI38b].

Einstein:1939:MVF

[EI39a] Albert Einstein and Leopold Infeld. Det moderne Verdensbillede: Fysikkens Udvikling fra Galilei og Newton til Relativitetsteori og Kvantemekanik. (Danish) [The Modern World View: The Development of Physics from Galileo and Newton to Relativity Theory and Quantum Mechanics]. Schultz, Copenhagen, Denmark, 1939. 287 pp. LCCN ???? Translation to Danish by Niels Arley of [EI38b].

Einstein:1939:EFD

[EI39b] Albert Einstein and Leopold Infeld. A evolução da fisica: o desenvolvimento das ideias desde os primitivos conceitos ate a relatividade e os quanta. (Portuguese) [The evolution of physics: the development of ideas from primitive concepts to relativity and quanta], volume 1 of Biblioteca do espirito moderno. Serie 2a, Ciencia. Companhia Editora Nacional, São Paulo, Brazil, 1939. 344 + 3 pp. LCCN ???? Translation to Japanese of [EI38b].

Einstein:1939:FSV

[EI39c] Albert Einstein and Leopold Infeld. Fysikken som virkelighetens eventyr. (Norwegian) [Physics as reality’s adventure]. Aschehoug, Oslo, Norway, 1939. 310 pp. LCCN ???? Translation to Norwegian by Gunnar Randers of [EI38b].

Einstein:1939:BWI

[EI40a] Albert Einstein and Leopold Infeld. Butsurigaku wa ikani hajimerareta ka. (Japanese) [How was it started?]. Iwanami Shoten, Tokyo, Japan, 1939–1940. ???? (two volumes) pp. LCCN ???? Translation to Japanese of [EI38b].

Einstein:1940:GEP

[EI40b] Albert Einstein and Leopold Infeld. The gravitational equations and the problem of motion. II. Annals of Mathematics (2), 41:455– 464, 1940. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic).

Einstein:1942:EPG

[EI42] Albert Einstein and Leopold Infeld. The Evolution of Physics. The Growth of Ideas from the Early Concepts to Relativity and Quanta. Simon and Schuster, New York, NY, USA, 1942. x + 319 + 1 pp. LCCN ????

Einstein:1947:EPE

[EI47] Albert Einstein and Leopold Infeld. The evolution of physics from early concepts to Relativity and quanta. A Touchstone book. Cambridge University Press, Cambridge, UK, 1947. x + 319 pp. LCCN QC7 .E5 1947.

Einstein:1948:LIP

[EI48a] Albert Einstein and Leopold Infeld. L’Évolution des idées en physique des premiers concepts aux théories de la relativité et des quanta. (French) [The evolution of ideas in physics from the first concepts to the theories of relativity and quanta]. Flammarion, Paris, France, eleventh edition, 1948. 298 pp. LCCN ???? Translation to French by Maurice Solovine of [EI38b].

Einstein:1948:LDF

[EI48b] Albert Einstein and Leopold Infeld. L’evoluzione della fisica (Italian). [The evolution of physics]. G. Einaudi, Torino, Italy, 1948. 310 pp. LCCN QC7 .E515.

Einstein:1949:EMG

[EI49a] Albert Einstein and Leopold Infeld. [Equations of motion in general relativity]. Cited in [Inf78e, page 149] as the authors’ last joint paper, but never finally completed or published., 1949.

Einstein:1949:MPG

[EI49b] Albert Einstein and Leopold Infeld. On the motion of particles in general relativity theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(3):209–241, 1949. CODEN CJMAAB. ISSN 0008-414X.

Einstein:1949:PAE

[EI49c] Albert Einstein and Leopold Infeld. Physik als Abenteuer der Erkenntnis. (German) [Physics as an adventure of knowledge]. Sijthoff, Leiden, The Netherlands, 1949. viii + 200 pp. LCCN ????

Einstein:1950:EPG

[EI50a] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics]. Zsolnay-Verlag, Wien, Austria, 1950. 352 pp. LCCN ????

Einstein:1950:EP

[EI50b] Albert Einstein and Leopold Infeld. The Evolution of Physics. Scientific Book Club, London, UK, 1950. x + 320 pp. URL https://

archive.org/details/evolutionofphysi033254mbp.

Einstein:1956:EPG

[EI56] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics]. Rowohlt, Reinbek bei Hamburg, Germany, 1956. ???? pp. LCCN ????

Einstein:1957:EFD

[EI57] Albert Einstein and Leopold Infeld. Evoluţia fizicii: desvoltarea ideilor de la primele concepte la teoria relativităţii şi teoria cuantelor. (Romanian) [Evolution of Physics: The development of ideas from the first concepts to the theory of relativity and quantum theory]. Editura Tehnică, Bucureşti, Romania, 1957. 240 pp. LCCN ???? Translation to Romanian of [EI38b].

Einstein:1958:FJD

[EI58] Albert Einstein and Leopold Infeld. Fyzika jako dobrodruzstvi poznani. (Czech) [Physics as a knowledge adventure]. Orbis, Praha, Czechoslovakia, 1958. 282 pp. LCCN ???? Translation to Czech by Jan Rey of [EI49c].

Einstein:1959:EFR

[EI59] Albert Einstein and Leopold Infeld. Ewolucja fizyki: rozwój pogladów od najdawniejszych pojeć do teorii wzgledności i kwantów. (Polish) [Evolution of physics: development of ideas from the earliest concepts to relativity and quantum theory]. Państwowy Wydawnictwo Naukowe, Warszawa, Poland, 1959. 327 pp. LCCN ???? Translation to Polish by Ryszard Gajewski (1930–2014) of [EI38b].

Einstein:1961:MVF

[EI61a] Albert Einstein and Leopold Infeld. Det moderne verdensbillede, fysikkens udvikling fra Galilei og Newton til relativitetsteori og kvantemekanik. (Danish) [The modern worldview: the development of physics from Galileo and Newton to relativity theory and quantum mechanics], volume 7 of Gyldendals uglebøger. Gyldendal, Oslo, Norway, 1961. 226 pp. LCCN ???? Translation to Danish by Niels Arley of [EI38b].

Einstein:1961:EPN

[EI61b] Albert Einstein and Leopold Infeld. Die Evolution der Physik : von Newton bis zur Quantentheorie. (German) [The Evolution of Physics: from Newton to quantum theory]. Rowohlt, Reinbek bei Hamburg, Germany, 1961. 210 pp. LCCN ????

Einstein:1961:EP

[EI61c] Albert Einstein and Leopold Infeld. The evolution of physics. Cambridge University Press, Cambridge, UK, second edition, 1961. ISBN 0-521-08371-0, 0-521-09687-1. 302 pp. LCCN QC 7 EIN.

Einstein:1961:EPG

[EI61d] Albert Einstein and Leopold Infeld. The evolution of physics: the growth of ideas from early concepts to Relativity and quanta. An Essandess paperback. Simon and Schuster, New York, NY, USA, 1961. xvi + 302 pp. LCCN QC7 .E5 1961.

Einstein:1962:EFR

[EI62a] Albert Einstein and Leopold Infeld. Ewolucja fizyki: rozwój pogladów od najdawniejszych pojeć do teorii wzgledności i kwantów. (Polish) [Evolution of physics: development of ideas from the earliest concepts to relativity and quantum theory]. Państwowy Wydawnictwo Naukowe, Warszawa, Poland, 1962. 263 + 1 pp. LCCN ???? Translation to Polish by Ryszard Gajewski (1930–2014) of [EI38b].

Einstein:1962:FKN

[EI62b] Albert Einstein and Leopold Infeld. Fysiikan kehitys : Newtonista kvanttiteoriaan. (Finnish) [Physics: From Newton to Quantum Theory], volume 11 of Jokamiehen korkeakoulu. Kustannusosakeyhtio Otava, Helsinki, Finland, 1962. 256 pp. LCCN ???? Translation to Finnish by Heikki Hirvinen of [EI50a].

Einstein:1962:RFO

[EI62c] Albert Einstein and Leopold Infeld. Razvoj fizike od Newtona do kvantne teorije. (Slovenian) [Development of Newton’s physics to quantum theory]. Mladinska Knjiga, Ljubljana, Yugoslavia, 1962. 223 pp. LCCN ????

Einstein:1963:LIP

[EI63] Albert Einstein and Leopold Infeld. L’Évolution des idées en physique des premiers concepts aux théories de la relativité et des quanta. (French) [The evolution of ideas in physics from the first concepts to the theories of relativity and quanta], volume 47. Payot, Paris, France, 1963. ???? pp. LCCN ???? Reprint of [EI38e].

Einstein:1964:RMV

[EI64] Albert Einstein and Leopold Infeld. The rise of the mechanical view. In Rapport and Wright [RW64], pages 13–47.

Einstein:1965:FAD

[EI65a] Albert Einstein and Leopold Infeld. La fisica, aventura del pensamiento: el desarrollo de las ideas desde los primeros conceptos hasta la relatividad y los cuantos. (Spanish) [Physics, adventure of thought: the development of ideas from the first concepts to relativity and quanta]. Editorial Losada, Buenos Aires, Argentina, 1965. 254 pp. LCCN ???? Translation by Rafael Grinfeld of [EI38b].

Einstein:1965:LDF

[EI65b] Albert Einstein and Leopold Infeld. L’evoluzione della fisica: sviluppo delle idee dai concetti iniziali alla relativita e ai quanti. (Italian) [The evolution of physics: development of ideas from initial concepts to relativity and quanta]. Boringhieri, Torino, Italia, 1965. 310 pp. LCCN ????

Einstein:1966:EFR

[EI66a] Albert Einstein and Leopold Infeld. Evoliutsiia fiziki: razvitie idei ot pervonachalnykh poniatii do teorii otnositel’nosti i kvantov. (Russian) [Evolution of physics: The development of ideas from the original concepts to the theory of relativity and quanta]. Molodaya gvardiya, Moscow, USSR, fourth edition, 1966. 266 + 4 pp. LCCN ???? Translation to Russian by S. G. Suvorova of [EI38b].

Einstein:1966:EPE

[EI66b] Albert Einstein and Leopold Infeld. The Evolution of Physics from Early Concepts to Relativity and Quanta. A Touchstone book. Simon and Schuster, New York, NY, USA, 1966. ISBN 0-671-20156-5. xvi + 302 pp. LCCN QC7 .E5 1961.

Einstein:1968:LFS

[EI68] Albert Einstein and Leopold Infeld. L’evolución de la física. (Spanish) [The Evolution of Physics], volume 18 of Biblioteca bàsica de cultura contemporània. Edicions 62, Barcelona, Spain, 1968. 244 pp. LCCN ????

Einstein:1971:EPG

[EI71] Albert Einstein and Leopold Infeld. The evolution of physics: the growth of ideas from early concepts to relativity and quanta. A Touchstone book. Cambridge University Press, Cambridge, UK, second edition, 1971. ISBN 0-521-09687-1, 0-521-08371-0. xvi + 302 pp. LCCN QC7 .E5 1971.

Einstein:1978:ETI

[EI78a] Albert Einstein and Leopold Infeld. Ē exelixē tōn ideōn stē physikē. (Greek) [The Evolution of Physics], volume 1 of Epistēmologikē bibliothēkē. Ekdoseis Dōdōnē, Athēna, Greece, 1978. 298 pp. LCCN ???? Translation to Greek by Eut. Mpitsakē.

Einstein:1978:LIP

[EI78b] Albert Einstein and Leopold Infeld. L’Évolution des idées en physique des premiers concepts aux théories de la relativité et des quanta. (French) [The evolution of ideas in physics from the first concepts to the theories of relativity and quanta], volume 47 of Petite bibliothèque Payot. Payot, Paris, France, 1978. ISBN 2-228-30473-5, 2-228-30474-3. 280 pp. LCCN ???? Translation to French by Maurice Solovine of [EI38b].

Einstein:1984:LFS

[EI84] Albert Einstein and Leopold Infeld. L’evolució de la física. (Spanish) [The Evolution of Physics], volume 12 of Clàssics del Pensament Modern. Edicions 62, Barcelona, Spain, 1984. ISBN 84-297-2107-X. 236 pp. LCCN ???? Translation to Spanish by Humbert Padellans of [EI38b]. Edited by David Jou. Preface by Antoni Lloret i Orriols.

Einstein:1987:EPG

[EI87] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics], volume 8342 of rororo-Sachbuch. Rowohlt, Reinbek bei Hamburg, Germany, 1987. ISBN 3-499-18342-0. 265 pp. LCCN ???? Translation to German by Werner Preusser.

Einstein:1988:EPG

[EI88] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics]. Rowohlt, Reinbek bei Hamburg, Germany, ninth edition, 1988. ISBN 3-499-18342-0. 265 pp. LCCN ????

Einstein:1989:EPG

[EI89a] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics]. Rowohlt, Reinbek bei Hamburg, Germany, thirteenth edition, 1989. ISBN 3-499-18342-0. 265 pp. LCCN ????

Einstein:1989:LIP

[EI89b] Albert Einstein and Leopold Infeld. L’Évolution des idées en physique des premiers concepts aux théories de la relativité et des quanta. (French) [The evolution of ideas in physics from the first concepts to the theories of relativity and quanta]. France Loisirs, Paris, France, 1989. ISBN 2-7242-4426-5. 445 pp. LCCN ???? Translation to French by Maurice Solovine of [EI38b]. Preface by Igor Bogdanov and Grichka Bogdanov.Klein.

Einstein:1992:EPG

[EI92] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics]. Weltbild-Verlag, Augsburg, Germany, 1992. ISBN 3-89350-161-4. 332 pp. LCCN ????

Einstein:1993:LIP

[EI93] Albert Einstein and Leopold Infeld. L’Évolution des idées en physique des premiers concepts aux théories de la relativité et des quanta. (French) [The evolution of ideas in physics from the first concepts to the theories of relativity and quanta]. Flammarion, Paris, France, 1993. ISBN 2-08-081119-3. 280 pp. LCCN ???? Reprint of [EI38e].

Einstein:1994:FEI

[EI94] Albert Einstein and Leopold Infeld. Fizigin evrimi: ilk kavramlardan iliskinlige ve kuantumlara. (Turkish) [The evolution of physics: from early concepts to relativity and quanta]. Onur Yayinlari, Ankara, Turkey, 1994. ISBN 975-351-006-3. 262 pp. LCCN ???? Translation to Turkish by Öner Ünalan of [EI47].

Einstein:1995:EPG

[EI95] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics], volume 19921 of Rororo. Rowohlt, Reinbek bei Hamburg, Germany, nineteenth edition, 1995. ISBN 3-499-19921-1. 286 pp. LCCN ???? Translation to German by Werner Preusser. With an introduction by Albrecht Fölsing.

Einstein:1998:EPG

[EI98a] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics], volume 19921 of Rororo. Rowohlt, Reinbek bei Hamburg, Germany, nineteenth edition, 1998. ISBN 3-499-19921-1. 286 pp. LCCN ???? Translation to German by Werner Preusser. With an introduction by Albrecht Fölsing.

Einstein:1998:EFR

[EI98b] Albert Einstein and Leopold Infeld. Ewolucja fizyki: rozwój pogladów od najdawniejszych pojeć do teorii wzgledności i kwantów. (Polish) [Evolution of physics: development of ideas from the earliest concepts to relativity and quantum theory]. Klasycy nauki. Prószyński i S-ka, Warszawa, Poland, fourth edition, 1998. ISBN 83-7180-985-9. 257 pp. LCCN QC7 .E5 1998.

Einstein:2000:FEB

[EI00] Albert Einstein and Leopold Infeld. Fisikaren eboluzioa. (Basque) [Evolution of physics]. Gaiak Argitaldaria, Donostia, Spain, 2000. ISBN 84-89772-44-4. 257 pp. LCCN ???? Translation to Basque by Juan Ignazio Abrisketa of [EI47].

Einstein:2004:EPG

[EI04] Albert Einstein and Leopold Infeld. Die Evolution der Physik. (German) [The Evolution of Physics], volume 19921 of Rororo. Rowohlt, Reinbek bei Hamburg, Germany, nineteenth edition, 2004. ISBN 3-499-19921-1. 286 pp. LCCN ????

Einstein:2008:EFP

[EI08a] Albert Einstein and Leopold Infeld. A evolução da física. (Portuguese) [The Evolution of Physics]. Jorge Zahar, Rio de Janeiro, Brazil, 2008. ISBN 85-378-0052-X. 248 pp. LCCN ???? Translation to Portuguese by Giasone Rebuá of [EI47].

Einstein:2008:EPE

[EI08b] Albert Einstein and Leopold Infeld. The evolution of physics: from early concepts to relativity and quanta. Simon and Schuster, New York, NY, USA, 2008. ISBN 1-4165-5945-0 (hardcover). xxx + 302 pp. LCCN QC7 .E5 2008. Reprint of [EI38b]. With a new foreword by Walter Isaacson.

Einstein:2011:FEI

[EI11] Albert Einstein and Leopold Infeld. Fiziğin evrimi: ilk kavramlardan ili kinliğe ve kuantumlara. (Turkish) [The evolution of physics: from early concepts to relativity and quanta]. Evrensel Basım Yayın, İstanbul, Turkey, 2011. ISBN 605-4156-85-3. 262 pp. LCCN ???? Translation to Turkish by Öner Ünalan of [EI47].

Einstein:2013:LDF

[EI13] Albert Einstein and Leopold Infeld. L’evoluzione della fisica: sviluppo delle idee dai concetti iniziali alla relatività e ai quanti. (Italian) [The evolution of physics: development of ideas from initial concepts to relativity and quanta], volume 20 of I grandi pensatori / [Bollati Boringhieri]. Boringhieri, Torino, Italia, second edition, 2013. ISBN 88-339-2258-8. 279 + 2 pp. LCCN ???? Translation to Italian by Adele Graziadei of [EI47]. Preface by Carlo Castagnoli.

Einstein:2015:LIP

[EI15] Albert Einstein and Leopold Infeld. L’Évolution des idées en physique des premiers concepts aux théories de la relativité et des quanta. (French) [The evolution of ideas in physics from the first concepts to the theories of relativity and quanta]. Champs. Flammarion, Paris, France, 2015. ISBN 2-08-137310-6 (paperback). ISSN 0151-8089. 344 pp. LCCN ???? URL http://catalogue.bnf. fr/ark:/12148/cb44450120n. Translation to French by Maurice Solovine of [EI38b]. Preface by Étienne Klein.

Einstein:1938:GEPa

[EIH38a] Albert Einstein, Leopold Infeld, and Banesh Hoffmann. The gravitational equations and the problem of motion. Annals of Mathematics (2), 39(1):65–100, January 1938. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL http://links.

jstor.org/sici?sici=0003-486X(193801)2:39:1<65:TGEATP>2.
0.CO%3B2-P.

Einstein:1938:GEPb

[EIH38b] Albert Einstein, Leopold Infeld, and Banesh Hoffmann. The gravitational equations and the problems of motion. Annals of Mathematics (2), 41(2):455–564, January 1938. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL http://links.

jstor.org/sici?sici=0003-486X(193801)2:39:1<65:TGEATP>2.
0.CO%3B2-P.

Ferrand:2017:ACU

[Fer17] Franck Ferrand. Au cœur de l’histoire: La vie fulgurante du mathématicien Évariste Galois. (French) [At the heart of the story: The meteoric life of the mathematician Évariste Galois]. 44-minute lecture video., December 3, 2017. URL https://www.youtube.com/

watch?v=QzlMaQrbHDI.

Florides:2008:JLS

[Flo08] Petros S. Florides. John Lighton Synge. 23 March 1897–30 March 1995. Biographical Memoirs of Fellows of the Royal Society, 54:401– 424, December 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Fock:1953:AUC

[Foc53a] V. A. Fock. Against unscientific criticism of modern physical theories. Voprosy filosofii, 1(??):168–174, ???? 1953.

Fock:1953:KPS

[Foc53b] V. A. Fock. Das Kopernikanische und das Ptolemäische System im Lichte der allgemeinen Relativitätstheorie. (German) [The Copernican and the Ptolemaic system in the light of the general theory of relativity]. Sowjetwissenschaft, ??(??):805–809, ???? 1953.

Fock:1953:U

[Foc53c] V. A. Fock. [unknown]. Mysl Filosoficzna, ??(??):??, ???? 1953. Reprint of yet-to-be-identified Russian original, cited in [Inf78e, page 80].

Fock:1954:MTR

[Foc54] V. A. Fock. Die moderne Theorie von Raum und Zeit. (German) [The modern theory of space and time]. Sowjetwissenschaft, ??(??): 525–544, ???? 1954.

Fock:1955:CUC

[Foc55] V. A. Fock. The concept of uniformity, covariance and relativity in the theory of space and time. Voprosy filosofii, 4(??):131–135, ???? 1955.

Fortunato:2002:BIT

[FOP02] D. Fortunato, L. Orsina, and L. Pisani. Born–Infeld type equations for electrostatic fields. Journal of Mathematical Physics, 43(11): 5698–5706, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Franklin:1981:BRQ

[Fra81] Allan Franklin. Book review: Quest: An Autobiography, 2nd ed. by Leopold Infeld. American Scientist, 69(3):351, May/June 1981. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

URL http://www.jstor.org/stable/27850522.

Gajewski:1999:LIR

[Gaj99] Ryszard Gajewski. Leopold Infeld, as I remember him. Acta Physica Polonica B, 30(10):3003–3008, October 1999. CODEN APOBBB. ISSN 0587-4254 (print), 1509-5770 (electronic).

Galois:1828:AAD

[Gal29] Évariste Galois. Analyse algébrique. Démonstration d’un théorème sur les fractions continues périodiques. (French) [Algebraic analysis. Demonstration of a theorem on periodic continued fractions]. Annales de Mathématiques Pures et Appliquées. [Annales de Gergonne]., 19:294–301, April 1828–1829. ISSN 1764-7843. URL http:

//www.numdam.org/item?id=AMPA_1828-1829__19__294_0.

Galois:1830:AMR

[Gal30a] Évariste Galois. Analyse d’un mémoire sur la résolution algébrique des équations. (French) [Analysis of a memoir on the algebraic resolution of equations]. Bulletin des Sciences mathématiques de M. Ferussac, 13(??):271–272, April 1830.

Galois:1830:NRE

[Gal30b] Évariste Galois. Note sur la résolution des équations numériques. (French) [Note on the resolution of numeric equations]. Bulletin des Sciences mathématiques de M. Ferussac, 13(??):413–414, June 1830.

Galois:1830:TNF

[Gal30c] Évariste Galois. Sur la théorie des nombres. (French) [On number theory]. Bulletin des Sciences mathématiques de M. Ferussac, 13 (??):428–??, ???? 1830.

Galois:1830:NQP

[Gal31a] Évariste Galois. Notes sur quelques points d’analyse. (French) [Notes on some points of analysis]. Annales de Mathématiques Pures et Appliquées. [Annales de Gergonne]., 21(??):182, December 1830–

Galois:1831:ESP

[Gal31b] Évariste Galois. Sur l’Enseignement des sciences. des professeurs. des ouvrages. (French) [On the teaching of science. Professors. Books]. La Gazette des Écoles, ??(??):??, January 2, 1831.

Galois:1832:EPQ

[Gal32a] Évariste Galois. Des équations primitives qui sont solubles par radicaux. (French) [Primitive equations that are soluble by radicals]. Unfinished paper, 1832.

Galois:1832:LGM

[Gal32b] Évariste Galois. Lettre de Galois à M. Auguste Chevalier. Unpublished., 1832. Reprinted in [Che32a, pages 568–576].

Galois:1846:OMF

[Gal46] Évariste Galois. Œuvres mathématiques. (French) [Mathematical works]. Journal de mathématiques pures et appliquées, 11(??):381– 444, ???? 1846. Eight papers of Galois, published or written from 1828 to 1832, collected, edited, and introduced by by Joseph Liouville after the death of Galois in a duel in 1832.

Galois:1984:OEG

[Gal84] Évariste Galois. Obra d’Évariste Galois. (Catalan) [Works of Évariste Galois], volume 1 of Monografies de la Secció de Ciéncies [Monographs of the Science Section]. Institut d’Estudis Catalans, Barcelona, Spain, 1984. ISBN 84-7283-060-8. 66 pp. With an introduction, selection and commentaries by Antoni Malet.

Galois:1989:OUM

[Gal89] Évariste Galois. Œuvres mathématiques. Éditions Jacques Gabay, Sceaux, France, 1989. ISBN 2-87647-052-7. 78 pp. Publiées en 1846 dans le Journal de Liouville, suivies d’une étude par Sophus Lie, “Influence de Galois sur le développement des mathématiques” (1895). [Published in 1846 in the Journal de Liouville, followed by a study by Sophus Lie, “Influence of Galois on the development of mathematics” (1895)], With a foreword by J. Liouville.

Galois:1990:DPP

[Gal90] Évariste Galois. Discussion on the progress of pure analysis. Mathesis, 6(3):273–280, 1990. ISSN 0185-6200. Mathesis.

Galois:1997:EMM

[Gal97] Évariste Galois. Écrits et mémoires mathématiques. Les Grands Classiques Gauthier-Villars. [Gauthier-Villars Great Classics]. Éditions Jacques Gabay, Paris, 1997. ISBN 2-87647-020-9. 606 pp. Édition critique intégrale des manuscrits et publications. [Integral critical edition of the manuscripts and publications], With a preface by Jean Dieudonné, Edited, with notes and commentary by Robert Bourgne and Jean-Pierre Azra, Reprint of the second (1976) edition.

Galois:2011:EGO

[Gal11a] Évariste Galois. Évariste Galois—opus. Element, Zagreb, Croatia, 2011. ISBN 978-953-197-9-2-3-8. viii + 87 pp. Translated from the 1989 French edition, edited and with a prologue by Leon Horvat.

Galois:2011:M

[Gal11b] Évariste Galois.  Manuscripts.  Online digitization by F.

Xavier Labrador., 2011.  URL http://www.bibliotheque-
institutdefrance.fr/numerisation/.

Galois:1830:MCR

[Gal0 ] Évariste Galois. Mémoire sur les conditions de résolubilité des equations par radicaux. (French) [Memoir on the conditions of solvability of equations by radicals]. Paper submitted to the Academy of Sciences, but rejected by Poisson., 1830 (??).

Garling:1986:GT

[Gar86] David J. H. Garling. Galois theory. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-31249-3, 0-521-32077-1. v + 167 pp. LCCN QA171 .G277 1986.

Garling:2021:GTA

[Gar21] David J. H. Garling. Galois Theory, and its Algebraic Background. Cambridge University Press, Cambridge, UK, second edition, 2021. ISBN 1-108-83892-8, 1-108-96908-9. LCCN ????

Gelbar:1968:KKF

[Gel68] M. Gelbar. Kehilat Konin be-Frihata u-ve-Hurbana. (Hebrew) [The Community of Konin: Its Flowering and Destruction]. Association of Konin Jews in Israel, Tel Aviv, Israel, 1968. ???? pp. Contains one chapter on Leopold Infeld.

Galois:1978:DPP

[GK78] Évariste Galois and Helen M. Kline. Discussion on the progress of pure analysis. American Mathematical Monthly, 85(7):565–567, 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2320865.

Goldsmith:1980:EFH

[GMW80] Maurice Goldsmith, Alan L. (Alan Lindsay) Mackay, and James Woudhuysen, editors. Einstein, the first hundred years. Pergamon, New York, NY, USA, 1980. ISBN 0-08-025019-X. xiii + 200 pp. LCCN QC16.E5 E52 1980.

Goldstein:2011:CHS

[Gol11] Catherine Goldstein. Charles Hermite’s stroll through the Galois fields. Revue d’Histoire des Math’ematiques = Journal for the History of Mathematics, 17(2):211–270, ???? 2011. ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/Publications/RevueHistoireMath/17/pdf/smf_rhm_17_211-
270.pdf.

Goodstein:1998:BRM

[Goo98] David Goodstein. Book review: Mathematics to Madness, and Back: A Beautiful Mind, by Sylvia Nasar, Simon & Schuster. 459 pages. $25. New York Times, ??(??):E9, June 11, 1998. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http:

//search.proquest.com/hnpnewyorktimes/docview/110033307.

Grodzins:1963:AAS

[GR63] Morton Grodzins and Eugene I. Rabinowitch, editors. The Atomic Age: Scientists in National and World Affairs. Articles from the Bulletin of the Atomic Scientists 1945–1962. Basic Books, New York, NY, USA, 1963. xviii + 616 pp. LCCN D842 .B78. With the assistance of Harvey Flaumenhaft and Lois Gradner.

Hendi:2008:RBB

[Hen08] S. H. Hendi. Rotating black branes in Brans–Dicke–Born–Infeld theory. Journal of Mathematical Physics, 49(8):082501, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/
i8/p082501_s1.

Hermann:1981:IHF

[Her81] Robert Hermann. Infeld–Hull factorization, Galois–Picard–Vessiot theory for differential operators. Journal of Mathematical Physics, 22(6):1163–1167, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i6/p1163_s1.

Hajirahimi:2011:BIB

[HF11] M. Hajirahimi and A. H. Fatollahi. Born–Infeld black-body radiation. Europhysics Letters, 96(4):41001, November 2011. CODEN EULEEJ. ISSN 0295-5075 (print), 1286-4854 (electronic).

Hoffmann:1937:CAF

[HI37] Banesh Hoffmann and Leopold Infeld. On the choice of the action function in the new field theory. Physical Review, 51(9):765–773, May 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Hull:1948:FMH

[HI48] T. E. Hull and Leopold Infeld. The factorization method, hydrogen intensities, and related problems. Physical Review, 74(8):905–909, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Hirondel:1974:SRG

[Hir74] D. Hirondel. Special Relativity in General Relativity? Journal of Mathematical Physics, 15(9):1471–1473, September 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/
i9/p1471_s1.

Howard:1989:EHG

[HS89] Don Howard and John J. Stachel, editors. Einstein and the History of General Relativity: Based on the Proceedings of the 1986 Osgood Hill Conference, North Andover, Massachusetts, 8–11 May 1986, volume 1 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1989. ISBN 0-8176-3392-8, 3-7643-3392-8. LCCN QC173.6 .E38 1989. URL http://

www.springer.com/us/book/9780817633929.

Infeld:1941:E

[IA41a] Leopold Infeld and R. K. Allen. E456. American Mathematical Monthly, 48(9):635–636, 1941. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2303880.

Infeld:1941:PSE

[IA41b] Leopold Infeld and R. K. Allen. Problems and solutions: Elementary problems: Solutions: E456. American Mathematical Monthly, 48(9): 635–636, 1941. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Infeld:1978:LIH

[IBBT78] Eryk Infeld, Iwo Białynicki-Birula, and Andrzej Trautman. Leopold Infeld: his life and scientific work. Polish Scientific Publishers, Warszawa, Poland, 1978. 178 + 2 + 2 pp. LCCN ????

Infeld:1951:FM

[IH51] Leopold Infeld and T. E. Hull. The factorization method. Reviews of Modern Physics, 23(1):21–68, 1951. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Infeld:1954:BMP

[IIK^{+}54] Leopold Infeld, R. Ingarden, M. Krzyzański, J. Rayski, W. Rubinowicz, and W. Wrona. Die Bedeutung der modernen Physik für die Entwicklung der Mathematik. In Die Hauptreferate des 8. Polnischen Mathematikerkongresses, Warschau, September 1953, pages 95–109. Deutscher Verlag der Wissenschaften, Berlin, 1954. Lecture at the Mathematical Congress in Warsaw, 1953.

Infeld:1966:MR

[IMT66] Leopold Infeld and Róża Michalska-Trautman. On mechanics of radiation. Annals of Physics, 40(3):374–394, 1966. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035x (electronic).

Infeld:1969:RSN

[IMT69a] Leopold Infeld and Róża Michalska-Trautman. Radiation from systems in nearly periodic motion. Annals of Physics, 55(3):576–586, 1969. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035x (elec-

tronic). URL http://adsabs.harvard.edu/abs/1969AnPhy..55.
.576I.

Infeld:1969:TBP

[IMT69b] Leopold Infeld and Róża Michalska-Trautman. The two-body problem and gravitational radiation. Annals of Physics, 55(3):561–575, 1969. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035x (elec-

tronic). URL http://adsabs.harvard.edu/abs/1969AnPhy..55.
.561I.

Infeld:1921:PLW

[Inf21a] Leopold Infeld. ???? (Polish) [Light waves in the theory of relativity]. Prace Matematyczno-Fizyczne, 32(??):33–67, ???? 1921. This paper is Leopold Infeld’s doctoral thesis.

Infeld:1921:LWT

[Inf21b] Leopold Infeld. Light waves in the Theory of Relativity. Ph.D. thesis, Uniwersytet Jagielloński, Kraków, Poland, 1921. First doctorate in theoretical physics in Poland.

Infeld:1927:MDT

[Inf27] Leopold Infeld. Sur la mesure du temps et la mesure de l’espace dans la physique classique et dans la théorie de la relativité, I partie. (French) [On the measurement of time and the measurement of space in classical physics and in the theory of relativity, Part I]. C.R. Soc. Polon. Phys., 3(??):5–16, ???? 1927. In Polish, with a French summary.

Infeld:1928:BAH

[Inf28a] Leopold Infeld. Bemerkungen zu der Arbeit von Herrn K. Hattori. (German) [Comments on the work of Mr K. Hattori]. Physikalische Zeitschrift, 29(??):810–811, ???? 1928. CODEN PHZTAO. ISSN 0369-982X.

Infeld:1928:EMT

[Inf28b] Leopold Infeld. Les équations de Maxwell dans la théorie commune à la gravitation et l’électricité. (French) [The equations of Maxwell in the theory common to gravity and electricity]. Comptes Rendus de l’Acad ’emie des Sciences, 186(??):1280–1282, ???? 1928.

Infeld:1928:MDT

[Inf28c] Leopold Infeld. Sur la mesure du temps et la mesure de l’espace dans la physique classique et dans la théorie de la relativité, I partie. (French) [On the measurement of time and the measurement of space in classical physics and in the theory of relativity, Part II]. C.R. Soc. Polon. Phys., 3(??):117–129, ???? 1928. In Polish, with a French summary.

Infeld:1928:PEF

[Inf28d] Leopold Infeld. Zum Problem einer einheitlichen Feldtheorie von Elektrizität und Gravitation. (German) [On the problem of a unified field theory of electricity and gravitation]. Zeitschrift für Physik, 50(1–2):137–152, January 1928. CODEN ZEPYAA. ISSN

0044-3328. URL https://link.springer.com/article/10.1007/
BF01328597.

Infeld:1928:FEG

[Inf28e] Leopold Infeld. Zur Feldtheorie von Elektrizität und Gravitation. (German) [On the field theory of electricity and gravitation]. Physikalische Zeitschrift, 29:145–147, 1928. CODEN PHZTAO. ISSN 0369-982X.

Infeld:1929:KQG

[Inf29] Leopold Infeld. Kausalgesetz und Quantenmechanik. (German) [Causation and quantum mechanics]. Zeitschrift für Physik, 57(5– 6):411–415, May 1929. CODEN ZEPYAA. ISSN 0044-3328. URL

https://link.springer.com/article/10.1007/BF01339628.

Infeld:1930:BAH

[Inf30a] Leopold Infeld. Eine Bemerkung zu der Arbeit von Herrn G. Wataghin: Über die Unbestimmtheitsrelationen der Quantentheorie. (German) [A note on the work of Mr G. Wataghin: On the uncertainty principle of quantum mechanics]. Zeitschrift für Physik, 66(9– 10):708, September 1930. CODEN ZEPYAA. ISSN 0044-3328. URL

https://link.springer.com/article/10.1007/BF01421134.

Infeld:1930:KQI

[Inf30b] Leopold Infeld. Kausalgesetz und Quantenmechanik. II. (German) [Causation and quantum mechanics. II]. Zeitschrift für Physik, 61(9–10):703–711, September 1930. CODEN ZEPYAA. ISSN

0044-3328. URL https://link.springer.com/article/10.1007/
BF01341182.

Infeld:1930:URQ

[Inf30c] Leopold Infeld. On the uncertainty relations in quantum mechanics and their connection with the problem of measurements and causality. In Polish. This is the author’s Docentship paper in Lwów., 1930.

Infeld:1931:SEG

[Inf31a] Leopold Infeld. Über die Struktur der Elektronenwelle. (German) [On the structure of the electron wave]. Bull. Acad. Polon. Sci., Set A, ??(??):201–231, ???? 1931.

Infeld:1931:INE

[Inf31b] Leopold Infeld. Über eine Interpretation der neuen Einsteinschen Weltgeometrie auf dem Boden der klassischen Mechanik. (German) [On an interpretation of the new Einstein’s world geometry on the ground of classical mechanics]. Physikalische Zeitschrift, 32:110–112, 1931. CODEN PHZTAO. ISSN 0369-982X.

Infeld:1931:NGG

[Inf31c] Leopold Infeld. Zur nichtholonomen Geometrie. (German) [On nonholonomic geometry]. Prace Matematyczno-Fizyczne, 39(??):1–9, ???? 1931.

Infeld:1932:VSD

[Inf32a] Leopold Infeld. Die verallgemeinerte Spinorenrechnung und die Diracschen Gleichungen. (German) [Generalized spinor calculation and the Dirac equations]. Physikalische Zeitschrift, 33:475–483, 1932. CODEN PHZTAO. ISSN 0369-982X.

Infeld:1932:GCS

[Inf32b] Leopold Infeld. The generalised calculation of spinores in Dirac’s equation. Physikalische Zeitschrift, 33:475–483, 1932. CODEN PHZ-TAO. ISSN 0369-982X.

Infeld:1932:RPT

[Inf32c] Leopold Infeld. Remarques sur le problème de la théorie unitaire des champs. (French) [Notes on the problem of unitary field theory]. Rend. Acad. Nazionale Lincei, 15(??):157–160, ???? 1932.

Infeld:1933:NDN

[Inf33] Leopold Infeld. Nowe drogi nauki: kwanty i materja (Polish) [New pathways of science: quanta and matter], volume 2 of Z Dziedziny Nauki i Techniki. nakł. Mathesis Polskiej, Warszawa, Poland, 1933. x + 284 pp. LCCN ????

Infeld:1934:DEG

[Inf34a] Leopold Infeld. Dirac’s equation in the general relativity theory. Acta Physica Polonica, 3(??):1–14, ???? 1934. CODEN APPOAK. ISSN 0001-673X.

Infeld:1934:WMS

[Inf34b] Leopold Infeld. The world of modern science: matter and quanta. G. P. Putnam’s Sons, New York, NY, USA, 1934. 287 pp. LCCN ???? Translation by Louis Infield of [Inf33]. Introduction by Albert Einstein.

Infeld:1936:NEF

[Inf36a] L. Infeld. The new electrodynamics and the fine structure constant. Nature, 137(3468):658, April 18, 1936. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Infeld:1936:NAF

[Inf36b] Leopold Infeld. The new action function and the unitary field theory. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 32:127–137, 1936. CODEN PCPSA4. ISSN 0008-1981.

Infeld:1936:WZY

[Inf36c] Leopold Infeld. Wu zhi yu liang zi. (Mandarin) [Material and quantum]. Shang wu yin shu guan, Shanghai, China, 1936. 1 + 5 + 260 (two volumes) pp. LCCN OC 9100 4400 Ser. 2, v. 233. Translation to Mandarin by He Yujie of [Inf34b].

Infeld:1937:LTN

[Inf37a] Leopold Infeld. The Lorentz transformations in the new quantum electrodynamics. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 158(A894):368–371, January 15, 1937. CODEN PRLAAZ. ISSN 0080-4630.

Infeld:1937:NGA

[Inf37b] Leopold Infeld. A new group of action functions in the unitary field theory. II. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 33:70–78, 1937. CODEN PCPSA4. ISSN 0008-1981.

Infeld:1938:EGRa

[Inf38a] Leopold Infeld. Electromagnetic and gravitational radiation. Physical Review, 53(10):836–841, May 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Infeld:1938:EGRb

[Inf38b] Leopold Infeld. Electromagnetic and gravitational radiation [abstract only]. Physical Review, 53(11):938, June 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

https://journals.aps.org/pr/pdf/10.1103/PhysRev.53.913.

Infeld:1938:FDR

[Inf38c] Leopold Infeld. The fourth dimension and relativity. A radio broadcast given by the Galois Institute of Mathematics over Station WNYC, Long Island University., 1938.

Infeld:1939:BRP

[Inf39] Leopold Infeld. Book review: The Principles of Statistical Mechanics by Richard C. Tolman. Philosophy of Science, 6(3):381, July 1939. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184744.

Infeld:1940:BRM

[Inf40a] Leopold Infeld. Book review: Modern Science by Hyman Levy. Science & Society, 4(2):235–237, Spring 1940. ISSN 0036-8237 (print), 1943-2801 (electronic). URL http://www.jstor.org/

stable/40399332.

Infeld:1940:TBM

[Inf40b] Leopold Infeld. On the theory of Brownian motion. University of Toronto Studies, Applied Mathematics Series, 4:1–42, 1940.

Infeld:1941:NTS

[Inf41a] Leopold Infeld. On a new treatment of some eigenvalue problems. Physical Review (2), 59(9):737–747, May 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Infeld:1941:QESa

[Inf41b] Leopold Infeld. Quest: the evolution of a scientist. Doubleday, Doran, and Co., New York, NY, USA, 1941. iv + 342 pp. LCCN QC16.I6 A3 1941.

Infeld:1941:QESb

[Inf41c] Leopold Infeld. Quest: the evolution of a scientist. Scientific Book Club, London, UK, 1941. 288 pp. LCCN ????

Infeld:1942:GFM

[Inf42a] Leopold Infeld. A generalization of the factorization method for solving eigenvalue problems. Trans. Roy. Soc. Canada. Sect. III. (3), 36:7–18, 1942.

Infeld:1942:QA

[Inf42b] Leopold Infeld. Quest: an autobiography. Readers Union, London, UK, 1942. ISBN 0-88738-796-9. viii + 272 pp. LCCN QC16.I6 A3 1942.

Infeld:1943:BRI

[Inf43a] Leopold Infeld. Book Review: Introduction to the theory of relativity. Bulletin of the American Mathematical Society, 49(7):527–529, 1943. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Infeld:1943:CRR

[Inf43b] Leopold Infeld. Clocks, rigid rods and relativity theory. American Journal of Physics, 11:219–222, 1943. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Infeld:1945:LNA

[Inf45] Leopold Infeld. Letter: A new approach to relativistic cosmology. Nature, 156(3952):114, July 28, 1945. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://adsabs.

harvard.edu/abs/1945Natur.156..114I.

Infeld:1946:AEW

[Inf46a] Leopold Infeld. Atomic energy and world government. Pamphlet, Canadian Institute of International Affairs, Toronto, ON, Canada, 1946.

Infeld:1946:GRS

[Inf46b] Leopold Infeld. General relativity and the structure of our universe. In Alfred E. Schild, editor, ????, page ?? ????, ????, 1946. LCCN ????

Infeld:1947:ANS

[Inf47a] Leopold Infeld. Al this newe science that men lere. The Scientific Monthly, 65(5):437–438, 1947. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://www.jstor.org/

stable/19475.

Infeld:1947:BRS

[Inf47b] Leopold Infeld. Book review: The Strange Story of the Quantum. The Scientific Monthly, 65(5):437–438, November 1947. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic).

Infeld:1947:IWG

[Inf47c] Leopold Infeld. The influence of the width of the gap upon the theory of antennas. Quarterly of Applied Mathematics, 5(2):113– 132, 1947. CODEN QAMAAY. ISSN 0033-569X (print), 1552-4485 (electronic).

Infeld:1947:PE

[Inf47d] Leopold Infeld. Portrait: Einstein. American Scholar, 16(3):337– 341, 1947. ISSN 0003-0937 (print), 2162-2892 (electronic). URL

http://www.jstor.org/stable/41204928.

Infeld:1947:RFC

[Inf47e] Leopold Infeld. Recurrence formulas for Coulomb wave function. Physical Review (2), 72:1125, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Infeld:1947:TRM

[Inf47f] Leopold Infeld. Theory of radioactivity in the momentum space [abstract only]. Physical Review, 72(2):154, ???? 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Infeld:1948:LOT

[Inf48a] Leopold Infeld. Letter: One Two Three … Infinity. The Scientific Monthly, 66:440–441, May 1948. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic).

Infeld:1948:RMB

[Inf48b] Leopold Infeld. Review: Mathematics—and beyond: One Two Three… Infinity by George Gamow. The Scientific Monthly, 66 (5):440–441, May 1948. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://www.jstor.org/

stable/19511.

Infeld:1948:WGL

[Inf48c] Leopold Infeld. Whom the Gods Love: the Story of Évariste Galois. Classics in mathematics education. McGraw-Hill, New York, NY, USA, 1948. ix + 323 pp. LCCN QA29 .G25 I5.

Infeld:1949:GRS

[Inf49a] Leopold Infeld. On the structure of our universe. In Schilpp [Sch49b], chapter 18, pages 477–499. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Infeld:1949:PII

[Inf49b] Leopold Infeld. Part IV: The impedance of a rectangular wave guide with a thin antenna. Canadian Journal of Research, Section A: Physical Sciences, 27a(4):109–124, July 1949. CODEN CNRAAU. ISSN 0366-7383.

Infeld:1949:VD

[Inf49c] Leopold Infeld. Visit to Dublin. Scientific American, 181(4): 11–15, October 1949. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v181/n4/pdf/scientificamerican1049-
11.pdf.

Infeld:1949:VE

[Inf49d] Leopold Infeld. Visit to England. Scientific American, 181(5): 40–43, November 1949. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v181/n5/pdf/scientificamerican1149-
40.pdf.

Infeld:1949:VP

[Inf49e] Leopold Infeld. Visit to Poland. Scientific American, 181(6): 40–43, December 1949. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v181/n6/pdf/scientificamerican1249-
40.pdf.

Infeld:1950:AEH

[Inf50a] Leopold Infeld. Albert Einstein: his work and its influence on our world. Scribner, New York, NY, USA, revised edition, 1950. v + 132 pp. LCCN ????

Infeld:1950:AES

[Inf50b] Leopold Infeld. Albert Einstein speaks his mind: Out of my Later Years, by Albert Einstein. New York Times, ??(??):146, May 28, 1950. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/111436294/.

Infeld:1950:BRP

[Inf50c] Leopold Infeld. Book review: The Principles of Scientific Research. The Scientific Monthly, 71:135–136, August 1950. CODEN SC-MOAA. ISSN 0096-3771 (print), 2327-7513 (electronic).

Infeld:1950:FMA

[Inf50d] Leopold Infeld. The factorization method and its application to differential equations in theoretical physics. Proc. Symposia Appl. Math., 2:58–65, 1950.

Infeld:1950:LNE

[Inf50e] Leopold Infeld. Letter: The new Einstein theory and the equations of motion. Nature, 166(4234):1075, December 23, 1950. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Infeld:1950:NET

[Inf50f] Leopold Infeld. The new Einstein theory and the equations of motion. Nature, 166(4234):1075, December 23, 1950. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http:/

/www.nature.com/nature/journal/v166/n4234/pdf/1661075a0.
pdf.

Infeld:1950:ENT

[Inf50g] Leopold Infeld. On Einstein’s new theory. American Scholar, 19(4): 423–433, Autumn 1950. ISSN 0003-0937 (print), 2162-2892 (elec-

tronic). URL http://www.jstor.org/stable/41206693.

Infeld:1950:WBL

[Inf50h] Leopold Infeld. Wybrańcy bogów …: powieść o życiu Ewarysta Galois. (Polish) [Goddesses …: Life Story of Ewarysta Galois], volume 7 of Biblioteka “Gazety Robotniczej”. Ksiaźka i Wiedza, Warszawa, Poland, second edition, 1950. 263 + 2 pp. LCCN ???? Translation to Polish by Stanisław Sielski of [Inf48c].

Infeld:1951:NET

[Inf51a] Leopold Infeld. The new Einstein theory and the equations of motion. Acta Physica Polonica, 10:284–293, 1951. CODEN APPOAK. ISSN 0587-4254 (print), 1509-5770 (electronic).

Infeld:1951:ENT

[Inf51b] Leopold Infeld. On Einstein’s new theory. In Annual report of the Board of Regents of the Smithsonian Institution, pages 189–197. Smithsonian Institution, Washington, DC, USA, 1951. LCCN QC 6

.I54 1952. URL https://babel.hathitrust.org/cgi/pt?id=uc1.

31210022604829. Reprint of [Inf50g].

Infeld:1951:WBL

[Inf51c] Leopold Infeld. Wybrańcy bogów …: powieść o życiu Ewarysta Galois. (Polish) [Goddesses …: Life Story of Ewarysta Galois], volume 7 of Biblioteka “Gazety Robotniczej”. Ksiażka i Wiedza, Warszawa, Poland, second edition, 1951. 309 + 1 pp. LCCN ???? Translation to Polish by Stanisław Sielski of [Inf48c].

Infeld:1952:LTA

[Inf52] Leopold Infeld. Letter: Is there an aether? Nature, 169(4304): 702, April 26, 1952. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Infeld:1953:AES

[Inf53a] Leopold Infeld. Albert Einstein: sein Werk und sein Einfluß auf unsere Welt. (German) [Albert Einstein: his work and his influence on our world]. Schönbrunn-Verlag, Wien, Austria, 1953. 173 pp. LCCN ????

Infeld:1953:CCE

[Inf53b] Leopold Infeld. The coordinate conditions and the equations of motion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(1):17–25, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Infeld:1953:DJS

[Inf53c] Leopold Infeld. [Death of Joseph Stalin]. Isvestia, ??(??):??, ???? 1953. Cited in [Inf78e, page 84].

Infeld:1953:ECH

[Inf53d] Leopold Infeld. An electronic cloud in a homogeneous electric and magnetic field according to Dirac’s theory. Bulletin de l’Académie polonaise des sciences. Classe troisième, 1(??):99–104, ???? 1953. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1953:LVF

[Inf53e] Leopold Infeld. Leonardo Da Vinci and the fundamental laws of science. Science & Society, 17(1):26–41, 1953. ISSN 0036-8237 (print), 1943-2801 (electronic). URL http://www.jstor.org/

stable/40400163.

Infeld:1953:UAM

[Inf53f] Leopold Infeld. On the use of an approximation method in Dirac’s electrodynamics. Bulletin de l’Académie polonaise des sciences.

Classe troisième, 1:18–22, 1953. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1953:RVF

[Inf53g] Leopold Infeld. [rebuttal to V. A. Fock’s view of Albert Einstein]. Mysl Filosoficzna, ??(??):??, ???? 1953. Cited in [Inf78e, page 80].

Infeld:1953:JEK

[Inf53h] Leopold Infeld. Über die jüngste Entwicklung der klassischen Elektrodynamik. (German) [On the recent development of classical electrodynamics]. Fortschritte der Physik = Progress of Physics, 1(??): 88–98, ???? 1953. CODEN FPYKA6. ISSN 0015-8208 (print), 1521-3978 (electronic).

Infeld:1953:WGL

[Inf53i] Leopold Infeld. Wen die Götter lieben: die Geschichte des Évariste Galois. (German) [Whom the gods love: the story of Évariste Galois]. Schönbrunn-Verlag, Wien, Austria, 1953. 330 pp. LCCN ????

Infeld:1954:EMN

[Inf54a] Leopold Infeld. Equations of motion and non-harmonic coordinate conditions. Bulletin de l’Académie polonaise des sciences. Classe troisième, 2:163–166, 1954. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1954:FTV

[Inf54b] Leopold Infeld. Frieden oder totale Vernichtung. (German) [Peace or total destruction]. ????, Berlin, West Germany, 1954. 7 pp. LCCN ????

Infeld:1954:DAE

[Inf54c] Leopold Infeld. [In defense of Albert Einstein]. Mysl Filosoficzna, ??(??):??, ???? 1954. Cited in [Inf78e, page 85].

Infeld:1954:MZK

[Inf54d] Leopold Infeld. Materiały z konferencji fizyków w Spałe. (Polish) [Materials from physics conference in Spałe]. Państwowe Wydawnictwo Naukowe, Warszawa, Poland, 1954. 366 pp. LCCN ???? Sprawozdanie z III konferencji ogólnopolskiej fizyków w Spałe, 1-14 września 1952 [Report of the 3rd National Physics Conference in Spałe, 1-14 September 1952].

Infeld:1954:MBG

[Inf54e] Leopold Infeld. On the motion of bodies in general relativity theory. Acta Physica Polonica, 13:187–204, 1954. CODEN APPOAK. ISSN 0587-4254 (print), 1509-5770 (electronic).

Infeld:1954:WGL

[Inf54f] Leopold Infeld. Wen die Götter lieben: die Geschichte des Évariste Galois. (German) [Whom the gods love: the story of Évariste Galois]. Schönbrunn-Verlag, Wien, Austria, 1954. 330 pp. LCCN ????

Infeld:1955:AE

[Inf55a] Leopold Infeld. [Albert Einstein]. Twórczość, ??(??):??, ???? 1955.

Infeld:1955:DAE

[Inf55b] Leopold Infeld. [Death of Albert Einstein]. L’Humanité, ??(??):??, ???? 1955. Cited in [Inf78e, page 85], and reprinted in several Polish newspapers. Later enlarged and published in [Inf55a].

Infeld:1955:GRG

[Inf55c] Leopold Infeld. Die Geschichte der Relativitätstheorie. (German) [The history of the Theory of Relativity]. Naturwissenschaften, 42 (15):431–436, August 1, 1955. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). English translation in [].

Infeld:1955:EBR

[Inf55d] Leopold Infeld. Einige Bemerkungen über die Relativitätstheorie. (German) [Some remarks on relativity theory]. Annalen der Physik 6 (Berlin, Germany), 16(5–8):229–240, 1955. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Infeld:1955:EML

[Inf55e] Leopold Infeld. Equations of motion for linear field-theories. Bulletin de l’Académie polonaise des sciences. Classe troisième, 3:213–216, 1955. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1955:HRT

[Inf55f] Leopold Infeld. The history of relativity theory. Rend. Mat. e Appl. (5), 13:270–281, 1955. ISSN 0034-4427. Translation to English of [Inf55c].

Infeld:1955:IRT

[Inf55g] Leopold Infeld. Istoriya razvitiya teorii otnositelnosti. (Russian) [History of the development of the theory of relativity]. Uspekhi Fizicheskikh Nauk, 57(2):193–203, 1955. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/ru/

articles/1955/10/c/.

Infeld:1955:OKD

[Inf55h] Leopold Infeld. Od Kopernika do Einsteina. Polish [From Copernicus to Einstein]. Kosmos, Series B, 1(??):209–226, 1955.

Infeld:1955:WGL

[Inf55i] Leopold Infeld. Wen die Götter lieben: die Geschichte des Évariste Galois. (German) [Whom the gods love: the story of Évariste Galois]. Die Buchgemeinde, Wien, Austria, 1955. 330 pp. LCCN ????

Infeld:1956:AEJ

[Inf56a] Leopold Infeld. Albert Einstein, jego dzieło i rola w nauce. (Polish) [Albert Einstein his work and its influence on our world]. Państwowe wydawnictwo naukowe, Warszawa, Poland, 1956. 192 pp. LCCN ???? Translation to Polish by Ryszard Gajewski (1930–2014) of [Inf50a].

Infeld:1956:AES

[Inf56b] Leopold Infeld. Albert Einstein. Sein Werk und sein Einfluß auf unsere Welt. (German) [Albert Einstein. His work and his influence on our world]. Akademie-Verlag, Berlin, Germany, 1956. 174 pp. LCCN ????

Infeld:1956:MVO

[Inf56c] Leopold Infeld. Moi Vospominaniya ob Einshteine. (Russian) [My memories of Einstein]. Uspekhi Fizicheskikh Nauk, 59(1):135–184, 1956. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (elec-

tronic). URL http://ufn.ru/ru/articles/1956/5/i/.

Infeld:1956:MWE

[Inf56d] Leopold Infeld. Moje wspomnienia o Einsteinie. (Polish) [My Memories of Einstein]. Iskry, Warszawa, Poland, 1956. 148 + 3 pp. LCCN ????

Infeld:1956:EMG

[Inf56e] Leopold Infeld. On equations of motion in general relativity theory. Helv. Phys. Acta Supp., 4(??):206–209, ???? 1956.

Infeld:1956:RPA

[Inf56f] Leopold Infeld. [Remarks for the Polish Academy of Sciences]. Przeglad Kulturalny, ??(??):??, ???? ??, 1956. Cited in [Inf78e, page 103].

Infeld:1957:AESb

[Inf57a] Leopold Infeld. Albert Einstein: A sua obra e asua influência no mundo contemporâneo. (Portuguese) [Albert Einstein: His work and influence in the contemporary world]. Publicações Europa-América, Lisboa, Portugal, 1957. 181 + 11 pp. LCCN ???? Translation to Portuguese by Fernando de Macedo of [Inf50a].

Infeld:1957:AESa

[Inf57b] Leopold Infeld. Albert Einstein: sein Werk und sein Einfluß auf unsere Welt. (German) [Albert Einstein: his work and his influence on our world]. Akademie-Verlag, Berlin, Germany, 1957. 173 pp. LCCN ????

Infeld:1957:BRG

[Inf57c] Leopold Infeld. Begriff des Raumes in der Geometrie. (German) [on the equations of motion]. Ber. Riemann-Tagung Forsch.-Inst. Math., 1:202–209, 1957.

Infeld:1957:EMGa

[Inf57d] Leopold Infeld. The equations of motion in general relativity theory and the action principle. Acta Physica Polonica, 16:177–210, 1957. CODEN APPOAK. ISSN 0587-4254 (print), 1509-5770 (electronic).

Infeld:1957:EMGb

[Inf57e] Leopold Infeld. Equations of motion in General Relativity theory and the action principle. Reviews of Modern Physics, 29 (3):398–411, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.29.398;   http://rmp.aps.org/
abstract/RMP/v29/i3/p398_1.

Infeld:1957:EGA

[Inf57f] Leopold Infeld. Évariste Galois, aimé des dieux. (French) [Évariste Galois, loved by the gods]. Éditions la Farandole, Paris, France, 1957. 351 pp. LCCN ???? Translation to French by Joseph Sully of [Inf48c].

Infeld:1957:LFO

[Inf57g] Leopold Infeld. The Lagrangian as a function only of coordinates and the mechanical spin of a particle. Bulletin de l’Académie polonaise des sciences. Classe troisième, 5:985–989, lxxxii, 1957. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1957:LHO

[Inf57h] Leopold Infeld. The Lagrangian with higher order derivatives and the mechanical spin of a particle. Bulletin de l’Académie polonaise des sciences. Classe troisième, 5:979–983, lxxxii, 1957. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1957:NDN

[Inf57i] Leopold Infeld. Nowe drogi nauki : materia, promieniowanie. (Polish) [New scientific paths: matter, radiation]. Wiedza Powszechna, Warszawa, Poland, second edition, 1957. 258 + 1 + 2 pp. LCCN ???? Prepared by Włodzimierz Kołos (1928–1996).

Infeld:1957:EM

[Inf57j] Leopold Infeld. On the equations of motion. Schr. Forschungsinst. Math., 1:202–209, 1957.

Infeld:1957:LSR

[Inf57k] Leopold Infeld. On the Lagrangian in special relativity theory. Bulletin de l’Académie polonaise des sciences. Classe troisième, 5:491– 495, xli, 1957. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1957:TOP

[Inf57l] Leopold Infeld. Tredici ore per l’immortalità. La vita del matematico Galois (Italian) [13 hours for immortality: the life of the mathematician Galois]. Feltrinelli Editore, Milano, Italia, 1957. 331 pp. Translation to Italian by Giovanni Galtieri of [Inf48c].

Infeld:1957:WGL

[Inf57m] Leopold Infeld. Wen die Götter lieben: die Geschichte des Évariste Galois. (German) [Whom the gods love: the story of Évariste Galois]. Rutten & Loening, Berlin, West Germany, 1957. 417 pp. LCCN ???? Translation to German by Doris Brehm of [Inf48c].

Infeld:1958:AEJ

[Inf58a] Leopold Infeld. Albert Einstein : jeho dílo a vliv na ná sv t. (Czech) [Albert Einstein: His work and influence on our world], volume 74 of Universita vojáka. Naše vojsko, Praha, Czechoslovakia, 1958. 117 + 4 pp. LCCN ???? Translation to Czech by Lenka Teigová-Stachová of the Polish original.

Infeld:1958:PFV

[Inf58b] Leopold Infeld. Planck’s Festschrift “On variational principles in relativistic dynamics”. VEB Verlag der Deutsch. Akad. der Wissenschaften, Berlin, ??(??):115–127, 1958.

Infeld:1958:PSL

[Inf58c] Leopold Infeld. Prof. Stanisław Loria. Nature, 182(4647):1411– 1412, November 22, 1958. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v182/n4647/pdf/1821411b0.pdf.

Infeld:1959:EMG

[Inf59a] Leopold Infeld. Equations of motion and gravitational radiation. Annals of Physics, 6(4):341–367, 1959. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035x (electronic).

Infeld:1959:MVN

[Inf59b] Leopold Infeld. Mé vzpomínky na Einstein. (Czech) [My Memories of Einstein], volume 171 of Knihovna eskoslovenské Spolenosti pro íení politických a vdeckých znalostí. Orbis, Praha, Czechoslovakia, 1959. 79 pp. LCCN ???? Translation to Czech by Bojana Fiedlerová of the Polish original.

Infeld:1960:AES

[Inf60a] Leopold Infeld. Albert Einstein: A sua obra e asua influência no mundo contemporâneo. (Portuguese) [Albert Einstein: His work and influence in the contemporary world]. Publicações Europa-América, Lisboa, Portugal, third edition, 1960. 187 pp. LCCN ???? Translation to Portuguese by Fernando de Macedo of [Inf50a].

Infeld:1960:EGK

[Inf60b] Leopold Infeld. Everist Galua. Molodaya gvardiya, Moscow, USSR, 1960. 365 + 3 + 6 pp. LCCN ???? Translation to Russian of [Inf48c].

Infeld:1960:NFG

[Inf60c] Leopold Infeld. A new form of the geodesic line equation. Bulletin de l’Académie Polonaise des Sciences, Série des Sciences, Mathématiques, Astronomiques et Physiques, 8:559–561, 1960. CO-DEN BAPMAM. ISSN 0001-4117.

Infeld:1961:AES

[Inf61a] Leopold Infeld. Albert Einstein: A sua obra e asua influência no mundo contemporâneo. (Portuguese) [Albert Einstein: His work and influence in the contemporary world]. Publicações Europa-América, Lisboa, Portugal, third edition, 1961. 181 pp. LCCN ???? Translation to Portuguese by Fernando de Macedo of [Inf50a].

Infeld:1961:AEH

[Inf61b] Leopold Infeld. Albert Einstein: his work and its influence on our world. Scribner, New York, NY, USA, revised edition, 1961. v + 134 pp. LCCN ????

Infeld:1961:AM

[Inf61c] Leopold Infeld. The EIH and the k-approximation methods. Bulletin de l’Académie Polonaise des Sciences, Série des Sciences, Mathématiques, Astronomiques et Physiques, 9:93–97, 1961. CO-DEN BAPMAM. ISSN 0001-4117.

Infeld:1961:ESO

[Inf61d] Leopold Infeld. Einstein, su obra y su influencia en nuestro mundo. (Spanish) [Einstein, his work and his influence in our world]. Lautaro, Buenos Aires, Argentina, 1961. 174 pp. LCCN ????

Infeld:1961:EG

[Inf61e] Leopold Infeld. Évariste Galois. Mladé letá, Bratislava, Czechoslovakia, 1961. 321 pp. LCCN ???? Translation to Czech by Ružena Jamrichová.

Infeld:1961:FSD

[Inf61f] Leopold Infeld. Freedom and scientific disciplines. [unknown Polish newspaper], ??(??):??, ???? 1961.

Infeld:1961:PCC

[Inf61g] Leopold Infeld. Is Planck’s constant a constant in a gravitational field? Bulletin de l’Académie Polonaise des Sciences, Série des Sciences, Mathématiques, Astronomiques et Physiques, 9:617–620, 1961. CODEN BAPMAM. ISSN 0001-4117.

Infeld:1961:LTR

[Inf61h] Leopold Infeld. Letter: Theory of red shift and nature of Plancks’s constant. Nature, 191(479):1184, September 16, 1961. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Infeld:1961:MCL

[Inf61i] Leopold Infeld. On the most Cartesian-like co-ordinate system. Bulletin de l’Académie Polonaise des Sciences, Série des Sciences, Mathématiques, Astronomiques et Physiques, 9:299–302, 1961. CO-DEN BAPMAM. ISSN 0001-4117.

Infeld:1961:TRS

[Inf61j] Leopold Infeld. Theory of the red shift and the nature of Planck’s constant. Nature, 191(4794):1184, September 16, 1961. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Infeld:1962:FS

[Inf62] Leopold Infeld. Freedom for science. Bulletin of the Atomic Scientists, 18(5):31, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Infeld:1962:PCT

[Inf63a] Leopold Infeld. Planck’s constant and the theory of the red shift. Zeitschrift für Physik, 171:34–43, 1962–1963. CODEN ZEPYAA. ISSN 0170-9739 (print), 1431-5858 (electronic).

Infeld:1963:AEU

[Inf63b] Leopold Infeld. Albert Einstein: l’uomo e lo scienziato: la teoria della relatività e la sua influenza sul mondo contemporaneo. (Italian) [Albert Einstein: man and scientist: the theory of relativity and its influence on the contemporary world], volume 23 of Piccola biblioteca Einaudi. Giulio Einaud, Torino, Italia, fourth edition, 1963. 145 pp. LCCN ???? Translation to Italian by Orazio Nicotra.

Infeld:1963:LGR

[Inf63c] Leopold Infeld. The Lagrangian in general relativity theory (G.R.T.) and radiation. Bulletin de l’Académie Polonaise des Sciences, Série des Sciences, Mathématiques, Astronomiques et Physiques, 11:399– 405, 1963. CODEN BAPMAM. ISSN 0001-4117.

Infeld:1963:PCT

[Inf63d] Leopold Infeld. Planck’s constant and theory of red shift. Zeitschrift für Physik, 171(1):34–43, February 1963. CODEN ZEPYAA. ISSN 0044-3328.

Infeld:1963:UAM

[Inf63e] Leopold Infeld. “Uniformly accelerated” motion and relativity. Acta Physica Polonica, 23:69–75, 1963. CODEN APPOAK. ISSN 0587-4254 (print), 1509-5770 (electronic).

Infeld:1964:PTG

[Inf64a] Leopold Infeld, editor. Proceedings on Theory of Gravitation: Conference in Warszawa and Jabłonna, 25–31 July, 1962. Conférence internationale sur les théories relativistes de la gravitation. Gauthier-Villars et cie, Paris, France, 1964. LCCN QC178 .C66 1962.

Infeld:1964:SZP

[Inf64b] Leopold Infeld. Szkice z przeszłości: wspomnienia. (Polish) [Sketches of the past: memories]. Państwowy Instytut Wydawniczy, Warszawa, Poland, 1964. 298 + 2 pp. LCCN QC16.I6 A32.

Infeld:1964:UDI

[Inf64c] Leopold Infeld. Uravnenie dvizhenija izluchajushhego jelektrona i ego lagranzhian. (Russian) [The equations of motion of a radiating electron and its Lagrangian]. Acta physica Academiae Scientiarum Hungaricae, 17:7–14, 1964. CODEN APAHAQ. ISSN 0231-4428.

Infeld:1965:SI

[Inf65a] Leopold Infeld. As I see it. Bulletin of the Atomic Scientists, 21(2): 7–14, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Infeld:1965:BVH

[Inf65b] Leopold Infeld. Bhautiki ka vikas. (Hindi) [The Evolution of Physics]. Argus Pub. Co., New Delhi, India, 1965. 237 pp. LCCN ???? Translation to Hindi by Ram Kumar Mishra of [EI38b].

Infeld:1965:RGP

[Inf65c] Leopold Infeld. Remarks on generalized problem of motion. Acta Physica Austriaca, 19(4):380–384, 1965. CODEN APASAP. ISSN 0001-6713.

Infeld:1966:SZP

[Inf66] Leopold Infeld. Szkice z przeszłości: wspomnienia. (Polish) [Sketches of the past: memories]. Państwowy Instytut Wydawniczy, Warszawa, Poland, second edition, 1966. 298 + 2 pp. LCCN QC16.I6 A32.

Infeld:1967:SFE

[Inf67] Leopold Infeld. On a simple formula for energy radiation. Bulletin de l’Académie Polonaise des Sciences — Serie des sciences mathématiques astronomiques et physiques, 15(4):293–295, 1967.

Infeld:1968:AEU

[Inf68a] Leopold Infeld. Albert Einstein: l’uomo e lo scienziato: la teoria della relatività e la sua influenza sul mondo contemporaneo. (Italian) [Albert Einstein: man and scientist: the theory of relativity and its influence on the contemporary world]. Giulio Einaud, Torino, Italia, sixth edition, 1968. 145 pp. LCCN ???? Translation to Italian by Orazio Nicotra.

Infeld:1968:KFJ

[Inf68b] Leopold Infeld. Kordian, fizyka i ja : wspomnienia. (Polish) [Kordian, physics and I: memories]. Państwowy Instytut Wydawniczy, Warszawa, Poland, 1968. 223 pp. LCCN ????

Infeld:1969:GNS

[Inf69a] Leopold Infeld. Garoa no shōgai, kkamigami no aideshi hito. (Japanese) []. Nihon hyōronsha, Tokyo, Japan, 1969. x + 388 pp. LCCN ???? Translation to Japanese of [Inf48c].

Infeld:1969:LEK

[Inf69b] Leopold Infeld. Leben mit Einstein : Kontur einer Erinnerung. (German) [Living with Einstein: Contours of a memory]. Europa-Verlag, Wien, Austria, 1969. 223 pp. LCCN ???? Translation to German of ???? by Walter Hacker.

Infeld:1970:LIB

[Inf70] Eryk Infeld. Leopold Infeld bibliography. General Relativity and Gravitation, 1(2):191–208, June 1970. CODEN GRGVA8. ISSN 0001-7701 (print), 1572-9532 (electronic). URL https://link.

springer.com/article/10.1007/BF00756897.

Infeld:1973:CE

[Inf73] Leopold Infeld. From Copernicus to Einstein. In Barbara Bieńkowska, editor, The Scientific World of Copernicus on the occasion of the 500th anniversary of his birth 1473–1973, pages 66–83. D. Reidel Publishing Company, Dordrecht, The Netherlands / Boston, MA, USA, 1973. ISBN 94-010-2618-1. LCCN QB36.C8. URL http:

//adsabs.harvard.edu/abs/1973swc..conf...66I. Foreword by

Zdeněk Kopal.

Infeld:1974:AEU

[Inf74a] Leopold Infeld. Albert Einstein: l’uomo e lo scienziato: la teoria della relatività e la sua influenza sul mondo contemporaneo. (Italian) [Albert Einstein: man and scientist: the theory of relativity and its influence on the contemporary world]. Giulio Einaud, Torino, Italia, 1974. 145 pp. LCCN ???? Translation to Italian by Orazio Nicotra.

Infeld:1974:EGK

[Inf74b] Leopold Infeld. Everist Galua: Kniga dlja tenija na francuzskom jazyke v 9-10 klassach srednej školy. (Russian) [Évariste Galois: book for reading in the French language in 9–10 middle-school classes]. Prosveščenije, Moscow, USSR, 1974. 143 pp. LCCN ????

Infeld:1978:EDH

[Inf78a] Leopold Infeld. El elegido de los dioses: la historia de Évariste Galois. (Spanish) [The chosen of the gods: the story of Évariste Galois]. El Hombre y sus obras. Siglo Veintiuno Editores, México, DF, México, 1978. ISBN 968-23-0045-2. 355 pp. LCCN ???? Translation to Spanish by Roberto Eugenio Bixio of [Inf48c].

Infeld:1978:REG

[Inf78b] Leopold Infeld. Le Roman d’Évariste Galois. (French) [The story of Évariste Galois]. Éditions la Farandole, Paris, France, second edition, 1978. ISBN 2-7047-0055-9. 361 pp. LCCN ???? Translation to French by Joseph Sully of [Inf48c].

Infeld:1978:O

[Inf78c] Leopold Infeld. Oppenheimer. In Why I left Canada: reflections on science and politics [Inf78e], pages 160–180. ISBN 0-7735-0272-6. LCCN QC16.I6 A3213. Infeld makes a strong rebuttal in support of Albert Einstein against a speech made by J. Robert Oppenheimer on 14 December 1965. He also records harsh criticism by Philip Morrison of Oppenheimer.

Infeld:1978:WGL

[Inf78d] Leopold Infeld. Whom the Gods Love: the Story of Évariste Galois, volume 7 of Classics in mathematics education. National Council of Teachers of Mathematics, Reston, VA, USA, 1978. ISBN 0-87353-125-6. xv + 323 pp. LCCN ???? Reprint of [Inf48c].

Infeld:1978:WLC

[Inf78e] Leopold Infeld. Why I left Canada: reflections on science and politics. McGill-Queen’s University Press, Montréal, Québec, Canada, 1978. ISBN 0-7735-0272-6. xii + 212 pp. LCCN QC16.I6 A3213. Translations from Polish to English by Helen Infeld. Edited with introduction and notes by Lewis Pyenson. Foreword by Alfred Schild (1921–1977).

Infeld:1978:WN

[Inf78f] Leopold Infeld. Władysław Natanson. In Why I left Canada: reflections on science and politics [Inf78e], pages 115–122. ISBN 0-7735-0272-6. LCCN QC16.I6 A3213. Translations from Polish to English by Helen Infeld. Edited with introduction and notes by Lewis Pyenson. Foreword by Alfred Schild (1921–1977).

Infeld:1979:AAN

[Inf79a] Leopold Infeld. Albert Ajnštajn: njegova dela i njihov uticaj na naš svet. (Serbian) [Albert Einstein: his work and its impact on our world], volume 7 of Biblioteka Zanimljiva nauk. Nolit, Beograd, Yugoslavia, 1979. 131 pp. LCCN ???? Translation to Serbian by Branislav Lalović of [Inf50a].

Infeld:1979:AE

[Inf79b] Leopold Infeld. Albert Einstein. Państwowy Wydawnictwo Naukowe, Warszawa, Poland, second edition, 1979. ISBN 83-01-00079-1. 188 + 4 + 1 pp. LCCN ???? Translation to Polish by Ryszard Gajewski (1930–2014) of [Inf50a].

Infeld:1980:QA

[Inf80] Leopold Infeld. Quest: an autobiography. Chelsea Pub. Co., New York, NY, USA, second edition, 1980. ISBN 0-8284-0309-0 (hardcover), 0-8218-4073-8. 361 pp. LCCN QC16.I6 A3 1980.

Infeld:1981:EGI

[Inf81] Leopold Infeld. Evarist Galua : Izbrannik bogov. (Russian) [Évariste Galois: The Chosen of the Gods]. Molodaya gvardiya, Moscow, USSR, third edition, 1981. 350 pp. LCCN ???? Translation to Russian by M. Kan of [Inf48c].

Infeld:1983:ESO

[Inf83] Leopold Infeld. Einstein, su obra y su influencia en nuestro mundo. (Spanish) [Einstein, his work and his influence in our world]. Editorial Leviatán, Buenos Aires, Argentina, 1983. 197 pp. LCCN ????

Infeld:1984:AE

[Inf84] Leopold Infeld. Albert Einstein. Państwowy Wydawnictwo Naukowe, Warszawa, Poland, third edition, 1984. ISBN 83-01-05210-4. 188 + 4 + 1 pp. LCCN ???? Translation to Polish by Ryszard Gajewski (1930–2014) of [Inf50a].

Infeld:1989:QES

[Inf89] Leopold Infeld. Quest: the evolution of a scientist. Transaction Publishers, New Brunswick, NJ, USA, 1989. ISBN 0-88738-796-9. ???? pp. LCCN QC16.I6 A318 1989. With a new introduction by Peter Bergmann.

Infeld:1992:AEU

[Inf92] Leopold Infeld. Albert Einstein: l’uomo e lo scienziato: la teoria della relatività e la sua influenza sul mondo contemporaneo. (Italian) [Albert Einstein: man and scientist: the theory of relativity and its influence on the contemporary world], volume 23 of Piccola biblioteca Einaudi. Giulio Einaudi, Torino, Italia, thirteenth edition, 1992. ISBN 88-06-04416-8. 145 pp. LCCN ????

Infeld:1998:WBP

[Inf98] Leopold Infeld. Wybrańcy bogów: powieść o życiu Ewarysta Galois. (Polish) [Goddesses …: Life Story of Ewarysta Galois]. Prószynski i S-ka, Warszawa, Poland, 1998. ISBN 83-7180-225-0. 319 + 4 pp. LCCN ???? Translation to Polish by Stanisław Sielski of [Inf48c].

Infeld:1999:MNE

[Inf99] Leopold Infeld. Der Mann neben Einstein : ein Leben zwischen Raum und Zeit. (German) [The man next to Einstein: a life between space and time]. Weymann-Bauer-Verlag, Rostock, Germany, 1999. ISBN 3-929395-42-8. 234 pp. LCCN ???? Translation to German by Kurt Kelm of original Polish edition [Inf64b].

Infeld:2006:QA

[Inf06] Leopold Infeld. Quest: an autobiography. AMS Chelsea Publishing, Providence, RI, USA, second edition, 2006. ISBN 0-8284-0309-0 (hardcover), 0-8218-4073-8. 361 pp. LCCN QC16.I6 A3 1980. URL

http://bookstore.ams.org/chel-309/.

Infeld:2015:MGE

[Inf15] Eva Infeld. My grandfather and Einstein. Web site, Jan-

uary 6, 2015. URL https://www.fiatphysica.com/blog/people/
einstein-theory-of-relativity-evolution-of-physics.

Infeld:1953:EP

[IP53] Leopold Infeld and Jerzy Plebański. Electrodynamics without potentials. Acta Physica Polonica, 12:123–134, 1953. CODEN APPOAK. ISSN 0587-4254 (print), 1509-5770 (electronic).

Infeld:1954:EP

[IP54] Leopold Infeld and Jerzy Plebański. Electrodynamics without potentials. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 222(1149):224–227, March 9, 1954. CODEN PRLAAZ. ISSN 0962-8444.

Infeld:1955:CCU

[IP55a] Leopold Infeld and Jerzy Plebański. On a certain class of unitary transformations. Acta Physica Polonica, 14:41–75, 1955. CODEN APPOAK. ISSN 0587-4254 (print), 1509-5770 (electronic).

Infeld:1955:UTS

[IP55b] Leopold Infeld and Jerzy Plebański. Unitary transformations and spinor calculus. Bulletin de l’Académie polonaise des sciences. Classe troisième, 3:95–99, 1955. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1956:ESF

[IP56a] Leopold Infeld and Jerzy Plebański. Expansion of singular functions associated with the Klein–Gordon equations. Acta Physica Polonica, 15(??):207–219, ???? 1956. CODEN APPOAK. ISSN 0001-673X.

Infeld:1956:CFE

[IP56b] Leopold Infeld and Jerzy Plebański. On a covariant formulation of the equations of motion. Bulletin de l’Académie polonaise des sciences. Classe troisième, 4:757–762, 1956. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1956:OMS

[IP56c] Leopold Infeld and Jerzy Plebański. On an operational method of solving the Klein–Gordon equation. Bulletin de l’Académie polonaise des sciences. Classe troisième, 4:215–219, 1956. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1956:MDF

[IP56d] Leopold Infeld and Jerzy Plebański. On modified Dirac δ-functions. Bulletin de l’Académie polonaise des sciences. Classe troisième, 4: 687–691, 1956. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1956:DPG

[IP56e] Leopold Infeld and Jerzy Plebański. On the “dipole procedure” in general relativity theory. Bulletin de l’Académie polonaise des sciences. Classe troisième, 4:763–767, 1956. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1956:SDE

[IP56f] Leopold Infeld and Jerzy Plebański. A simple derivation of the equations of motion in classical electrodynamics. Bulletin de l’Académie polonaise des sciences. Classe troisième, 4:347–351, 1956. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1957:FMD

[IP57] Leopold Infeld and Jerzy Plebański. On a further modification of Dirac’s δ-functions. Bulletin de l’Académie polonaise des sciences. Classe troisième, 5:51–54, vi, 1957. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1960:MR

[IP60] Leopold Infeld and Jerzy Plebański. Motion and Relativity. Polish Academy of Sciences. Physical monographs. Pergamon Press, Oxford, UK, 1960. ISBN 1-4832-8038-1 (print), 0-08-009436-8 (e-book). 229 pp. LCCN QC6 .I57 1960. URL http://www.

sciencedirect.com/science/book/9780080094366;   https://
www.elsevier.com/books/motion-and-relativity/infeld/978-
0-08-009436-6.

Infeld:1962:DRD

[IP62a] Leopold Infeld and Jerzy Plebański. Dvizhenie i reljativizm: dvizhenie tel v obshhej teorii otnositel’nosti. (Russian) [Motion and relativity: the motion of bodies in the general theory of relativity]. Izdvo Inostrannoj Literatury, Moskva, USSR, 1962. 204 pp. LCCN

???? URL http://booksshare.net/index.php?id1=4&category=
physics&author=infeld-l&book=1962&page=63.

Infeld:1962:MR

[IP62b] Leopold Infeld and Jerzy Plebański. Motion and relativity. Verlag für ausländische Literatur, Moscow, USSR, 1962. 204 pp. Translation by D. V. Belov from English original. Edited by D. Ivanenko.

Infeld:1949:CTW

[IPS^{+}49] L. Infeld, J. R. Pounder, A. F. Stevenson, W. Z. Chien, and J. L. Synge. Contributions to the theory of wave guides. Canadian Journal of Research, Section A: Physical Sciences, 27a(4):69–129, July 1949. CODEN CNRAAU. ISSN 0366-7383.

Infeld:1989:NWS

[IR89] E. (Eryk) Infeld and G. (George) Rowlands. Nonlinear Waves, Solitons, and Chaos. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-32111-5 (hardcover), 0-521-37937-7 (paperback). xi + 423 pp. LCCN QC174.26.W28 I55 1989.

Infeld:1931:ISC

[IS31] Leopold Infeld and S. Szczeniowski. The influence of space charge on the structure of de Broglie waves. Acta Physica Polonica, 1(??): 37–46, ???? 1931. CODEN APPOAK. ISSN 0001-673X.

Infeld:1945:NAK

[IS45a] Leopold Infeld and Alfred Schild. A new approach to kinematic cosmology. Physical Review (2), 68(11–12):250–272, December 1 and 15, 1945. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://adsabs.harvard.edu/abs/1945PhRv..
.68..250I.

Infeld:1945:NKP

[IS45b] Leopold Infeld and Alfred Schild. A note on the Kepler problem in a space of constant negative curvature. Physical Review (2), 67 (3–4):121–122, 1945. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Infeld:1946:NAK

[IS46] Leopold Infeld and Alfred E. Schild. A new approach to kinematic cosmology. B. Physical Review (2), 70(5–6):410–425, 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://adsabs.harvard.edu/abs/1946PhRv...70..410I.

Infeld:1949:GPF

[IS49a] Leopold Infeld and Alfred Schild. Geodesic postulate and field equations in g.r. [abstract only]. Physical Review, 75(??):1323, ???? 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Infeld:1949:MTP

[IS49b] Leopold Infeld and Alfred Schild. On the motion of test particles in general relativity. Reviews of Modern Physics, 21(3):408–413, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://adsabs.harvard.edu/abs/
1949RvMP...21..408I.

Infeld:1950:GRM

[IS50] Leopold Infeld and Adrian E. Scheidegger. Gravitational radiation and motion [abstract only]. Physical Review, 75(??):201, ???? 1950. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Infeld:1951:RGE

[IS51] Leopold Infeld and Adrian E. Scheidegger. Radiation and gravitational equations of motion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(2):195–207, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Infeld:1956:LGP

[IS56] Leopold Infeld and J. L. Synge. Letter: Gap problem in antenna theory. Journal of Applied Physics, 27(3):310, 1956. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850.

Infeld:1947:SSB

[ISC47] Leopold Infeld, V. G. Smith, and W. Z. Chien. On some series of Bessel functions. Journal of Mathematics and Physics / Massachusetts Institute of Technology, 26(1–4):22–28, April 1947. CO-DEN JMPHA9. ISSN 0097-1421.

Infeld:1949:PRS

[ISS49] L. Infeld, A. F. Stevenson, and J. L. Synge. Part I: Radiation from a source inside a perfectly conducting wave guide of rectangular section. Canadian Journal of Research, Section A: Physical Sciences, 27a(4):70–91, July 1949. CODEN CNRAAU. ISSN 0366-7383.

Infeld:1933:WEE

[IvdW33] Leopold Infeld and Bartel Leendert van der Waerden. Die Wellengleichung des Elektrons in der allgemeinen Relativitätstheorie. (German) [The wave equation of the electron in the general relativity theory]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, pages 380–402, 1933. CODEN SPWPAI. ISSN 0371-2435.

Infeld:1940:EME

[IW40] Leopold Infeld and P. R. Wallace. The equations of motion in electrodynamics. Physical Review (2), 57(9):797–806, May 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Kac:1979:BRM

[Kac79] Mark Kac. Book review: Memoir of a physicist: Why I Left Canada. Reflections on Science and Politics, Leopold Infeld. Science, 204 (4388):49–50, 1979. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.jstor.org/stable/1747967.

Kaempffert:1938:BRE

[Kae38] Waldemar Kaempffert. Book review: Einstein Himself Explains: With Remarkable Clarity and Simplicity He and His Collaborator Trace the Ideas Behind Relativity. The Evolution of Physics. The Growth of Ideas Front Early Concepts to Relativity and Quanta. By Albert Einstein and Leopold Infeld. x + 319 pp. New York: Simon & Schuster. $2.50. New York Times, ??(??):86, April 10, 1938. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

URL http://search.proquest.com/hnpnewyorktimes/docview/
102650329/.

Kotarbinski:1960:PIK

[KI60] Tadeusz Kotarbiński and Leopold Infeld. Przykład indywidualny kształtowania sie postawy wolnomyślicielskiej / Religia i ja. (Polish). [An example of individual formation of a free-thinking attitude/ Religion and I]. Ksiaźka i Wiedza, Warszawa, Poland, 1960. 47 pp. LCCN ????

Kiessling:2011:CPP

[Kie11] M. K.-H. Kiessling. Convergent perturbative power series solution of the stationary Maxwell–Born–Infeld field equations with regular sources. Journal of Mathematical Physics, 52(2):022902, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i2/p022902_s1.

Kotarbinski:1962:PIK

[KIR62] Tadeusz Kotarbiński, Leopold Infeld, and Bertrand Russell. Przykład indywidualny kształtowania sie postawy wolnomyślicielskiej / Religia i ja / Dlaczego nie jestem chrześcijaninem (Polish). [An example of individual formation of a free-thinking attitude / Religion and I / Why I am not a Christian]. Ksiaźka i Wiedza, Warszawa, Poland, second edition, 1962. 74 pp. LCCN ???? Translation to Polish by Amelia Kurlandzka of an as-yet-unknown English original.

Kotarbinski:1981:RJP

[KIR81] Tadeusz Kotarbiński, Leopold Infeld, and Bertrand Russell. Religia i ja. (Polish). [Religion and I]. Ksiaźka i Wiedza, Warszawa, Poland, fourth edition, 1981. ISBN 83-05-11450-3. 77 + 3 pp. LCCN ????

Kisin:2007:WLG

[Kis07] Mark Kisin. What is … a Galois representation? Notices of the American Mathematical Society, 54(6):718–719, June/July 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kostant:1995:EGT

[Kos95a] Bertram Kostant. Erratum: “The graph of the truncated icosahedron and the last letter of Galois”. Notices of the American Mathematical Society, 42(10):1118–1119, October 1995. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). See [Kos95b].

Kostant:1995:GTI

[Kos95b] Bertram Kostant. The graph of the truncated icosahedron and the last letter of Galois. Notices of the American Mathematical Society, 42(9):959–968, September 1995. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). See erratum [Kos95a].

Lenzen:1939:BRE

[Len39] V. F. Lenzen. Book review: The Evolution of Physics by Albert Einstein; Leopold Infeld. Isis, 30(1):124–125, February 1939. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/225593.

Li:2016:HRT

[Li16] Gu-Qiang Li. Hawking radiation via tunneling from Born–Infeld AdS black hole. Canadian Journal of Physics = Journal canadien de physique, 94(12):1369–1371, 2016. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (electronic).

Lie:1895:IGD

[Lie95] Sophus Lie. Influence de Galois sur le développement des mathématiques. (French) [Influence of Galois on the development of mathematics]. In Souvenirs du Centenaire de l’École Normale: 1795–1895, page ?? Librairie Hachette, Paris, France, 1895.

Lieber:1932:GTG

[Lie32] Lillian R. (Lillian Rosanoff) Lieber. Galois and the theory of groups. The Science Press Printing Company, Lancaster, PA, USA, 1932. 4 + 58 + 4 pp. LCCN QA171 .L75 1932.

Livio:2005:ECS

[Liv05] Mario Livio. The Equation That Couldn’t Be Solved: How Mathematical Genius Discovered the Language of Symmetry. Simon and Schuster, New York, NY, USA, 2005. ISBN 0-7432-5820-7 (hardcover), 0-7432-5821-5 (e-book), 0-7432-7462-8 (e-book). x + 353 pp. LCCN QA174.2 .L58 2005.

Lizon:2016:UDE

[Liz16] Francisco J. Sánchez Lizón. El último duelo de Évariste Galois. (Spanish) [The final duel of Évariste Galois]. Tandaia, O Milladoiro, Ames, A Coruñ a, Spain, 2016. ISBN 84-945443-1-4. 334 pp. LCCN ????

Lombardo:2005:NSS

[Lom05] D. J. Cirilo Lombardo. New spherically symmetric monopole and regular solutions in Einstein–Born–Infeld theories. Journal of Mathematical Physics, 46(4):042501, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i4/p042501_s1.

Malisoff:1941:BRS

[Mal41] William Marias Malisoff. Book review: Scientist’s Progress: Quest. The Evolution of a Scientist. By Leopold Infeld. 342 pp. New York: Doubleday, Doran & Co. $3. New York Times, ??(??): BR18, March 23, 1941. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/105527374/.

Mallett:1979:HMI

[Mal79] Ronald L. Mallett. Higgs mechanism and the inverse Einstein– Infeld–Hoffman problem. Journal of Mathematical Physics, 20(5):

921–922, May 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i5/p921_s1.

McCrea:1963:BRR

[McC63] W. H. McCrea. Book review: Recent Developments in General Relativity, by S. Bazański and others. Mathematical Gazette, 47(361): 267–268, October 1, 1963. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mazharimousavi:2012:GSH

[MH12] S. Habib Mazharimousavi and M. Halilsoy. Ground state H-atom in Born–Infeld theory. Foundations of Physics, 42(4):524–530, April 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1007/
s10701-011-9623-7.

Montagu:1941:BRQ

[Mon41] M. F. Ashley Montagu. Book review: Quest. The Evolution of a Scientist by Leopold Infeld. Isis, 33(3):344–345, September 1941. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330794.

Muller:2014:DIR

[MP14] Stefan Müller and Mariapia Palombaro. On a differential inclusion related to the Born–Infeld equations. SIAM Journal on Mathematical Analysis, 46(4):2385–2403, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

N:1938:RPA

[N.38] E. N. Book review: Physik als Abenteuer der Erkenntnis by Albert Einstein, Leopold Infeld. The Journal of Philosophy, 35 (18):500–501, September 1, 1938. ISSN 0022-362X (print), 1939-

8549 (electronic). URL http://links.jstor.org/sici?sici=
0022-362X(19380901)35:18<500:PAADE>2.0.CO%3B2-O;   http:
//www.jstor.org/stable/2017661.

Nagel:1948:BRU

[Nag48] Ernest Nagel. Book review: The Unhappy Life — Story of a Mathematical Genius: Whom the Gods Love. The Story of Évariste Galois. By Leopold Infeld. 393 pp. New York: Whittlesey House. $3.50. New York Times, ??(??):209, May 2, 1948. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/108212544/.

Nagel:1950:SRC

[Nag50] Ernest Nagel. Species and relativity: Charles Darwin. The Naturalist as a Cultural Force, by Paul B. Sears, and Albert Einstein. His Work and Its Influence on our World, by Leopold Infeld. New York Times, ??(??):BR4, February 26, 1950. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/111510278.

Natanson:1932:DLM

[Nat32] Władysław Natanson. Dzieciece lata Maxwella. (polish) [Ten years in Maxwell’s life]. Przeglad Współczesny, 126(??):??, October 1932.

Neumann:2011:MWE

[Neu11] Peter M. Neumann. The Mathematical Writings of Évariste Galois. European Mathematical Society, Zürich, Switzerland, 2011. ISBN 3-03719-104-X. x + 410 pp. LCCN QA3 .G18 2011.

Olsen:1961:NPP

[Ols61a] Arthur J. Olsen. Noted Polish physicist urges more freedom: Calls liberty indispensable for communist victory: Professor heads Warsaw’s Atomic Energy Agency. New York Times, ??(??): 6, November 27, 1961. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/115352270.

Olsen:1961:PSD

[Ols61b] Arthur J. Olsen. Polish scientists debate ‘freedom’: Intellectual ferment rising in non-political fields. New York Times, ?? (??):4, December 18, 1961. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/115296200.

Page:1965:BRI

[Pag65] Thornton Page. Book review: International Conference on the Theory of Gravitation, edited by Leopold Infeld. American Journal of Physics, 33(8):664, August 1965. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Penrose:1999:LHL

[Pen99] Sir Roger Penrose. Lecture in honour of Leopold Infeld (extended outline only): Spinors in general relativity. Acta Physica Polonica B, 30(10):2979–2987, October 1999. CODEN APOBBB. ISSN 0587-4254 (print), 1509-5770 (electronic). URL http://www.actaphys.

uj.edu.pl/fulltext?series=Reg&vol=30&page=2979.

Picard:1897:OUM

[Pic97] Émile Picard. Œuvres mathématiques d’Évariste Galois. (French) [Mathematical works of Évariste Galois]. Gauthier-Villars et cie, Paris, France, 1897. ???? pp. See also second edition [Pic51].

Picard:1951:OUM

[Pic51] Émile Picard. Œuvres mathématiques d’Évariste Galois. (French) [Mathematical works of Évariste Galois]. Gauthier-Villars et cie, Paris, France, 1951. ???? pp. Revision of [Pic97] with a new article by G. Verriest, Évariste Galois et la théories des equations. See also first edition [Pic97].

Pierpont:1898:EHG

[Pie98] James Pierpont. Early history of Galois’ theory of equations. Bulletin of the American Mathematical Society, 4(7):332–340, 1898. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Peng:2013:RPB

[PR13] Yue-Jun Peng and Jérémy Ruiz. Riemann problem for the Born–Infeld system without differential constraints. IMA Journal of Applied Mathematics, 78(1):102–131, February 2013. CO-DEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

URL http://imamat.oxfordjournals.org/content/78/1/102.
full.pdf+html.

Picard:2001:OUM

[PT01] Émile Picard and Jules Tannery. Œuvres Mathématiques d’Évariste Galois publiés in 1897 et Évariste Galois: Manuscrits et Papiers inédits publiés en 1906–1907. (French) [Mathematical Works of Évariste Galois, published in 1897 and Évariste Galois: Manuscripts and Unpublished Papers, published in 1906–1907]. Éditions Jacques Gabay, Paris, France, 2001. ???? pp.

Quinn:1940:BRE

[Qui40] R. B. Quinn. Book review: The Evolution of Physics, by Albert Einstein and Leopold Infeld. Popular Astronomy, 48:286, Decem-

ber 1940. URL http://adsabs.harvard.edu/abs/1940PA.....
48..286E;   http://adsabs.harvard.edu/abs/1940PA.....48.
.286Q.

Russell:1963:AAW

[RE63] Bertrand Russell and Albert Einstein. An appeal for the abolition of war. In Grodzins and Rabinowitch [GR63], pages 539–541. LCCN D842 .B78. Reprint of [BEI^{+}55b].

Richmond:1995:KQ

[Ric95] Theo Richmond. Konin: a quest. Jonathan Cape, London, UK, 1995. ISBN 0-224-03890-7. xvii + 1 + 520 + 8 pp. LCCN ????

Richmond:1996:KQ

[Ric96] Theo Richmond. Konin: a quest. Vintage Books, New York, NY, USA, 1996. ISBN 0-09-940981-X (hardcover). xxi + 520 pp. LCCN ????

Richmond:1997:KSN

[Ric97] Theo Richmond. Konin: auf der Suche nach der Stadt meiner Eltern. (German) [Konin: looking for the city of my parents]. C. Bertelsmann, München, Germany, 1997. ISBN 3-570-12276-X. 573 + 16 pp. LCCN ???? Translation to German by Elke Hosfeld of [Ric95].

Richmond:1999:LIH

[Ric99] Theo Richmond. Leopold Infeld: hope lost, hope regained. Acta Physica Polonica B, 30(10):2863–2873, October 1999. CO-DEN APOBBB. ISSN 0587-4254 (print), 1509-5770 (electronic).

URL http://www.actaphys.uj.edu.pl/fulltext?series=Reg&
vol=30&page=2863.

Robertson:1938:NPP

[Rob38] H. P. Robertson. Note on the preceding paper (“The gravitational equations and the problem of motion” [MR1503389] by A. Einstein, Leopold Infeld and B. Hoffmann): the two body problem in General Relativity. Annals of Mathematics (2), 39(1):101–104, 1938. CO-DEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic).

Roeder:1981:BRQ

[Roe81] R. C. Roeder. Book review: Quest — an Autobiography. Journal of the Royal Astronomical Society of Canada, 75:268–270, October 1981. CODEN JRASA2. ISSN 0035-872X. URL http://adsabs.

harvard.edu/abs/1981JRASC..75..268R.

Roxburgh:2007:SHB

[Rox07] I. W. Roxburgh. Sir Hermann Bondi KCB. 1 November 1919– 10 September 2005: Elected FRS 1959. Biographical Memoirs of Fellows of the Royal Society, 53:45–61, December 2007. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2007.0008.

Rapport:1964:P

[RW64] Samuel Rapport and Helen Wright, editors. Physics. New York University Press, New York City, NY, USA, 1964. xiii + 333 pp.

Salvia:2019:ECP

[Sal19] Stefano Salvia. Embattled cooperation(s): Peaceful atoms, pacifist physicists, and partisans of peace in the Early Cold War (1947– 1957). Physics in Perspective (PIP), 21(1):43–62, March 2019. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Sarton:1937:EG

[Sar37] George Sarton. Évariste Galois. Osiris, 3(??):241–259, ???? 1937. CODEN OSIRAX. ISSN 0369-7827 (print), 1933-8287 (electronic).

URL http://www.jstor.org/stable/301589.

Stevenson:1952:PFD

[SB52] A. F. Stevenson and W. A. Bassali. On the possible forms of differential equation which can be factorized by the Schrödinger–Infeld method. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):385–395, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Seitz:1961:PSE

[SBS^{+}61] Frederick Seitz, Peter G. Bergmann, S. F. Singer, Howard Laster, George Weiss, Derek J. De Solla Price, W. James Lyons, Elizabeth Cuthill, William D. O’Gorman, William D. O’Gorman, and William D. O’Gorman. Physical sciences and engineering. Science, 133(3460):1247–1251, April 21, 1961. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/1707254.

Sawilowsky:2006:JLM

[SC06] Shlomo S. Sawilowsky and John L. Cuzzocrea. Joseph Liouville’s ‘Mathematical Works of Évariste Galois’. Journal of Modern Applied Statistical Methods, 5(2):??, ???? 2006. CODEN ???? ISSN

1538-9472. URL https://jmasm.com/index.php/jmasm/article/
view/275.

Schilpp:1949:AEPa

[Sch49a] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher-Scientist, volume 1 of The Library of Living Philosophers. Cambridge University Press, Cambridge, UK, 1949. ISBN 0-87548-286-4. ISSN 0075-9139. xviii + 781 pp. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Schilpp:1949:AEPb

[Sch49b] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher-Scientist, volume 2 of The Library of Living Philosophers. Cambridge University Press, Cambridge, UK, 1949. ISBN 0-87548-286-4. ISSN 0075-9139. xviii + 781 pp. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Schlegel:1950:BRL

[Sch50a] Richard Schlegel. Book review: Leopold Infeld, Albert Einstein: His Work and Its Influence On Our World. American Journal of Physics, 18(5):327, May 1950. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.aip.org/link/

?AJP/18/327/1.

Schrag:1950:MHF

[Sch50b] Lex Schrag. Mathematician hard to figure: Prof. Infeld in Poland but doubt he took A-secrets. Globe and Mail, ??(??):??, September 22, 1950. Reprinted in [Inf78e, pages 57–59].

Schwartz:1954:RBB

[Sch54] Harry Schwartz. Red bloc building own atomic ‘pool’: Soviet has a virtual monopoly and benefits most — China may have reactor plant. New York Times, ??(??):3, September 7, 1954. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http:

//search.proquest.com/hnpnewyorktimes/docview/112937767.

Szczeniowski:1931:ICEa

[SI31] S. Szczeniowski and L. Infeld. The influence of a cloud of electrons on the structure of de Broglie waves. Bulletin International de l’Académie Polonaise des Sciences et des Lettres. Classe des sciences mathématiques et naturelles. Série A: Sciences mathématiques, 1931 (6):482–488, 1931. CODEN BIAPAC. ISSN 0366-0303.

Szczeniowski:1932:WCE

[SI32] S. E. Szczeniowski and Leopold Infeld. O wpływie chmury elektronowej na strukture fali de Broglie’a. (Polish). [On the effect of the electron cloud on de Broglie’s wave structure]. Acta Physica Polonica, 1(1–2):37–46, ???? 1932. CODEN APPOAK. ISSN 0001-673X.

Schild:1949:GPF

[SI49] Alfred Schild and Leopold Infeld. Geodesic postulate and field equations in general relativity. Physical Review, 75(8):1323, 1949. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Scheidegger:1950:GRM

[SI50] Adrian E. Scheidegger and Leopold Infeld. Gravitational radiation and motion. Physical Review, 79(1):201, 1950. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Silver:2010:DDH

[Sil10] Daniel S. Silver. Duel at dawn: heroes, martyrs, and the rise of modern mathematics [book review of MR2654369]. Notices of the American Mathematical Society, 57(10):1297–1299, November 2010. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/notices/201302/rnoti-p154.pdf.

Speck:2012:NST

[Spe12] Jared Speck. The nonlinear stability of the trivial solution to the Maxwell–Born–Infeld system. Journal of Mathematical Physics, 53 (8):083703, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i8/p083703_s1.

Stachel:1999:EIS

[Sta99] John Stachel. Einstein and Infeld, seen through their correspondence. Acta Physica Polonica B, 30(10):2879–2908, October 1999. CODEN APOBBB. ISSN 0587-4254 (print), 1509-5770

(electronic). URL http://www.actaphys.uj.edu.pl/fulltext?

series=Reg&vol=30&page=2879. The appendix Scientific Writings of Leopold Infeld is an shortened version of the annotated bibliography in [Inf70].

Stewart:1982:BRQ

[Ste82] Albert B. Stewart. Book review: Quest. The Evolution of a Scientist by Leopold Infeld. The Antioch Review, 40(1):113, 1982. ISSN 0003-5769 (print), 2326-9707 (electronic). URL http://www.jstor.org/

stable/4638538.

Suffczynski:1953:ETW

[STW53] M. Suffczyński, R. Teisseyre, and J. Werle. Elektrodynamika teoretyczna / według wykładów Leopolda Infelda oprac. (Polish) [Theoretical electrodynamics / according to lectures by Leopold Infeld]. Uniwersytet Warszawsk, Warszawa, Poland, 1953. 334 pp. LCCN ????

Suffczynski:1950:EOW

[SW50] Maciej Suffczyński and Józef Werle. Elektrodynamika i optyka: wg wykładów Leopolda Infelda oprac. (Polish) [Electrodynamics and optics / According to Leopold Infeld’s lectures], volume 1. Uniwersytet Warszawsk, Warszawa, Poland, 1950. 263 + 2 pp. LCCN ????

Schneider:1979:LOG

[SW79] C. K. E. Schneider and Raj Wilson. Ladder operators of group matrix elements. Journal of Mathematical Physics, 20(12):2380– 2390, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i12/p2380_s1.

Szczeniowski:1933:RIL

[Szc33] S. Szczeniowski. [recenzja:] Infeld, Leopold dr., doc. Uniw. J. K. Nowe Drogi Nauki. Kwanty i materja. Str. x + 284 z 27 fig. w tekście i na 3 tabl. Nakładem “Mathesis Polskiej”. Warszawa. 1933. Mathesis Polska, 8(5–6):90–91, ???? 1933.

Sears:1953:MMS

[SZI53] Paul Bigelow Sears, Gregory Zilboorg, and Leopold Infeld. Makers of modern science: a twentieth century library trilogy. Scribner, New York, NY, USA, 1953. 124 + 132 + 134 pp. LCCN ???? Edited by Hiram Collins Haydn.

Tannery:1906:MPIa

[Tan06a] Jules Tannery. Manuscrits et papiers inédits de Évariste Galois. (French) [manuscripts and unpublished papers of Évariste Galois]. Bulletin des Sciences Mathématiques, XXX(??):226–248, 1906.

Tannery:1906:MPIb

[Tan06b] Jules Tannery. Manuscrits et papiers inédits de Évariste Galois. (French) [manuscripts and unpublished papers of Évariste Galois]. Bulletin des Sciences Mathématiques, XXX(??):225–263, 1906.

Tannery:1907:MPI

[Tan07] Jules Tannery. Manuscrits et papiers inédits de Évariste Galois (suite). (French) [manuscripts and unpublished papers of Évariste Galois (continued)]. Bulletin des Sciences Mathématiques, XXX, XXXI(??):275–308, 1907.

Tannery:1908:MEG

[Tan08] Jules Tannery. Manuscrits de Évariste Galois. (French) [Manuscripts of Évariste Galois]. Gauthier-Villars et cie, Paris, France, 1908. ???? pp. Includes [Tan06a, Tan06b, Tan07].

Taylor:1979:BRW

[Tay79] Loretta M. Taylor. Book review: Whom the Gods Love: The Story of Évariste Galois by Leopold Infeld. The Mathematics Teacher, 72 (1):76, 1979. ISSN 0025-5769 (print), 2330-0582 (electronic). URL

http://www.jstor.org/stable/27961535.

Tempczyk:2001:LIP

[Tem01] Michał Tempczyk. Leopold Infeld: The problem of matter and field. Poznań Studies in the Philosophy of the Sciences and the Humanities, 74(??):207, 2001. ISSN 0303-8157 (print), 1875-7243 (electronic).

Tits:1982:EGS

[Tit82] Jacques Tits. Évariste Galois, son œuvre, sa vie, ses rapports avec l’Académie: exposé fait en la séance du 7 juin 1982 à l’occasion du 150e anniversaire de la mort d’Évariste Galois. (French) [E‘variste Galois, his work, his life, his relations with the Academy: presentation made at the meeting of 7 June 1982 on the occasion of the 150th Anniversary of the death of Évariste Galois]. Institut de France, Paris, France, 1982. 10 pp. LCCN ????

TotiRigatelli:1993:MSB

[Tot93] Laura Toti Rigatelli. Matematica sulle barricate: vita di Évariste Galois. (Italian) [Mathematics on barricades: the life of Évariste Galois]. Sansoni saggi. Sansoni, Firenze, Italia, 1993. ISBN 88-383-1512-4. 143 pp. LCCN QA29.G25 T67 1993.

TotiRigatelli:1996:EG

[Tot96] Laura Toti Rigatelli. Évariste Galois, 1811–1832, volume 11 of Vita mathematica. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1996. ISBN 3-7643-5410-0 (Basel), 0-8176-5410-0 (Boston). 162 pp. LCCN QA29.G25 T6713 1996.

Troutman:1979:BRW

[Tro79] Andria P. Troutman. Book review: Whom the Gods Love: The Story of Évariste Galois by Leopold Infeld. Arithmetic Teacher, 27 (4), 1979. CODEN ARITBF. ISSN 0004-136X (print), 2327-0756

(electronic). URL http://www.jstor.org/stable/41189512.

Vazquez:1977:LSNa

[Vaz77] Luis Vazquez. Localized solutions of a nonlinear electromagnetic field. Journal of Mathematical Physics, 18(6):1259–1263, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/
i6/p1259_s1.

Verriest:1934:EGT

[Ver34] Gustave Verriest. Évariste Galois et la théorie des équations algébriques. (French) [Évariste Galois and the theory of algebraic equations]. Gauthier-Villars et cie, Paris, France, 1934. 56 pp.

Krzywoblocki:1963:RPP

[vK63] M. Z. v. Krzywoblocki. Recent publications and presentations: Recent Developments in General Relativity. American Mathematical Monthly, 70(9):1029–1030, November 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Weber:2011:EGO

[WA11] Anne-Gaëlle Weber and Andrea Albrecht. Évariste Galois ou le roman du mathématicien. (French) [Évariste Galois or the mathematician’s novel]. Revue d’Histoire des Math’ematiques = Journal for the History of Mathematics, 17(2):403–435, ???? 2011. ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/Publications/RevueHistoireMath/17/pdf/smf_rhm_17_403-
435.pdf.

Whitney:1955:NNS

[Whi55] Peter D. Whitney. Nine noted scientists urge war ban: Scientists ask abolition of war; Einstein among signers of plea. Warning on nuclear peril was signed by Einstein. New York Times, ??(??):1, July 1955. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

URL http://search.proquest.com/hnpnewyorktimes/docview/
113191041.

Werle:1968:LIH

[WHS68] J. Werle, F. Hund, and F. Sauter. Leopold Infeld / Hertha Sponer-Franck / Josef Meixner 60 Jahre. Physikalische Blätter, 24(4):165– 167, April 1968. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Wolf:1991:CSB

[Wol91] C. Wolf. Charge screening in Born–Infeld theory by magnetic charge of another U(1) gauge group. Canadian Journal of Physics = Journal canadien de physique, 69(10):1304–1308, 1991. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (electronic).

Wolf:1993:BIS

[Wol93] C. Wolf. Born–Infeld solution for a point charge in D dimensions. Canadian Journal of Physics = Journal canadien de physique, 71 (9–10):484–487, 1993. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (electronic).